### v1

Abstract
Deep neural networks (DNNs) have exhibited substantial success in the field of Natural Language Processing (NLP). As these systems are increasingly integrated into real-world applications, ensuring their safety and reliability becomes a primary concern. There are safety critical contexts where such models must be robust to variability or attack, and give guarantees over their output. Computer Vision had pioneered the use of formal verification for neural networks for such scenarios and developed common verification standards and pipelines. In contrast, NLP verification methods have only recently appeared in the literature. While presenting sophisticated algorithms on their own right, these papers have not yet crystallised into a common methodology, they are often light on the pragmatical issues of NLP verification, and the area remains fragmented.
In this paper, we make an attempt to distil and evaluate general components of an NLP verification pipeline, that emerges from the progress in the field to date. Our contributions are two-fold.
Firstly, we give a general (i.e. algorithm-independent) characterisation of verifiable subspaces that result from embedding sentences into continuous spaces. We identify, and give an effective method to deal with, the technical challenge of semantic generalisability of verified subspaces; and propose it as a standard metric in the NLP verification pipelines (alongside with the standard metrics of model accuracy and model verifiability).
Secondly, we propose a general methodology to analyse the effect of the embedding gap – a problem that refers to the discrepancy between verification of geometric subpspaces on the one hand, and semantic meaning of sentences which the geometric subspaces are supposed to represent, on the other hand. In extreme cases, poor choices in embedding of sentences may invalidate verification results. We propose a number of practical NLP methods that can help to identify the effects of the embedding gap; and in particular we propose the metric of falsifiability of semantic subpspaces that we propose as another fundamental metric to be reported as part of the NLP verification pipeline.
We believe that together these general principles pave the way towards a more consolidated and effective development of this new domain.
0.1 Introduction
Deep neural networks (DNNs) have demonstrated remarkable success at addressing challenging problems in various areas, such as Computer Vision (CV) [102] and Natural Language Processing (NLP) [113, 49]. However, as DNN-based systems are increasingly deployed in safety-critical applications [10, 125, 11, 34, 13, 35], ensuring their safety and security becomes paramount. Current NLP systems cannot guarantee either truthfulness, accuracy, faithfulness, or groundedness of outputs given an input query, which can lead to different levels of harm.
Contexts which necessitate guaranteed outputs.
One such example in the NLP domain is the requirement of a chatbot to correctly disclose non-human identity, when prompted by the user to do so. Recently there have been several pieces of legislation proposed that will enshrine this requirement in law [63, 68]. In order to be compliant with these new laws, in theory the underlying DNN of the chatbot (or the sub-system responsible for identifying these queries) must be 100% accurate in its recognition of such a query. However, a central theme of generative linguistics going back to von Humboldt, is that language is ‘an infinite use of finite means’, i.e there exists many ways to say the same thing. In reality the questions can come in a near infinite number of different forms, all with similar semantic meanings. For example: “Are you a Robot?”, “Am I speaking with a person?”, “Am i texting to a real human?”, “Aren’t you a chatbot?”. Failure to recognise the user’s intent and thus failure to answer the question correctly could potentially have legal implications for designers of these systems [63, 68].
Similarly, as such systems become widespread in their use, it may be desirable to have guarantees on queries concerning safety critical domains, for example when the user asks for medical advice. Research has shown that users tend to attribute undue expertise to systems [2, 34] potentially causing real world harm [12] (e.g. ‘Is it safe to take these painkillers with a glass of wine?’). However, a question remains on how to ensure that NLP systems can give formally guaranteed outputs, particularly for scenarios that require maximum control over the output.
Formal verification of neural networks.
[…]
Challenges of NLP verification.
However, existing verification approaches primarily focus on computer vision (CV) tasks, where images are seen as vectors in a continuous space and every point in the space corresponds to a valid image. In contrast, sentences in NLP form a discrete domain [^2] In this paper, we work with textual representations of sentences. Raw audio input can be seen as continuous, but this is out of scope of this paper., making it challenging to apply traditional verification techniques effectively. In particular, taking an NLP dataset $\mathcal{Y}$ to be a set of sentences $s_{1},\ldots,s_{q}$ written in natural language, an embedding $\mathbb{E}$ is a function that maps a sentence to a vector in ${}^{m}$. The resulting vector space is called the embedding space. Due to discrete nature of the set $\mathcal{Y}$, the reverse of the embedding function $\mathbb{E}^{-1}:^{m}\rightarrow\mathcal{Y}$ is not total or undefined. This problem is known as the “problem of the embedding gap". Sometimes, one uses the term to more generally refer to any discrepancies that $\mathbb{E}$ introduces, for example, when it maps dissimilar sentences close in ${}^{m}$. We use the term in both mathematical and NLP sense.
Mathematically, the general (geometric) “DNN robustness verification" approach of defining and verifying subspaces of ${}^{m}$ should work, and some prior works exploit this fact. However, pragmatically, because of the embedding gap only a tiny fraction of vectors contained in the verified subspaces maps back to valid sentences. When a verified subspace contains no or very few sentence embeddings, we say that verified subspace has low generalisability. Low generalisability may render verification efforts obsolete for practical applications.
From the NLP perspective, there are other, more subtle, examples when the embedding gap can manifest. Suppose we succeeded in verifying a DNN robust on some subspace of ${}^{m}$. Consider an example when the subspace contains sentences that are semantically similar to the sentence: i really like too chat to a human. are you one?. Verification gives us guarantees that the DNN will always identify these sentences as questions about human/robot identity. But suppose the embedding function $\mathbb{E}$ wrongly embedded sentences belonging to an opposite class into this subspace. For example, an LLM Vicuna generates the following sentence as a rephrase of the previous one: Do you take pleasure in having a conversation with someone?. Suppose our verified subspace contained an embedding of this sentence, too, and thus our verified DNN identifies this second sentence to belong to the the same class as the first one. However, the second sentence is not a question about human/robot identity of the agent! When we can find such an example, we say that it falsifies the verification guarantee for the subspace it is contained in. Alternatively, we say that the subpspace is falsifiable.
Contributions
Our main aim is to provide a general and principled verification methodology that bridges the embedding gap when possible; and gives precise metrics to evaluate and report its effects in any case. The contributions split into two main groups, depending whether the embedding gap is approached from mathematical or NLP perspective.
[…]
Contributions Part 1: Characterisation of verifiable subspaces and NLP verification pipeline.
[…]
An alternative method to purely geometric approach that suffers from the embedding gap is to construct subspaces of the embedding space based on the semantic perturbations of sentences [54, 50, 146]. Concretely, the idea is to form each $\mathcal{S}_{i}$ by embedding a sentence $i$ and its $n$ semantic perturbations into the real vector space and enclosing them inside some geometric shape. Ideally, such a shape should be given by a convex hull around these $n+1$ embedded sentences, however calculating convex hulls with sufficient precision is computationally infeasible for high number of dimensions. Thus, simpler shapes, such as hyper-cubes and hyper-rectangles are used in the literature. We propose a novel refinement of these ideas, by including the method of a hyper-rectangle rotation in order to increase the shape precision (see Figure 1). We will call the resulting shapes semantic subspaces (in contrast to those obtained purely geometrically).
A few questions have been left un-answered in the previous work. Firstly, because generalisability of the verified subspaces is not reported in the literature, we cannot know whether the prior semantically-informed approaches are better in that respect than purely geometric methods. If they are better in both verifiability and generalisability, it is unclear whether the improvement should be attributed to:
  • the fact that verified semantic subspaces simply have an optimal volume (for the verifiability-generalisability trade-off), or
  • the improved precision of verified subspaces that comes from using the semantic knowledge.
This paper provides a strong argument for including generalisability as a standard metric in reporting NLP verification results in the future. Through a series of experiments, we confirm that verified semantic subspaces are more verifiable and more generalisable than their geometric counterparts. Moreover, by comparing the volumes of the obtained verified semantic and geometric subspaces, we show that the improvement is partly due to finding an optimal size of subspaces (for the given embedding space), and partly due to improvement in shape precision.
The second group of unresolved questions concerns robust training regimes in NLP verification that is used as means of improving verifiability of subspaces in prior works [54, 50, 146]. It was not clear what made robust training successful:
  • was it because additional examples generally improved the precision of the decision boundary? (in which case data set augmentation would have a similar effect);
[…]
  • did the knowledge of semantic subspaces play the key role?
[…]
As a result, we arrive at a fully parametric approach to NLP verification that disentangles the four components:
  • choice of the semantic attack (on the NLP side),
  • semantic subspace formation in the embedding space (on the geometric side),
  • semantically robust training (on the machine learning side),
  • choice of the verification algorithm (on the verification side).
We argue that, together with the new generalisability metric, this approach opens the way for more principled evaluation of performance of NLP verification methods that accounts for the effects of the embedding gap; and generation of more transparent NLP verification benchmarks. We implement a tool that generates NLP verification benchmarks based on the above choices. This paper is the first to use a complete SMT-based verifier (namely Marabou [131]) for NLP verification.
Contributions Part 2: NLP Verification pipeline in use: an NLP perspective on the embedding gap.
We test the theoretical results by suggesting an NLP verification pipeline, a general methodology that starts with NLP analysis of the dataset and obtaining semantically similar perturbations that together characterise the semantic meaning of a sentence; proceeds with embedding of the sentences into the real vector space and defining semantic subspaces around embeddings of semantically similar sentences; and culminates with using these subspaces for both training and verification. This clear division into stages allows us to formulate practical NLP methods for minimising the effects of the embedding gap. In particular, we show that the quality of the generated sentence perturbations maybe improved through the use of human evaluation, cosine similarity and ROUGE-N. We introduce the falsifiability metric as an effective practical way to measure the quality of the embedding functions. Through a detailed case study, we show how geometric and NLP intuitions can be put together at work towards obtaining DNNs that are more verifiable over better generalisable and less falsifiable semantic subspaces. Perhaps more importantly, the proposed methodology opens the way for transparency in reporting NLP verification results, – something that this domain will benefit from if it reaches the stage of practical deplyoment of NLP verification pipelines.
Paper Outline.
From here, the paper proceeds as follows. Section 0.2 gives an extensive literature review encompassing DNN verification methods generally, and NLP verification methods in particular. The section culminates with distilling a common “NLP verification pipeline" encompassing the existing literature. Based on the understanding of major components of the pipeline, the rest of the paper focuses on improving understanding or implementation of its components. Section 0.3 formally defines the components of the pipeline in a general mathematical notation, which abstracts away from particular choices of sentence perturbation, sentence embedding, training and verification algorithms. The central notion the section introduces is that of geometric and semantic subpspaces. The next Section 0.4 makes full use of this general definition, and shows that semantic subpspaces play a pivotal role in improving verification and training of DNNs in NLP. This section formally defines the generalisability metric and considers the problem of generalisability-verifiability trade-off. Through thorough empirical evaluation, it shows that a principled approach to defining semantic subpaces can help to improve both generalisability and verifiability of DNNs, thus reducing the effects of the trade-off. The final Section 0.5 further tests the NLP verification pipelines using state-of-the-art NLP tools, and analyses the effects of the embedding gap from the NLP persective, in particular it introduces a method of measuring falsifiability of semantic subpaces and reporting this metric alongside verifiability and generalisability. Section 0.6 concludes the paper and discusses future work.
0.2 Related Work
0.2.1 DNN Verification
Formal verification is an active field across several domains including hardware [64, 97], software languages [58], network protocols [87] and many more [129], however it was only recently that this became applicable to the field of machine learning [59]. Several verifiers have been popular in DNNs verification and competitions [5, 6, 81, 108]. We can divide them into 2 main categories: complete and incomplete verifiers. When the verification approach guarantees that, if a query fails, then the query is false and it produces a counter-example to prove it, then we call it complete verification, otherwise we call it incomplete verification. Furthermore, while complete verifiers are always deterministic, incomplete verifiers may also be probabilistic. Unlike deterministic verification, probabilistic verification is guaranteed to output ‘not verified’ with a certain probability (e.g., 99.9%).
Complete verifiers
can be based on Satisfiability Modulo Theories (SMT), Mixed-Integer Linear Programming (MILP) or Branch-and-Bound (Bab). SMT-based verification [99, 59, 131] is built upon the observation that feed-forward neural networks are defined by the sequential composition of affine trasformations and ReLU operations. Both these transformations and operations can be encoded by a conjunction of linear inequalities, thus general-purpose SMT solvers can be directly applied to solve the satisfiability problem, yielding a solution to complete verification. A state-of-the-art SMT-based tool is Marabou [131], which answers queries about neural networks and their properties in the form of constraint satisfaction problems. Marabou takes as input networks with piece-wise linear activation functions and with fully connected topology. It first applies multiple pre-processing steps to infer bounds for each node in the network. Next it applies a combination of Simplex [31] search over linear constraints with SMT techniques directing the search over non-linear constraints.
MILP-based approaches [23, 82, 114] encode the verification problem as a mixed-integer linear programming problem, in which the constraints are linear inequalities and the objective is represented by a linear function. Differently from linear programming (LP), in MILP it is possible to constrain some variables to take only integer values instead of real numbers, allowing the constraints to encode the non-linear ReLU operations. Thus, the verification problem can be precisely encoded as an MILP problem. A representative tool for this category is ERAN [109], which is mainly based on abstract interpretation (see ‘Incomplete verifiers’ below ) but can also leverage the efficient MILP solver GUROBI [48]. ERAN combines abstract domains with custom multi-neuron relaxations to support fully-connected, convolutional, and residual networks with ReLU, Sigmoid, Tanh, and Maxpool activations. Both these methods suffer from scalability, which is their main limitation, but their strength is that they precisely encode the constraint.
BaB-based verification [43, 19, 18, 41, 57, 118, 140] relies on the piecewise-linear property of DNNs: since each ReLU neuron outputs ReLU($x$) = max{$x$, 00}, it is always locally linear within some region around input $x$. Furthermore, since feed-forward ReLU networks are the composition of these piecewise linear neurons and (linear) affine transformations, the output is locally linear w.r.t. input $x$. This property is formally stated and proved in [57] and it serves as the foundation for BaB verification. A BaB verification approach, as the name suggests, consists of two parts: branching and bounding. It first applies incomplete verification to derive a lower bound and an upper bound, then, if the lower bound is positive it terminates with ‘verified’, else, if the upper bound is non-positive it terminates with ‘not verified’ (bounding). Otherwise, the approach recursively chooses a neuron to split into two branches (branching), resulting in two linear constraints. Then bounding is applied to both constraints and if both are satisfied the verification terminates, otherwise the other neurons are split recursively. When all neurons are split, the branch will contain only linear constraints, and thus the approach applies linear programming to compute the precise constraint and verify the branch. Multi-Neuron Guided Branch-and-Bound (MN-BaB) [41] is a state-of-the-art neural network verifier that builds on the tight multi-neuron constraints proposed in PRIMA [95] and leverages these constraints within a BaB framework to yield an efficient, GPU based dual solver. Another state-of-the-art tool is $\alpha,\beta$-CROWN [133, 118], a neural network verifier based on an efficient linear bound propagation framework and branch-and-bound. It can be accelerated efficiently on GPUs and can scale to relatively large convolutional networks (e.g., $10^{7}$ parameters). It also supports a wide range of neural network architectures (e.g., CNN, ResNet, and various activation functions). BaB-based methods are more scalable than solver-based approaches, however they introduce a level of abstraction and sacrifice precision in favor of scalability. For example GCP-CROWN [140] extracts convex constraints from MILP solvers and integrates them in linear inequality propagation, which can be viewed as leveraging multi-neuron relaxations in branch-and-bound complete verification.
Deterministic incomplete verifiers
mainly use linear relaxations on ReLU neurons, resulting in an over-approximation of the initial constraint. In general, they define a lower bound and an upper bound of the output of each ReLU neuron as linear constraints, which define a region called ReLU polytope that gets propagated through the network. A predominant approach is the use of interval bound propagation (IBP) [128, 46, 83, 88]. The strength of IBP-based methods lies in their efficiency; they are faster than alternative approaches and demonstrate superior scalability. However, their primary limitation lies in the inherently loose bounds they produce [46]. This drawback becomes particularly pronounced in the case of deeper neural networks, typically those with more than 10 layers [73], where they cannot certify non-trivial robustness due to the amplification of over-approximation. Other methods that are less efficient but produce tighter bounds are based on polyhedra abstraction, such as CROWN [141] and DeepPoly [111], or based on multi-neuron relaxation, such as PRIMA [93]. One of the most mature tool in this category is ERAN [109], which can be used for complete verification, but its main purpose is deterministic incomplete verification through abstract interpretation (DeepPoly) and multi-neuron relaxation (PRIMA).
Probabilistic incomplete verification
approaches add random noise to smooth models, and then derive certified robustness for these smoothed models. This field is commonly referred to as Randomised Smoothing, given that these approaches provide probabilistic guarantees of robustness, and all current probabilistic verification techniques are tailored for smoothed models [67, 71, 38, 138, 103, 89]. Given that this work focuses on deterministic approaches, here we only report the existence of this line of work without going into details.
Note that these existing verification approaches primarily focus on computer vision tasks, where images are seen as vectors in a continuous space and every point in the space corresponds to a valid image, while sentences in NLP form a discrete domain, making it challenging to apply traditional verification techniques effectively.
In this work we use both an abstract interpretation-based incomplete verifier (ERAN [109]) and an SMT-based complete verifier (Marabou [131]) in order to demonstrate the effect that the choice of a verifier may bring, and demonstrate common trends.
0.2.2 Robust Training
Verifying DNNs poses significant challenges if they are not appositely trained. The fundamental issue lies in the failure of DNNs, including even sophisticated models, to meet essential verification properties, such as robustness [22]. To enhance robustness, various training methodologies have been proposed. It is noteworthy that, although robust training by projected gradient descent [45, 85, 62] predates verification, contemporary approaches are often related to, or derived from, the corresponding verification methods by optimizing verification-inspired regularization terms or injecting specific data augmentation during training. In practice, after robust training, the model usually achieves higher certified robustness and is more likely to satisfy the desired verification properties [22]. Thus, robust training is a strong complement to robustness verification approaches.
Robust training
techniques can be classified into several large groups:
  • data augmentation [100],
  • adversarial training [45, 85],
  • IBP training [46, 139] and other forms of certified training [94], or
  • a combination thereof [145].
Data augmentation involves the creation of synthetic examples through the application of diverse transformations or perturbations to the initial training data. These generated instances are then incorporated into the original dataset to enhance the training process. Adversarial training entails identifying worst-case examples at each epoch during the training phase and calculating an additional loss on these instances. State of the art adversarial training involve projected gradient descent algorithms such as FGSM [45] and PGD [85]. Certified training methods focus on providing mathematical guarantees about the model’s behaviour within certain bounds. Among them, we can name IBP training [46, 139] techniques, which impose intervals or bounds on the predictions or activations of the model, ensuring that the model’s output lies within a specific range with high confidence.
Note that all techniques mentioned above can be categorised based on whether they primarily augment the data (such as data augmentation) or augment the loss function (as seen in adversarial, IBP and certified training). Augmenting the data tends to enhance generalisation and is efficient, albeit it may not help against the most severe adversarial attacks. Conversely, methods that manipulate the loss functions directly confront the toughest adversarial attacks but often come with higher computational costs. Ultimately, the choice between altering data or loss functions depends on the specific requirements of the application and the desired trade-offs between performance, computational complexity, and robustness guarantees.
NLP robustness.
There exists a substantial body of research dedicated to enhancing the adversarial robustness of NLP systems [142, 121, 122, 75, 149, 150, 36]. These efforts aim to mitigate the vulnerability of NLP models to adversarial attacks and improve their resilience in real-world scenarios [121, 122] and mostly employ data augmentation techniques [40, 33]. In NLP, perturbations can occur at the character, word, or sentence level [24, 52, 20] and may involve deletion, insertion, swapping, flipping, substitution with synonyms, concatenation with characters or words, or insertion of numeric or alphanumeric characters [76, 39, 69]. For instance, in character level adversarial attacks, [9] introduces natural and synthetic noise to input data, while  [42, 72] identify crucial words within a sentence and perturbs them accordingly. For word level attacks, they can be categorised into gradient-based [76, 104], importance-based [51, 56], and replacement-based [3, 66, 98] strategies based on the perturbation method employed. In addition, in sentence level adversarial attacks, some attacks [53, 124] are created so that they do not impact the original label of the input and can be incorporated as a concatenation in the original text. In such scenarios, the expected behaviour from the model is to maintain the original output, and the attack can be deemed successful if the label/output of the model is altered. By augmenting the training data with these perturbed examples, models are exposed to a more diverse range of linguistic variations and potential adversarial inputs. This helps the models to generalise better and become more robust to different types of adversarial attacks. To help with this task, the NLP community has gathered a dataset of adversarial attacks named AdvGLUE [117], which aims to be a principled and comprehensive benchmark for NLP robustness measurements.
In this work we employ a PGD-based adversarial training as the method to increase the robustness and verifiability of our models, and we create character and word level perturbations as in [91] and sentence level perturbations with PolyJuice [132] and Vicuna [25].
0.2.3 Previous NLP Verification Approaches
Method Verification algorithm Verification characteristics Datasets NLP perturbations Embeddings Architectures (# of parameters) Robust training
Ours
SMT-based, Abstract interpretation-based,
BaB-based
Complete,
Precise,
Deterministic
RUARobot, Medical
[…]
Sentence: S-BERT, S-GPT
FFNN ($10^{4}$)
PGD-based
Jia et al. (2019) [54]
IBP-based
Incomplete,
Imprecise,
Deterministic
IMDB, SNLI
Word substitution
Word: GloVe
LSTM, CNN, BoW, Attention-based, ($10^{5}$)
IBP-based
Huang et al. (2019) [50]
IBP-based
Incomplete,
Imprecise,
Deterministic
AGNews, SST
Char and word substitution
Word: GloVe
CNN ($10^{5}$)
IBP-based
Welbl et al. (2020) [126]
IBP-based
Incomplete,
Imprecise,
Deterministic
SNLI, MNLI
Word deletion
Word: GloVe
Attention-based ($10^{5}$)
Data augmentation, random and beam search adversarial training, IBP-based
Zhang et al. (2021) [146]
IBP-based
Incomplete,
Imprecise,
Deterministic
IMDB, SST, SST2
Word perturbations
Word: not specified
LSTM ($10^{5}$)
IBP-based
Wang et al. (2023) [123]
IBP-based
Incomplete,
Imprecise,
Deterministic
IMDB, YELP, SST2
Word substitution
Word: GloVe
CNN ($10^{5}$)
IBP-based: Embedding Interval Bound Constraint (EIBC) triplet loss
Ko et al. (2019) [61]
Abstract interpretation-based
Incomplete,
Imprecise,
Deterministic
CogComp QC
[…]
Word: not specified
RNN, LSTM ($10^{5}$)
Shi et al. (2020) [107]
Abstract interpretation-based
Incomplete,
Imprecise,
Deterministic
YELP, SST
[…]
Word: not specified
Transformer ($10^{6}$)
Du et al. (2021) [37]
Abstract interpretation-based
Incomplete,
Imprecise,
Deterministic
Rotten Tomatoes Movie Review, Toxic Comment
[…]
Word: GloVe
RNN, LSTM ($10^{5}$)
IBP-based
Bonaert et al. (2021) [14]
Abstract interpretation-based
Incomplete,
Imprecise,
Deterministic
SST, YELP
[…]
Word: not specified
Transformer ($10^{6}$)
Ye et al. (2020) [134]
Randomised smoothing ($\alpha=0.01,n=5000$)
Incomplete,
Imprecise,
Probabilistic
IMDB, Amazon
Word substitution
Word: GloVe
Transformer ($10^{8}$)
Data
augmentation
Wang et al. (2021) [120]
Differential privacy-based
Incomplete,
Imprecise,
Probabilistic
IMDB, AGNews
Word substitution
Word: GloVe
LSTM ($10^{5}$)
Data
augmentation
Zhao et al. (2022) [148]
Randomised smoothing ($\alpha=0.001,n=30050$)
Incomplete,
Imprecise,
Probabilistic
AGNews, SST
Word substitution
Word: GloVe
Transformer ($10^{8}$)
Data
augmentation
and IBP-based
Zeng et al. (2023) [137]
Randomised smoothing ($\alpha=0.05,n=5000$)
Incomplete,
Imprecise,
Probabilistic
IMDB, YELP
Char and word substitution
Word: not specified
Transformer ($10^{8}$)
Data
augmentation
Ye et al. (2023) [135]
Randomised smoothing ($\alpha=0.001,n=9000$)
Incomplete,
Imprecise,
Probabilistic
IMDB, SST2, YELP, AGNews
Word substitution
Word: not specified
Transformer ($10^{8}$)
Data
augmentation
Zhang et al. (2023) [144]
Randomised smoothing ($\alpha=0.001,n=20000$)
Incomplete,
Imprecise,
Probabilistic
IMDB, Amazon, AGNews
Word perturbations
Word: GloVe
LSTM, Transformer ($10^{8}$)
Data
augmentation
Zhang et al. (2023) [147]
Randomised smoothing ($\alpha=0.05,n=5000$)
Incomplete,
Imprecise,
Probabilistic
SST2, AGNews
Word perturbations
Word: not specified
Transformer ($10^{9}$)
Table 1: Summary of the main features of the existing NLP verification approaches (first half) and randomised smoothing approaches (second half). In bold are SoA methods.
Although DNN verification studies have predominantly focused on computer vision, there is a growing body of research exploring the verification of NLP. This research can be categorised into three main approaches: IBP, abstract interpretation, and randomised smoothing. Table 1 shows a comparison of these approaches. To the best of our knowledge, this paper is the first one to use an SMT-based verifier for this purpose, and compare it with an abstract interpretation-based verifier on the same benchmarks.
Verification via Interval Bound Propagation.
The first technique successfully adapted from the computer vision domain for verifying NLP models was the IBP. In the NLP approaches, IBP is used for both training and verification. Its aim is to minimise the upper bound on the maximum difference between the classification boundary and the input perturbation region by augmenting the loss function. This facilitates the minimisation of the perturbation region in the last layer, ensuring it remains on one side of the classification boundary. As a result, the adversarial region becomes tighter and can be considered certified robust. Notably, Jia et al. [54] proposed certified robust models on word substitutions in text classification. The authors employed IBP to optimise the upper bound over perturbations, providing an upper bound over the discrete set of perturbations in the word vector space. Furthermore, Huang et al. [50] introduced a verification and verifiable training method for neural networks in NLP, proposing a tighter over-approximation in the form of a ‘simplex’ in the embedding space for input perturbations. To make the network verifiable, they defined the convex hull of all the original unperturbed inputs as a space of perturbations. By employing the IBP algorithm, they generated robustness bounds for each neural network layer. Later on, Welbl et al. [126] differentiated from the previous approaches by using IBP to address the under-sensitivity issue. They designed and formally verified the ‘under-sensitivity specification’ that a model should not become more confident as arbitrary subsets of input words are deleted. Recently, Zhang et al. [146] introduced Abstract Recursive Certification (ARC) to verify the robustness of LSTMs. ARC defines a set of programmatically perturbed string transformations to construct a perturbation space. By memorising the hidden states of strings in the perturbation space that share a common prefix, ARC can efficiently calculate an upper bound while avoiding redundant hidden state computations. Finally, Wang et al. [123] improved on the work of Jia et al. by introducing Embedding Interval Bound Constraint (EIBC). EIBC is a new loss that constraints the word embeddings in order to tighten the IBP bounds.
The strength of IBP-based methods is their efficiency and speed, while their main limitation is the bounds’ looseness, further accentuated if the neural network is deep.
Verification via Abstract Interpretation.
Another popular verification technique applied to various NLP models is based on abstract interpretation. Abstract interpretation was first developed by Cousot and Cousot [29] in 1977. It formalises the idea of abstraction of mathematical structures, in particular those involved in the specification of properties and proof methods of computer systems [28] and it has since been used in many applications [30]. Specifically, for DNN verification, this technique can model the behaviour of a network using an abstract domain that captures the possible range of values the network can output for a given input. This abstract domain can then be used to reason about the network’s behaviour under different conditions, such as when the network receives inputs that are adversarially perturbed. One notable contribution in this area is POPQORN [61], which is the first work that gives robustness guarantees for RNN-based networks. They handle the challenging non-linear activation functions of complicated RNN structures (like LSTMs and GRUs) by bounding them with linear functions. Later on, Du et al. improve on POPQORN by introducing Cert-RNN [37], a robust certification framework for RNNs that overcomes the limitations of POPQORN. The framework maintains inter-variable correlation and accelerates the non-linearities of RNNs for practical uses. Cert-RNN utilised Zonotopes [44] to encapsulate input perturbations and can verify the properties of the output Zonotopes to determine certifiable robustness. This results in improved precision and tighter bounds, leading to a significant speedup compared to POPQORN. Differently, Shi et al. [107] focus on transformers with self-attention layers. They developed a verification algorithm that can provide a lower bound to ensure the probability of the correct label is consistently higher than that of the incorrect labels. Analogously, Bonaert et al. [14] propose DeepT, a certification method for large transformers. It is specifically designed to verify the robustness of transformers against synonym replacement-based attacks. DeepT employes multi-norm Zonotopes to achieve larger robustness radii in the certification and can work with networks much larger than Shi et al.
[…]
Verification via Randomised Smoothing.
Randomised smoothing [27] is another technique for verifying the robustness of deep language models that has recently grown in popularity due to its scalability [134, 120, 148, 137, 135, 144, 147]. Its basic idea is to leverage randomness during inference to create a smoothed classifier that is more robust to small perturbations in the input. This technique can also be used to give certified guarantees against adversarial perturbations within a certain radius. Generally, randomized smoothing begins by training a regular neural network on a given dataset. During the inference phase, to classify a new sample, noise is randomly sampled from the predetermined distribution multiple times. These instances of noise are then injected into the input, resulting in noisy samples. Subsequently, the base classifier generates predictions for each of these noisy samples. The final prediction is determined by the class with the highest frequency of predictions, thereby shaping the smoothed classifier. To certify the robustness of the smoothed classifier against adversarial perturbations within a specific radius centered around the input, randomised smoothing calculates the likelihood of agreement between the base classifier and the smoothed classifier when noise is introduced to the input. If this likelihood exceeds a certain threshold, it indicates the certified robustness of the smoothed classifier within the radius around the input.
The main advantage of randomised smoothing-based methods is their scalability, indeed recent approaches are tested on larger transformer such as BERT and Alpaca. However, their main issue is that they are probabilistic approaches, meaning they give certifications up to a certain probability (e.g., 99.9%). In this work we focus on deterministic approaches, hence we only report these works in Table 1 for completeness without delving deeper into each paper here. All randomised smoothing-based approaches use data augmentation obtained by semantic perturbations.
0.2.4 Data Sets and Use Cases used in NLP Verification
Existing NLP verification data sets.
Table 2 summarises the main features and tasks of the datasets used in NLP verification. Despite their diverse origins and applications, the use of these datasets in the NLP verification literature converge on text classification, with a predominant focus on binary or multi-class categorisation. Furthermore, datasets can be sensitive to perturbations, i.e. perturbations can have non-trivial impact on label consistency. For example, Jia et al. [54] use IBP with the SNLI [15] [^3] A semantic inference dataset that labels whether one sentence entails, contradicts or is neutral to another sentence. dataset (see Tables 1 and 2) to show that word perturbations (e.g. ‘good’ to ‘best’) can change whether one sentence entails another. Some works such as [54] try to address this label consistency, while others do not.
Additionally, we find that the previous research on NLP verification does not utilise safety critical datasets (which strongly motivates the choice of datasets in alternative verification domains), with the exception of [37]. The papers do not provide detailed motivation as to why the dataset choices were made, however it could be due to the datasets being commonly used in NLP benchmarks (IMDB …). Instead, we focus on datasets motivated by safety critical applications.
Dataset Safety Critical Category Tasks Size Classes
IMDB [84]
$\times$
Sentiment analysis
Document-level and sentence-level classification
25,000
2
SST [112]
$\times$
Sentiment analysis
Sentiment classification, hierarchical sentiment classification, sentiment span detection
70,042
5
SST2 [112]
$\times$
Sentiment analysis
Sentiment classification
70,042
2
YELP [106]
$\times$
Sentiment analysis
Sentiment classification
570,771
2
Rotten Tomatoes Movie Review [96]
$\times$
Sentiment analysis
Sentiment classification
48,869
3/4
Amazon [86]
$\times$
Sentiment analysis
Sentiment classification, aspect-based sentiment analysis
34,686,770
5
SNLI [15]
$\times$
Semantic inference
Natural language inference, semantic similarity
570,152
3
MNLI [127]
$\times$
Semantic inference
Natural language inference, semantic similarity, generalisation
432,702
3
AGNews [143]
$\times$
Text analysis
Text classification, sentiment classification
127,600
4
CogComp QC [74]
$\times$
Text analysis
Question classification, semantic understanding
15,000
6/50
Toxic Comment [26]
Text analysis
Toxic comment classification, fine-grained toxicity analysis, bias analysis
18,560
6
Table 2: Summary of the main features of the datasets used in NLP verification.
Data Sets Proposed in This Paper
In this paper we propose to use two datasets that have not been used in NLP verification literature before. Both are driven by real-world use cases of safety-critical NLP applications, i.e. applications for which law enforcement and safety demand formal guarantees of “good" DNN behaviour.
Chatbot Disclosure Dataset.
First case concerns new legislation which states that a chatbot must not mislead people about its artificial identity [68, 63]. Given that the regulatory landscape surrounding NLP models (particularly LLMs and generative AI) is rapidly evolving, similar legislation could be widespread in the future – with recent calls for the US Congress to formalise such disclosure requirements [90]. The prohibition on deceptive conduct act may apply to the outputs generated by NLP systems if used commercially [4], and at minimum a system must guarantee a truthful response when asked about its agency [47, 1]. Furthermore, the burden of this should be placed on the designers of NLP systems, and not on the consumers.
Our first safety critical case is the R-U-A-Robot dataset [47], a written English dataset consisting of 6800 variations on queries relating to the intent of ‘Are you a robot?’, such as ‘I’m a man, what about you?’. The dataset was created via a context-free grammar template, crowd-sourcing and pre-existing data sources. It consists of 2,720 positive examples (where given the query, it is appropriate for the system to state its non-human identity), 3,400 negative/adversarial examples and 680 ‘ambiguous-if-clarify’ examples (where it is unclear whether the system is required to state its identity). The dataset was created to promote transparency which may be required when the user receives unsolicited phone calls from artificial systems. Given systems like Google Duplex [70], and the criticism it received for human-sounding outputs [77], it is also highly plausible for the user to be deceived regarding the outputs generated by other NLP-based systems [4]. Thus we choose this dataset to understand how to enforce such disclosure requirements. We collapse the positive and ambiguous examples into one label, following the principle of ‘better be safe than sorry’, i.e. prioritising a high recall system.
Medical Safety Dataset.
Another scenario one might consider is that inappropriate outputs of NLP systems have the potential to cause harm to human users [12]. For example, a system may give a user false impressions of its ‘expertise’ and generate harmful advice in response to medically related user queries [34]. In practice it may be desirable for the system to avoid answering such queries. Thus we choose the Medical safety dataset [2], a written English dataset consisting of 2,917 risk-graded medical and non medical queries (1,417 and 1,500 examples respectively). The dataset was constructed via collecting questions posted on reddit, such as r/AskDocs. The medical queries have been labelled by experts and crowd annotators for both relevance and levels of risk (i.e. non-serious, serious to critical) following established World Economic Forum (WEF) risk levels designated for chatbots in healthcare [130]. We merge the medical queries of different risk-levels into one class, given the high scarcity of the latter 2 labels to create an in-domain/out-of-domain classification task for medical queries. Additionally, we consider only the medical queries that were labelled as such by expert medical practitioners. Thus this dataset will facilitate discussion on how to guarantee a system recognises medical queries, in order to avoid generating medical output.
An additional benefit of these two datasets is that they are distinct semantically, i.e. the R-U-A-Robot dataset contains several semantically similar, but lexically different queries, while the medical safety dataset contains semantically diverse queries. For both datasets, we utilise the same data splits as given in the original papers, and refer to the final binary labels as positive and negative. The positive label in the R-U-A-Robot dataset implies a sample where it is appropriate to disclose non-human identity, while in the medical safety dataset it implies an in-domain medical query.
0.2.5 Our Work (NLP Verification Pipelines)
To show relation of our work to the body of already existing work, we distill an “NLP verification pipeline” that is common across many related papers. Figure 2 shows the pipeline diagrammatically. It proceeds in stages:
  1. 1.
    Given an NLP dataset, generate semantic perturbations on sentences that it contains. The semantic perturbations can be of different kinds: character, word or sentence level. IBP and randomised smoothing use word and character perturbations, abstract interpretation papers usually do not use any semantic perturbations. Our method allows to use all existing semantic perturbations, in particular, we implement character and word level perturbations as in [91], sentence level perturbations with PolyJuice [132] and Vicuna.
  2. 2.
    Embed the semantic perturbations into continuous spaces. The cited papers use the word embeddings GloVe [98], we use the sentence embeddings S-BERT and S-GPT.
  3. 3.
[…]
    Given a notion of geometric/semantic subspaces, one can use it in two different ways:
    • 4.
[…]
    • 5.
      $\ldots$ verify the classifier’s behaviour within those subspaces. IBP papers use IBP algorithms which are incomplete and imprecise, abstract interpretation gives incomplete and imprecise methods, and we use SMT-based tool Marabou (complete and precise) and abstract-interpretation tool ERAN (incomplete and imprecise).
Table 1 summarises differences and similarities of the above NLP verification approaches against ours. To the best of our knowledge, we are the first to use complete and precise methods in NLP verification. This paper is the first to employ SMT-based verifiers and to show how they out-perform abstract interpretation-based verification, which produce tighter bounds than IBP-based methods.
Furthermore, our study is the first to demonstrate that the construction of semantic subspaces can happen independently of the choice of the training and verification algorithms. Likewise, although training and verification build upon the defined (semantic) subpaces, the actual choice of the training and verification algorithms can be made independently of the method used to define the semantic subspaces. This separation, and the general modularity of our approach, facilitates a comprehensive examination and comparison of the two key components involved in any NLP verification process:
  • effects of the verifiability-generalisability trade-off for verification with geometric and semantic subspaces;
  • relation between the volume/shape of semantic subpaces and verifiability of neural networks obtained via semantic training with these subpaces.
These two aspects have not been considered in the literature before.
0.3 The Parametric NLP Verification Pipeline
Figure 2: Visualisation of the NLP verification pipeline followed in our approach.
This section presents a parametric NLP verification pipeline, shown in Figure 2 diagrammatically. We call it “parametric” because each component within the pipeline operates independently of the others and can be taken as a parameter when studying other components. The parametric nature of the pipeline allows for the seamless integration of state-of-the-art methods at every stage, and for more sophisticated experiments with those methods. Note that the outlined pipeline can be seen as a filter which can be applied on top of an NLP system or LLM (such as S-BERT and S-GPT) to certify intended DNN behavior for safety-critical input queries. The following section provides a detailed exposition of the methodological choices made at each step of the pipeline.
0.3.1 Semantic Perturbations
As discussed in Section 0.2.5, we require semantic perturbations for creating semantic subspaces. To do so, we consider three kinds of perturbations – i.e. character, word and sentence level. This systematically accounts for different variations of the samples.
Character and word level perturbations
are created via a rule-based method proposed in [91], to simulate different kinds of noise one could expect from spelling mistakes, typos etc. These perturbations are non-adversarial and can be generated automatically. The authors found that NLP models are sensitive to such small errors, while in practice this should not be the case. Character level perturbations types include randomly inserting, deleting, replacing, swapping or repeating a character of the data sample. At the character level, we do not apply letter case changing, given it does not change the sentence-level representation of the sample, nor do we apply perturbations to commonly misspelled words, given only a small overlap of those words occur in our datasets. Perturbations types at the word level include randomly repeating or deleting a word, changing the ordering of the words, the verb tense, singular verbs to plural verbs or adding negation to the data sample. At the word level, we omit replacement with synonyms, as this is accounted for via sentence rephrasing. Negation is not done on the medical safety dataset, as it creates label ambiguities (e.g. ‘pain when straightening knee’ $\rightarrow$ ‘no pain when straightening knee’), as well as singular plural tense and verb tense, given human annotators would experience difficulties with this task (e.g. rephrase the following in plural/ with changed tense – ‘peritonsillar abscess drainage aftercare.. please help’).
Further examples of character and word rule-based perturbations can be found in Tables 3 and 4.
Method Description Original sentence Altered sentence
Insertion
A character is randomly selected and inserted in a random position.
Are you a robot?
Are yovu a robot?
Deletion
A character is randomly selected and deleted.
Are you a robot?
Are you a robt?
Replacement
A character is randomly selected and replaced by an adjacent character on the keyboard.
Are you a robot?
Are you a ronot?
Swapping
A character is randomly selected and swapped with the adjacent right or left character in the word.
Are you a robot?
Are you a rboot?
Repetition
A character in a random position is selected and duplicated.
Are you a robot?
Arre you a robot?
Table 3: Character-level perturbations: their types and examples of how each type acts on a given sentence from the R-U-A-Robot dataset [47]. Perturbations are selected from random words that have 3 or more characters, first and last characters of a word are never perturbed.
Method Description Original sentence Altered sentence
Deletion
Randomly selects a word and removes it.
Can u tell me if you are a chatbot?
Can u tell if you are a chatbot?
Repetition
Randomly selects a word and duplicates it.
Can u tell me if you are a chatbot?
Can can u tell me if you are a chatbot?
Negation
Identifies verbs then flips them (negative/positive).
Can u tell me if you are a chatbot?
Can u tell me if you are not a chatbot?
Singular/ plural verbs
Changes verbs to singular form, and conversely.
Can u tell me if you are a chatbot?
Can u tell me if you is a chatbot?
Word order
Randomly selects consecutive words and changes the order in which they appear.
Can u tell me if you are a chatbot?
Can u tell me if you are chatbot a?
Verb tense
Converts present simple or continuous verbs to their corresponding past simple or continuous form.
Can u tell me if you are a chatbot?
Can u tell me if you were a chatbot?
Table 4: Word-level perturbations: their types and examples of how each type acts on a given sentence from the R-U-A-Robot dataset [47].
Sentence level perturbations.
We experiment with two types of sentence level perturbations, particularly due to the complicated nature of the medical queries (e.g. it is non-trivial to rephrase queries such as this – ‘peritonsillar abscess drainage aftercare.. please help’). We do so by either using Polyjuice [132] or vicuna-13b [^4] Using the following API: https://replicate.com/replicate/vicuna-13b/api.. Polyjuice is a general-purpose counterfactual generator that allows for control over perturbation types and locations, trained by fine-tuning GPT-2 on multiple datasets of paired sentences. Vicuna is a state-of-the-art open source chatbot trained by fine-tuning LLaMA [115] on user-shared conversations collected from ShareGPT  [^5] https://sharegpt.com/. For Vicuna, we use the following prompt to generate variations on our data samples ‘Rephrase this sentence 5 times: “[Example]”.’ For example, from the sentence “How long will I be contagious?", we can obtain “How many years will I be contagious?" or “Will I be contagious for long?" and so on.
We will use notation $\mathcal{P}$ to refer to a perturbation algorithm abstractly.
0.3.2 NLP Embeddings
Next component of the pipeline are embeddings. Embeddings play a crucial role in NLP verification as they map textual data into continuous vector spaces, in a way that should capture semantic relationships and contextual information.
Given an NLP dataset $\mathcal{Y}$ as a set of sentences $s_{1},\ldots,s_{q}$ written in natural language, an embedding $\mathbb{E}$ is a function that maps a sentence to a vector in ${}^{m}$. The vector space ${}^{m}$ is called the embedding space. Ideally, $\mathbb{E}$ should reflect the semantic similarities between sentences in $\mathcal{Y}$, i.e. the more semantically similar two sentences $s_{i}$ and $s_{j}$ are, the closer the distance between $\mathbb{E}(s_{i})$ and $\mathbb{E}(s_{j})$ should be in ${}^{m}$. Of course, defining semantic similarity in precise terms may not be tractable (the number of unseen sentences maybe infinite, the similarity maybe understood subjectively and/or depending on the context). This is why, the state-of-the-art NLP relies on machine learning methods to capture the notion of semantic similarity approximately.
Currently, the most common approach to obtain an embedding function $\mathbb{E}$ is by training transformers [32, 101]. Transformers are a type of DNNs that can be trained to map sequential data into real vector spaces and are capable of handling variable-length input sequences. They can also be used for other tasks, such as classification or sentence generation, but in those cases, too, training happens at the level of embedding spaces. In this work, a transformer is trained as a function $\mathbb{E}:\mathcal{Y}\rightarrow^{m}$ for some given $m$. The key feature of the transformer is the “self-attention mechanism", which allows the network to weigh the importance of different elements in the input sequence when making predictions, rather than relying solely on the order of elements in the sequence. This makes them good at learning to associate semantically similar words or sentences. In this work we initially use Sentence-BERT [101] and later add Sentence-GPT [92] to embed sentences. Unfortunately, the relation between the embedding space and the NLP dataset is not bijective: i.e. each sentence is mapped into the embedding space, but not every point in the embedding space has a corresponding sentence. This problem is well-known in NLP literature [65] and, as shown in this paper, is one of the reasons why verification of NLP is tricky. Given an NLP dataset $\mathcal{Y}$ that should be classified into $n$ classes, the standard approach is to construct a function $N:^{m}\rightarrow^{n}$ that maps the embedded inputs to the classes. In order to do that, a domain specific classifier $N$ is trained on the embeddings $\mathbb{E}(\mathcal{Y})$ and the final system will then be the composition of the two subsystems, i.e. $N\circ\mathbb{E}$.
0.3.3 Geometric Analysis of Embedding Spaces
We now formally define geometric and semantic subpaces of the embedding space. Our goal is to define subpaces on the embedding space ${}^{m}$ by using an effective algorithmic procedure.
We start with an observation that, given an NLP dataset $\mathcal{Y}$ that contains (a finite number of sentences) $s_{1},\ldots,s_{q}$, and an embedding function $\mathbb{E}:\mathcal{Y}\rightarrow^{m}$, we can define an embedding matrix $\mathcal{X}\in^{q\times m}$, where each row $j$ is given by $\mathbb{E}(s_{j})$. Treating embedded sentences as matrices, rather than as points in the real vector space, makes many computations easier. This motivates the following definition of a hyper-rectangle for $\mathcal{X}$.
Definition 1 (Hyper-rectangle for an Embedding Matrix).
Given an embedding matrix $\mathcal{X}\in^{q\times m}$, the $m$-dimensional hyper-rectangle for $\mathcal{X}$ is defined as:
$\mathbb{H}_{\mathcal{X}}\coloneqq\{x\in^{m}:\forall j\in[1,...,m].\>x^{j}\in[\min\nolimits_{i}(\mathcal{X}^{ij}),\max\nolimits_{i}(\mathcal{X}^{ij})]\}.$
where $\min\nolimits_{i}(\mathcal{X}^{ij})$ stands for the minimum value of the $j$-th element $x^{j}$ across all rows $i\in\{0,\ldots,q\}$ in $\mathcal{X}$; similarly for $\max\nolimits_{i}(\mathcal{X}^{ij})$.
Definition 2 (Subspace of the Embedding Space).
Given a NLP dataset $\mathcal{Y}$, an embedding function $\mathbb{E}:\mathcal{Y}\rightarrow^{m}$, and $\mathcal{Y}^{\prime}\subseteq\mathcal{Y}$ (with $\mathcal{Y}^{\prime}=\{s_{1},\ldots,s_{q}\}$), we say a hyper-rectangle $\mathbb{H}_{\mathcal{X}^{\prime}}$ is a subspace of the embedding space ${}^{m}$ around $\mathcal{Y}^{\prime}$ if the matrix $\mathcal{X}^{\prime}\in^{q\times m}$ satisfies the following condition: each $j$th row is given by $\mathbb{E}(s_{j})$.
[…]
$\mathbb{B}(\hat{x},\epsilon)\coloneqq\{x\in^{m}:||\hat{x}-x||\leq\epsilon\}.$
[…]
Of course, as we have already discussed in the introduction and Figure 1, hyper-rectangles are not very precise, geometrically. A more precise shape would be a convex hull around $q$ given points in the embedding space. Indeed literature has some definitions of convex hulls [7, 105, 55]. However, none of them is suitable as they are computationally too expensive due to the time complexity of $O(q^{m/2})$ where $q$ is the number of inputs and $m$ is the number of dimensions [7]. Approaches that use under-approximations to speed up the algorithms [105, 55] do not work well in NLP scenarios, as under-approximated subspaces are so small that they contain near zero sentence embeddings.
A computationally efficient way of making hype-rectangles more precise is to rotate them to align to the position of the given points in the embedding space. This motivates us to introduce the Eigenspace rotation.
Eigenspace Rotation.
To construct the tightest possible hyper-rectangle, we define a specific method of eigenspace rotation. As shown in Figure 1 (C and D), our approach is to calculate a rotation matrix $A$ such that the rotated matrix $\mathcal{X}_{\text{rot}}=\mathcal{X}A$ is better aligned with the axes than $\mathcal{X}$, and therefore $\mathbb{H}_{\mathcal{X}_{\text{rot}}}$ has a smaller volume. By a slight abuse of terminology, we will refer to $\mathbb{H}_{\mathcal{X}_{\text{rot}}}$ as the rotated hyper-rectangle, even though strictly speaking, we are rotating the data, not the hyper-rectangle itself. In order to calculate the rotation matrix $A$, we use singular value decomposition [60]. The singular value decomposition of $\mathcal{X}$ is defined as $\mathcal{X}=U\Sigma V^{*}$, where $U$ is a matrix of left-singular vectors, $\Sigma$ is a matrix of singular values and $V^{*}$ is a matrix of right-singular vectors and $\cdot^{*}$ denotes the conjugate transpose. Intuitively, the right-singular vectors $V^{*}$ describe the directions in which $\mathcal{X}$ exhibits the most variance. The main idea behind the definition of rotation is to align these directions of maximum variance with the standard canonical basis vectors. Formally, using $V^{*}$, we can compute the rotation (or change-of-basis) matrix $A$ that rotates the right-singular vectors onto the canonical standard basis vectors $I$, where $I$ is the identity matrix. To do this, we observe that $V^{*}A=I\rightarrow V^{*}=IA^{-1}\rightarrow V^{-1}=A^{-1}\rightarrow V=A$. We thus obtain $\mathcal{X}_{rot}=\mathcal{X}A$ as desired.
All hyper-rectangles constructed and used in this paper are rotated.
Geometric and Semantic Subspaces.
[…]
We will finish this section with defining semantic subspaces formally. Given a data set $\mathcal{Y}$, and some defined algorithm for sentence perturbation $\mathcal{P}$, we can form $\mathcal{Y}^{\prime}=\{s^{*}_{1},\ldots,s^{*}_{q}\}$, such that each $s^{*}_{i}=\mathcal{P}(s_{i})$. Intuitively, we want to construct a subspace for $\mathcal{Y}^{\prime}$ as described in Definition 2. However, in the later sections, we will need to refer to different kinds of perturbations (e.g. character, word or sentence level as explained in Section 0.3.1) or different types of perturbations (e.g. insertion, deletion, replacement, etc.) as illustrated in Tables 3 and 4). This motivates the following definitions. Notation $\mathcal{P}_{k}^{t}(s,a)$ will denote a perturbation of kind $k$ and type $t$, applied in the sentence $s$, in a position $a$. As all algorithms we use choose $a$ randomly, we will omit $a$ and write just $\mathcal{P}_{k}^{t}(s)$. When we say that a sentence $s$ has $b$ semantic perturbations of kind $k$, we refer to a set $\mathcal{A}^{k}_{s}=\{\mathcal{P}_{k}^{t_{1}}(s),\mathcal{P}_{k}^{t_{2}}(s),\ldots,\mathcal{P}_{k}^{t_{b}}(s)\}$, for some given choice of $b$ perturbations of types $t_{1},\ldots,t_{b}$. When the choice is unimportant or clear from the context, we may omit mentioning the chosen types of perturbation.
Definition 3 (Semantic Subspace for a Sentence).
Suppose we are given an NLP dataset $\mathcal{Y}=\{s_{1},\ldots,s_{q}\}$, and an embedding function $\mathbb{E}:\mathcal{Y}\rightarrow^{m}$.
Given a sentence $s_{i}$ and a corresponding set $\mathcal{A}^{k}_{s_{i}}$ of its semantic perturbations (of kind $k$), a semantic subspace for $s_{i}$ is a subspace around $\mathcal{Y}^{\prime}=\mathcal{A}^{k}_{s_{i}}$ constructed according to Definition 2.
It will be useful in the later sections to use notation $\mathbb{H}_{\mathcal{X}_{s_{i}}}$ to refer to the hyper-rectangle used in constructing the semantic subspace for $s_{i}$. We will refer to a set of such hyper-rectangles $\mathbb{H}^{*}_{k}=\bigcup_{i=1}^{q}\mathbb{H}^{*}_{\mathcal{X}_{s_{i}}}$, where $s_{i}\in\{s_{1},\ldots,s_{q}\}$ and $k$ is a chosen perturbation kind.
If $\mathbb{H}^{*}_{k}$ is based on perturbations of only one type, we will sometimes use that type as an index to $\mathbb{H}^{*}_{\_}$.
Example 2 (Construction of Semantic Subspaces).
To illustrate this construction, let us consider the sentence $s_{i}=$“Can u tell me if you are a chatbot?”. This sentence is one of $3400$ original sentences of the positive class in the dataset. From this single sentence, we can create six sentences (word-level perturbations), see Table 4. Once these seven sentences are embedded into the vector space, they form the hyper-rectangle $\mathbb{H}_{\mathcal{X}_{s_{i}}}$. By repeating this construction for the remaining $3399$ sentences, we obtain the set of hyper-rectangles $\mathbb{H}^{*}_{word}$ for the dataset.
0.3.4 Training
As outlined in Section 0.2.2, robust training is essential for bolstering the robustness of DNNs; without it, their verifiability would be significantly diminished. This study employs two robust training methods, namely data augmentation and a custom PGD adversarial training, with the goal of discerning the factors contributing to the success of robust training and compare the effectiveness of these methods.
Data Augmentation.
In this training method, we statically generate semantic perturbations at the character, word, and sentence levels before training, which are then added to the dataset. The network is subsequently trained on this augmented dataset using the standard stochastic gradient descent algorithm.
Adversarial Training.
In this training method, we modify the traditional Projected Gradient Descent (PGD) algorithm [85], defined as follows. Given a loss function $\mathcal{L}$, a step size $\gamma\in$ and a starting point $\hat{x}_{0}$ then the output of the PGD algorithm $x(l)$ after $l$ iterations is defined as:
$x(0)$ $=\hat{x}_{0}$
$x(t+1)$ $=\text{proj}_{\mathbb{H}}\Big{[}x(t)+\gamma\cdot\text{sign}(\nabla_{x(t)}\mathcal{L}(x(t),y))\Big{]}$
[…]
The resulting customised PGD training seeks to identify the worst perturbations within the custom-defined subspace, and trains the given neural network to classify those perturbations correctly, in order to make the network robust to adversarial inputs in the chosen subspace.
0.3.5 Choice of Verification Algorithm
As stated earlier, our approach in this study involves the utilization of cutting-edge tools for DNN verification. Initially, we employ ERAN [110], a state-of-the-art abstract interpretation-based method. This choice is made over IBP due to its ability to yield tighter bounds. Subsequently, we conduct comparisons and integrate Marabou [131], a state-of-the-art complete and precise SMT-based verifier. This enables us to attain the highest verification percentage, maximizing the tightness of the bounds.
We will use notation $\mathcal{V}$ to refer to a verifier abstractly.
0.4 Characterisation of Verifiable Subspaces
In this Section, we provide key results in support of Contribution 1 formulated in the introduction:
  • We start with introducing the metric of generalisability of (verified) subspaces, and introducing the problem of the generalisability-verifiability trade-off.
  • We show that the use of semantic subspaces helps to find a better balance between generalisability and verifiability, as compared to the use of geometric subspaces.
  • Finally, we show that adversarial training based on semantic subspaces results in DNNs that are both more verifiable and more generalisable than those obtained with other forms of robust training.
0.4.1 Generalisability-Verifiability Trade-off
This subsection defines the new metric of generalisability and shows its effect on (naively defined) geometric subspaces.
Generalisability of Verified Subspaces: Formal Definition
Let us start with recalling the existing standard metrics used in DNN verification. Recall that we are given an NLP dataset $\mathcal{Y}=\{s_{1},\ldots,s_{q}\}$, moreover we assume that each $s_{i}$ is assigned a correct class from $C=\{c_{1},\ldots,c_{n}\}$. We restrict to the case of binary classification in this paper for simplicity, so we will assume $C=\{c_{1},c_{2}\}$. Furthermore, we are given an embedding function $\mathbb{E}:\mathcal{Y}\rightarrow^{m}$, and a DNN $N:^{m}\rightarrow^{n}$. Usually $n$ corresponds to the number of classes, and thus in case of binary classification, we have $N:^{m}\rightarrow^{2}$. Classification, or assignment of a vector $v$ in ${}^{m}$ to a class, depends on the highest value in the vector $N(v)$. The most popular metric is accuracy of $N$, which is measured as a percentage of $q$ vectors in ${}^{m}$ that are assigned to a correct class by $N$. Note that this metric checks a finite number of points in ${}^{m}$ given by the data set.
The most popular metric in DNN verification is verifiability. Recall that we can define subspaces $\{\mathcal{S}_{1},\ldots,\mathcal{S}_{s}\}$ of ${}^{m}$; in such a way that each of them is associated with a class (to which all points in that subspace should belong). A DNN verifier $\mathcal{V}$ takes an $\mathcal{S}_{i}$, its designated class $c_{i}$ and $N$ as an input, and returns as an output an answer whether all points in $\mathcal{S}_{i}$ are guaranteed to be assigned to the class $c_{i}$ by $N$. Verifiability is a percentage of subspaces in $\{\mathcal{S}_{1},\ldots,\mathcal{S}_{s}\}$ that are successfully verified. All DNN verification papers report this measure. Note that each $\mathcal{S}_{i}$ contains an infinite number of points.
We are introducing a third metric – generalisability of (verified) subspaces. Suppose we have a subspace $\mathcal{S}_{i}$ that verifiably consists only of vectors that are assigned to a class $c_{i}$ by $N$. Because of the embedding gap, we cannot know in advance how many valid sentences in $\mathcal{Y}$ (or outside of $\mathcal{Y}$ !) will be mapped into $\mathcal{S}_{i}$ by $\mathbb{E}$. Checking the former is rather easy, but may not reflect the idea of generalisability to unseen similar sentences. Checking the latter is hard. We propose the following effective heuristic. By the constructions of the previous section, any $\mathcal{S}_{i}$ necessarily contains an embedding of at least one element $s_{i}$ from $\mathcal{Y}$. Choosing any perturbation algorithm $\mathcal{P}$ (of kind $k$), we can form a set $\mathcal{A}^{k}_{s_{i}}$, as described in Section 0.3.3. Note that $\mathcal{P}$ (and $k$) can be given by a collection of different perturbation algorithms and their kinds. The key assumption is that $\mathcal{A}^{k}_{s_{i}}$ contains valid sentences semantically similar to $s_{i}$ and belonging to the same class.
Consider a set $\mathcal{V}_{i}$ of vectors in ${}^{m}$ such that each vector $v_{j}\in\mathcal{V}_{i}$ is defined as $v_{j}=\mathbb{E}(s_{j})$, with $s_{j}\in\mathcal{A}^{k}_{s_{i}}$. This is the set of embeddings of sentences contained in $\mathcal{A}^{k}_{s_{i}}$. Percentage of elements of $\mathcal{V}_{i}$ contained in $\mathcal{S}_{i}$ gives us the generalisability of $\mathcal{S}_{i}$. Note that each $\mathcal{V}_{i}$ is finite, because each $\mathcal{A}^{k}_{s_{i}}$ is finite.
Note that, unlike accuracy and verifiability, the generalisability metric does not explicitly depend on any DNN. However, in this paper we only study generalisability of verifiable subspaces, and thus the existence of a verified DNN $N$ will be assumed.
Because verifiability is reported as a percentage of $\{\mathcal{S}_{1},\ldots,\mathcal{S}_{s}\}$, it will be convenient to report generalisability over the set of subspaces $\mathcal{S}=\bigcup\mathcal{S}_{i}$. For this, we take $\mathcal{A}=\bigcup\mathcal{A}^{k}_{s_{i}}$ and the corresponding $\mathcal{V}=\bigcup\mathcal{V}_{i}$ for all $s_{i}\in\mathcal{Y}$ and compute generalisability of $\mathcal{S}$ as a percentage of elements of $\mathcal{V}$ contained in all subspaces of $\mathcal{S}$ [^6] Note that this calculation allows an element of $\mathcal{A}^{k}_{s_{i}}$ to count towards generalisability if it belongs to $\mathcal{S}_{j}$, regardless of whether $i=j$ or $i\neq j$. Different approaches might involve restraining the validity to $i=j$, or computing generalisability of each $\mathcal{S}_{j}$ individually and taking an average. The current choice favours a global view by calculating how generalisable the whole collection $\mathcal{S}$ is. . When it is important to emphasise the kind of perturbation, we will use the notation $\mathcal{A}_{k}$ for $\bigcup\mathcal{A}^{k}_{s_{i}}$.
Base Line Experiments: Understanding the Properties of Embedding Spaces
Figure 3: An example of hyper-rectangle drawn around all points of the same class (left), shrunk hyper-rectangle $\mathbb{H}_{sh}$ that is obtained by excluding all points from the opposite class (center) and clustered hyper-rectangles (right) in 2-dimensions. The red dots represent sentences in the embedding space of one class, while the blue dots are embedded sentences that do not belong to that class.
The methodology defined thus far has given basic intuitions about the relative nature of the NLP verification pipeline. Bearing this in mind, it is important to start our analysis with the general study of the embedding (sub)spaces and suitable base line settings.
Benchmark data sets will be abbreviated as “RUAR" and “Medical". For a benchmark network $N:^{m}\rightarrow^{2}$, we train a medium-sized fully-connected DNN, using stochastic gradient descent and cross-entropy loss. The main requirement for a benchmark network is its sufficient accuracy, see Table 5.
Model Adversarial training Train Accuracy RUAR Test Accuracy RUAR Train Accuracy Medical Test Accuracy Medical
$N_{base}$
No
0.9387 ± 0.0014
0.9357 ± 0.0018
0.9632 ± 0.0005
0.9449 ± 0.0026
Table 5: Accuracy of the baseline DNN on the R-U-A-Robot and the Medical datasets.
For the choice of subspaces, which is our main interest in this paper, two extreme (and trivially defined) benchmarks, are the following geometric subspaces:
  • a subspace drawn around all embedded sentences of the same class in $\mathcal{Y}$. This is the largest subspace one may wish to verify but we should assume that verifiability of such a subspace would be near $0\%$. It is illustrated in the first graph of Figure 3.
[…]
We need to first understand exact geometric properties (e.g. volume, $\epsilon$ values) and exact verifiability figures for these two extremes. Let us start with understanding the volumes.
Volumes of Subspaces.
To start with the largest hyper-rectangle, we need to ensure that it is shrunk to exclude all samples from $\mathcal{Y}$ from the wrong class. We define the shrinking method that involves reducing the size of the hyper-rectangle to eliminate each input within the hyper-rectangle that does not pertain to the chosen class. Second graph of Figure 3 gives a visual intuition of how this is done.
Formally, suppose we have a hyper-rectangle $\mathbb{H}_{\mathcal{X}}$ and a vector $\hat{x}\in\mathbb{H}_{\mathcal{X}}$ that violates the class requirement. Recall that by definition:
$\mathbb{H}_{\mathcal{X}}\coloneqq\{x\in^{m}:\forall j\in[1\;,\;...\;,\;m]\;.\;\>x^{j}\in[\min\nolimits_{i}(\mathcal{X}^{ij}),\;\;\max\nolimits_{i}(\mathcal{X}^{ij})]\}.$
The algorithm starts with choosing a dimension $j\in[1,...,m]$, and computing whether $|\hat{x}^{j}-\min\nolimits_{i}(\mathcal{X}^{ij})|\leq|\hat{x}^{j}-\max\nolimits_{i}(\mathcal{X}^{ij})|$. If the inequality holds, take $d_{\min\nolimits}=|\hat{x}^{j}-\min\nolimits_{j}(\mathcal{X}^{ij})|$ and compute a new minimum for the dimension $j$: $\min\nolimits^{new}_{i}(\mathcal{X}^{ij})=\min\nolimits_{i}(\mathcal{X}^{ij})+d_{min}+\delta$, where $\delta$ is a small positive number (we use $e^{-100}$). We then recompute $\mathbb{H}_{sh}=\mathbb{H}_{\mathcal{X}}$ with $\min\nolimits_{i}(\mathcal{X}^{ij})=\min\nolimits^{new}_{i}(\mathcal{X}^{ij})$ for the chosen dimension $j$.
If the inequality does not hold, take $d_{\max\nolimits}=|\hat{x}^{j}-\max\nolimits_{i}(\mathcal{X}^{ij})|$ and compute a new maximum for the dimension $j$: $\max\nolimits^{new}_{i}(\mathcal{X}^{ij})=\max\nolimits_{i}(\mathcal{X}^{ij})-d_{max}-\delta$, with $\delta$ defined as above. We then recompute $\mathbb{H}_{sh}=\mathbb{H}_{\mathcal{X}}$ with $\max\nolimits_{i}(\mathcal{X}^{ij})=\max\nolimits^{new}_{i}(\mathcal{X}^{ij})$ for the chosen dimension $j$.
Note that modifying just one dimension of $\hat{x}$ like this guarantees that it is not in $\mathbb{H}_{sh}$. The algorithm we apply when computing $\mathbb{H}_{sh}$ involves a heuristic for choosing the optimal dimension along which we modify the hyper-rectangle. The idea is to make the most conservative choice, that excludes as few other embedded sentences as possible [^7] Note that this algorithm shrinks exactly one dimension by a minimal amount to exclude the unwanted embedded sentence. This choice keeps the algorithm fast while guaranteeing the subspace to retain the highest number of wanted inputs. However, it is not necessarily the best choice for verification: there might be cases where perturbations of the unwanted input are left inside after shrinking and, if the network classifies them correctly, the subspace can never be verified. For large subspaces, our algorithm might render verification unachievable and more clever algorithms should be explored and discussed.. It amounts to computing how many inputs from the chosen class remain inside of the hyper-rectangle after applying the shrinking algorithm to one $j\in[1,...,m]$ at a time, separately. We choose the dimension $j$ for which the hyper-rectangle retains the most inputs of the correct class. The algorithm is run iteratively for every element of $\mathbb{H}_{sh}$ that has the wrong class, until $\mathbb{H}_{sh}$ contains only embeddings of sentences of one class.
[…]
To give us a sense of progression from the largest hyper-rectangle to the smallest, we also cluster $\mathbb{H}_{sh}$, see the last graph of Figure 3. For the clustering method, we employ the k-means algorithm to generate multiple ensembles of hyper-rectangles around each cluster. This approach guarantees that the hyper-rectangles contain solely points from the specific class, while also reducing their volume. We obtain a set of $50$, $100$, $200$, $250$ clusters denoted as $\mathbb{H}^{*}_{50}$$\mathbb{H}^{*}_{250}$.
Table 6 shows average volumes of the subspaces that we obtain, for two data sets. Notice consistent reduction of volume, from $\mathbb{H}_{sh}$ to $\mathbb{H}^{*}_{50}$ $\mathbb{H}^{*}_{250}$ and ultimately to $\mathbb{H}^{*}_{\epsilon=0.005}$. There are several orders of magnitude between the largest and the smallest subspace.
Hyper-rectangles Construction method Avg. volume of $\mathbb{H}$ RUAR Set cardinality of $\mathbb{H}^{*}$ RUAR Avg. volume of $\mathbb{H}$ Medical Set cardinality of $\mathbb{H}^{*}$ Medical
$\mathbb{H}_{sh}$
Hyper-rectangle shrunk to exclude all negative examples
7.55e-11
1
2.6e-09
1
$\mathbb{H}^{*}_{50}$
Set of hyper-rectangles on 50 clusters
1.02e-16
50
6.56e-15
50
$\mathbb{H}^{*}_{100}$
Set of hyper-rectangles on 100 clusters
6.23e-18
100
3.25e-17
100
$\mathbb{H}^{*}_{200}$
Set of hyper-rectangles on 200 clusters
3.31e-20
200
4.67e-19
200
$\mathbb{H}^{*}_{250}$
Set of hyper-rectangles on 250 clusters
6.42e-22
250
2.42e-20
250
$\mathbb{H}^{*}_{\epsilon=0.005}$
Set of $\epsilon$-cubes around all positive dataset sentences, $\epsilon=0.005$
1.00e-60
3400
1.00e-60
989
Table 6: Sets of geomtetric subspaces used in the experiments, their cardinality and average volumes of hyper-rectangles. All shapes are eigenspace rotated for better precision.
Verifiability Range for Subspaces.
For each set of hyper-rectangles and the given network, we pass them to the verifier (ERAN) and measure verifiability. Table 7 shows that the shrunk hyper-rectangle $\mathbb{H}_{sh}$ achieves zero verifiability, and the various clustered hyper-rectangles ($\mathbb{H}^{*}_{50}$, $\mathbb{H}^{*}_{100}$ $\mathbb{H}^{*}_{200}$, $\mathbb{H}^{*}_{250}$) achieve at most negligible verifiability. We call this effect low verifiability of the subspaces. In contrast, the baseline $\mathbb{H}^{*}_{\epsilon=0.005}$ achieves up to $99.60\%$ verifiability. This suggests that $\epsilon=0.005$ is a good benchmark for a different extreme. Table 6 can give us an intuition of why $\mathbb{H}^{*}_{\epsilon=0.005}$ has notably higher verifiability than the other hyper-rectangles: the volume of $\mathbb{H}^{*}_{\epsilon=0.005}$ is several orders of magnitude smaller.
Dataset Model $\boldsymbol{\mathbb{H}_{sh}}$ $\boldsymbol{\mathbb{H}^{*}_{50}}$ $\boldsymbol{\mathbb{H}^{*}_{100}}$ $\boldsymbol{\mathbb{H}^{*}_{200}}$ $\boldsymbol{\mathbb{H}^{*}_{250}}$ $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.005}}$
RUAR $N_{base}$ 0.00 0.00 1.33 0.52 0.41 0.00 88.67
Medical $N_{base}$ 0.00 0.00 0.00 2.10 4.08 5.00 97.86
Table 7: Verifiability of the baseline DNN on the R-U-A-Robot and the Medical datasets, for a selection of geometric subspaces; for ERAN verifier.
Verifiability-Generalisability Trade-off for Geometric Subspaces
Tables 6 and 7 suggest that smaller subspaces are more verifiable. One may also conjecture that they are less generalisable (as they will contain fewer embedded sentences). We confirm this via experiments; we are particularly interested in understanding how quickly generalisability deteriorates as verifiability increases.
To test generalisability, for each data set, we algorithmically generate a new data set $\mathcal{A}$ containing its semantic perturbations, using the method described in Section 0.3.1. It will be convenient to refer to parts of that set that correspond to perturbations of sentences belonging to positive and negative classes, respectively. In such cases, we will be using the notation $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$. Note that in this section, we will work only with $\mathcal{A}^{pos}$. This choice is motivated by the nature of the chosen data sets: both Medical and R-U-A-Robot sentences split into:
  • a positive class, that contains sentences with one intended semantic meaning (they are medical queries, or they are questions about robot identity); and
  • a negative class that represents “all other sentences". These “other sentences" are not grouped by any specific semantic meaning and therefore do not form one coherent semantic category.
However Section 0.5 will make use of $\mathcal{A}^{neg}$ in the context of falsifiability of verified subspaces.
[…]
For generalisability, Table 8 shows that the most verifiable subspace $\mathbb{H}^{*}_{\epsilon=0.005}$ is the least generalisable. This means $\mathbb{H}^{*}_{\epsilon=0.005}$ may not contain any valid new sentences apart from the one for which it was formed! At the same time, $\mathbb{H}^{*}_{\epsilon=0.05}$ has up to $48\%$ of generalisability at the expense of only up to $5\%$ of verifiability (cf. Table 7). The effect of the generalisability vs verifiability trade-off can thus be rather severe for geometric subspaces.
This experiment demonstrates the importance of using the generalisability metric, as with only taking into account the verifiability of the subspaces one would choose $\mathbb{H}^{*}_{\epsilon=0.005}$, obtaining mathematically sound but pragmatically useless results. We argue that this is a strong argument for including generalisability as a standard metric in reporting NLP verification results in the future.
Dataset Hyper-rectangles Avg. Volume of $\mathbb{H}$ Generalisability (%) Generalisability (# of sentences) Total Sentences in $\mathcal{A}^{pos}$
RUAR
$\mathbb{H}^{*}_{\epsilon=0.005}$
1.00e-60
1.95
2821
144500
$\mathbb{H}^{*}_{\epsilon=0.05}$
1.00e-30
38.47
55592
144500
$\mathbb{H}_{sh}$
7.55e-11
50.91
73561
144500
Medical
$\mathbb{H}^{*}_{\epsilon=0.005}$
1.00e-60
0.09
10
11209
$\mathbb{H}^{*}_{\epsilon=0.05}$
1.00e-30
28.49
3194
11209
$\mathbb{H}_{sh}$
2.6e-09
37.13
4162
11209
Table 8: Generalisability of the selected geometric subspaces $\mathbb{H}^{*}_{\epsilon=0.005}$, $\mathbb{H}^{*}_{\epsilon=0.05}$ and $\mathbb{H}_{sh}$, measured on the sets of semantic perturbations $\mathcal{A}^{pos}_{RUAR}$ and $\mathcal{A}^{pos}_{medical}$.
0.4.2 Semantic Subspaces and Generalisability-Verifiability Trade-off
The previous subsection has shown that the generalisability-verifiability trade-off is not resolvable by geometric manipulations alone. In this section we argue that using semantic subspaces can help to improve the effects of the trade-off. The main hypothesis that we are testing is: semantic subspaces permit drawing more precise shapes, and this in turn must improve both verifiability and generalisability.
Verifiability of Semantic Subspaces
We will use the construction given in Definition 3. As Table 9 illustrates, we construct several semantic hyper-rectangles on sentences of the positive class using character-level ($\mathbb{H}^{*}_{char}$, $\mathbb{H}^{*}_{del.}$, $\mathbb{H}^{*}_{ins.}$, $\mathbb{H}^{*}_{rep.}$, $\mathbb{H}^{*}_{repl.}$, $\mathbb{H}^{*}_{swap.}$), word-level ($\mathbb{H}^{*}_{word}$) and sentence-level perturbations ($\mathbb{H}^{*}_{pj}$). The subscripts ${}_{char}$ and ${}_{word}$ refer to the kind of perturbation algorithm, while ${}_{del.}$, ${}_{ins.}$, ${}_{rep.}$, ${}_{repl.}$, ${}_{swap.}$ and ${}_{pj}$ refer to the type of perturbation, where pj stands for Polyjuice (see Section 0.3.1). Notice comparable volumes of all these shapes, and compare with $\mathbb{H}^{*}_{\epsilon=0.05}$.
Hyper-rectangles Construction method Avg. volume of $\mathbb{H}$ RUAR Set cardinality of $\mathbb{H}^{*}$ RUAR Avg. volume of $\mathbb{H}$ Medical Set cardinality of $\mathbb{H}^{*}$ Medical
$\mathbb{H}^{*}_{char}$
Set of hyper-rectangles for character perturbations
1.54e-30
3400
7.66e-31
989
$\mathbb{H}^{*}_{word}$
Set of hyper-rectangles for word perturbations
1.28e-30
3400
$\mathbb{H}^{*}_{pj}$
Set of hyper-rectangles for polyjuice sentence perturbations
2.01e-28
989
$\mathbb{H}^{*}_{swap.}$
Set of hyper-rectangles for swapping perturbations
1.57e-31
3400
3.42e-31
989
$\mathbb{H}^{*}_{repl.}$
Set of hyper-rectangles for replacement perturbations
9.84e-31
3400
3.43e-31
989
$\mathbb{H}^{*}_{del.}$
Set of hyper-rectangles for deletion perturbations
3.46e-31
3400
1.24e-32
989
$\mathbb{H}^{*}_{ins.}$
Set of hyper-rectangles for insertion perturbations
3.21e-31
3400
9.11e-33
989
$\mathbb{H}^{*}_{rep.}$
Set of hyper-rectangles for repetition perturbations
1.56e-31
3400
1.06e-32
989
$\mathbb{H}^{*}_{\epsilon=0.05}$
Set of $\epsilon$-cubes generated to match the volume of the hyper-rectangles
1e-30
3400
1e-30
989
Table 9: Sets of semantic subspaces used in the experiments, their cardinality and average volumes of hyper-rectangles. All shapes are eigenspace rotated for better precision.
We pass each set of hyper-rectangles and the network $N_{base}$ to the verifiers ERAN and Marabou to measure verifiability of the subspaces. Table 10 illustrates the verification results obtained using ERAN. From the table, we can infer that the verifiability of our semantic hyper-rectangles is indeed higher than that of the geomtrically-defined hyper-rectangles (Table 7). Furthermore, our semantic hyper-rectangles, while unable to reach the verifiability of $\mathbb{H}^{*}_{\epsilon=0.005}$, achieve notable higher verification than its counterpart of comparable volume $\mathbb{H}^{*}_{\epsilon=0.05}$. From this experiment, we conclude that not only volume, but also precision of the subspaces has an impact on their verifiability.
Dataset Model $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}^{*}_{word}}$ $\boldsymbol{\mathbb{H}^{*}_{char}}$ $\boldsymbol{\mathbb{H}^{*}_{del.}}$ $\boldsymbol{\mathbb{H}^{*}_{ins.}}$ $\boldsymbol{\mathbb{H}^{*}_{rep.}}$ $\boldsymbol{\mathbb{H}^{*}_{repl.}}$ $\boldsymbol{\mathbb{H}^{*}_{swap.}}$ $\boldsymbol{\mathbb{H}^{*}_{pj}}$
RUAR $N_{base}$ 0.00 1.80 0.87 1.62 2.63 1.66 0.94 2.07 -
Medical $N_{base}$ 5.00 - 39.71 39.62 44.66 48.71 37.49 42.60 50.09
Table 10: Verifiability of the baseline DNN on the R-U-A-Robot and the Medical datasets, for a slection of semantic subspaces; for ERAN verifier.
Following these results, Table 11 reports the verification results using Marabou instead of ERAN. As shown, Marabou is able to verify up to $66.83\%$ ($\mathbb{H}^{*}_{rep.}$), while ERAN achieves at most $50.09\%$. This shows that complete and precise verification (Marabou) outperforms abstract interpretation (ERAN). Overall, the Marabou experiment confirms the trends of improved verifiability shown by ERAN and thus confirms our hypothesis about importance of shape precision.
Dataset Model $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}^{*}_{word}}$ $\boldsymbol{\mathbb{H}^{*}_{char}}$ $\boldsymbol{\mathbb{H}^{*}_{del.}}$ $\boldsymbol{\mathbb{H}^{*}_{ins.}}$ $\boldsymbol{\mathbb{H}^{*}_{rep.}}$ $\boldsymbol{\mathbb{H}^{*}_{repl.}}$ $\boldsymbol{\mathbb{H}^{*}_{swap.}}$ $\boldsymbol{\mathbb{H}^{*}_{pj}}$
RUAR $N_{base}$ 1.79 11.69 4.88 4.35 9.72 9.46 5.65 8.07 -
Medical $N_{base}$ 37.96 - 64.03 64.15 64.65 66.83 64.75 64.36 61.57
Table 11: Verifiability of the baseline DNN on the R-U-A-Robot and the Medical datasets, for a slection of semantic subspaces; for Marabou verifier.
Generalisability of Semantic Subspaces
It remains to establish whether the more verifiable semantic subspaces are also more generalisable. Table 12 augments Table 8 with results for best-verifiable semantic subspaces, $\mathbb{H}^{*}_{word}$ and $\mathbb{H}^{*}_{pj}$. It shows that these semantic subpspaces are also the most generalisable, containing, respectively, $47.67\%$ and $28.74\%$ of the unseen sentences.
[…]
Dataset Hyper-rectangles Avg. Volume of $\mathbb{H}$ Generalisability (%) Generalisability (# of sentences) Total Sentences in $\mathcal{A}^{pos}$
RUAR
$\mathbb{H}^{*}_{\epsilon=0.005}$
1.00e-60
1.95
2821
144500
$\mathbb{H}^{*}_{\epsilon=0.05}$
1.00e-30
38.47
55592
144500
$\mathbb{H}^{*}_{word}$
1.28e-30
47.67 68882
144500
Medical
$\mathbb{H}^{*}_{\epsilon=0.005}$
1.00e-60
0.09
10
11209
$\mathbb{H}^{*}_{\epsilon=0.05}$
1.00e-30
28.49
3194
11209
$\mathbb{H}^{*}_{pj}$
2.01e-28
28.74 3222
11209
Table 12: Generalisability of the selected geometric subspaces $\mathbb{H}^{*}_{\epsilon=0.005}$ and $\mathbb{H}^{*}_{\epsilon=0.05}$ and the semantic subspaces $\mathbb{H}^{*}_{word}$ and $\mathbb{H}^{*}_{pj}$, measured on the sets of semantic perturbations $\mathcal{A}^{pos}_{RUAR}$ and $\mathcal{A}^{pos}_{medical}$. Note that the generalisability of $\mathbb{H}^{*}_{sh}$ (Table 8), despite it having the volume 19 order of magnitudes bigger, is only $3\%$ greater than $\mathbb{H}^{*}_{word}$.
0.4.3 Adversarial Training on Semantic Subspaces
In this section we aim to study the effects that adversarial training methods have on the verifiability of the previously defined subspaces. By comparing a range of different training approaches, we show in this section that our new method of adversarial training based on semantic subspaces is the most efficient.
Three kinds of training are deployed in this section:
  1. 1.
    We use $N_{base}$ as the baseline network from the previous experiments for comparison with networks trained for robustness.
  2. 2.
    Data augmentation. We obtain three augmented datasets $\mathcal{Y}\cup\mathcal{A}^{pos}_{char}$, $\mathcal{Y}\cup\mathcal{A}^{pos}_{word}$ and $\mathcal{Y}\cup\mathcal{A}^{pos}_{pj}$. Note that $\mathcal{A}^{pos}_{char}$, $\mathcal{A}^{pos}_{word}$ and $\mathcal{A}^{pos}_{pj}$ are in correspondence with the sets defined in Section 0.4.2. The subscripts char and word denote the kind of perturbation, for which the types are detailed in Tables 3 and 4, while the subscript pj refers to the sentence level perturbations generated with Polyjuice. We train the baseline architecture, using the standard stochastic gradient descent and cross entropy loss, on the augmented datasets, and obtain DNNs $N_{char-aug}$, $N_{word-aug}$ and $N_{pj-aug}$.
  3. 3.
    Customised PGD adversarial training. As defined in Section 0.3.4, we vary the parameter $\mathbb{H}$. In particular, we take for $\mathbb{H}$ different choices of geometric and semantic subspaces we have defined so far. We denote a network trained with PGD algorithm on a subspace $\mathbb{H}^{*}_{name}$ as $N_{name}$. For example, taking $\mathbb{H}=\mathbb{H}_{sh}$ we get $N_{sh}$ as a result of adversarially training the benchmark architecture on $\mathbb{H}_{sh}$. See Table 13 for full listing of the networks we obtain in this way.
We call DNNs of second and third type robustly trained networks. We keep the geometric and semantic subspaces from the previous experiments (shown in Table 9) to compare how training affects their verifiability. Following the same evaluation methodology of experiments as in Sections 0.4.1 and 0.4.2, we use the verifiers ERAN and Marabou to measure verifiability of the subspaces. Table 13 reports accuracy of the robustly trained networks, while the verification results are presented in Tables 14 and 15. From Table 13 we can see that networks trained with data augmentation achieve similar nominal accuracy to networks trained with adversarial training. However, the most prominent difference is exposed in Tables 14 and 15: adversarial training effectively improves the verifiability of the networks, while data augmentation actually decreases it.
Specifically, the adversarially trained networks trained on semantic subspaces ($N_{char}$, $N_{word}$, $N_{pj}$) achieved high verifiability, reaching up to $45.87\%$ for R-U-A-Robot and up to $83.48\%$ for the Medical dataset. This constitutes a significant improvement of the verifiability results compared to $N_{base}$. Looking at nuances, there does not seem to be a single winner subspace when it comes to adversarial training, and indeed in some cases $\mathbb{H}^{*}_{\epsilon=0.05}$ wins over more precise subspaces. All of the subspaces in Table 9 have very similar volume, which accounts for improved performance across all experiments. The particular peaks in performance then come down to particularities of a specific semantic attack that was used while training. For example, the best performing networks are those trained with Polyjuice attack, the strongest form of attack in our range. Thus, if the kind of attack is known in advance, the precision of hyper-rectangles can be further tuned.
Model Dataset Adversarial training Train Accuracy RUAR Test Accuracy RUAR Train Accuracy Medical Test Accuracy Medical
$N_{char-aug}$
$\mathcal{Y}\cup\mathcal{A}^{pos}_{char}$
No
0.9562 ± 0.0026
0.9320 ± 0.0035
0.9908 ± 0.0006
0.9346 ± 0.0030
$N_{word-aug}$
$\mathcal{Y}\cup\mathcal{A}^{pos}_{word}$
No
0.9857 ± 0.0006
0.9459 ± 0.0036
$N_{pj-aug}$
$\mathcal{Y}\cup\mathcal{A}^{pos}_{pj}$
No
0.9819 ± 0.0009
0.9319 ± 0.0039
$N_{char}$
$\mathcal{Y}$
Yes
0.9326 ± 0.0019
0.9251 ± 0.0038
0.9627 ± 0.0005
0.9509 ± 0.0016
$N_{word}$
$\mathcal{Y}$
Yes
0.9368 ± 0.0016
0.9237 ± 0.0029
$N_{pj}$
$\mathcal{Y}$
Yes
0.9505 ± 0.0019
0.9349 ± 0.0032
$N_{\epsilon=0.05}$
$\mathcal{Y}$
Yes
0.9401 ± 0.0017
0.9224 ± 0.0019
0.9605 ± 0.0009
0.9504 ± 0.0024
Table 13: Accuracy of the robustly trained DNNs on the R-U-A-Robot and the Medical datasets. $\mathcal{Y}$ stands for R-U-A-Robot or Medical depending on the column.
Dataset Model $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}^{*}_{word}}$ $\boldsymbol{\mathbb{H}^{*}_{char}}$ $\boldsymbol{\mathbb{H}^{*}_{del.}}$ $\boldsymbol{\mathbb{H}^{*}_{ins.}}$ $\boldsymbol{\mathbb{H}^{*}_{rep.}}$ $\boldsymbol{\mathbb{H}^{*}_{repl.}}$ $\boldsymbol{\mathbb{H}^{*}_{swap.}}$ $\boldsymbol{\mathbb{H}^{*}_{pj}}$
RUAR $N_{char-aug}$ 0.00 0.24 0.00 0.51 1.38 1.09 0.35 1.06 -
$N_{word-aug}$ 0.00 0.24 0.00 0.42 0.31 0.57 0.25 0.92 -
$N_{char}$ 0.00 8.97 4.43 4.81 9.86 11.3 6.91 8.51 -
$N_{word}$ 0.04 10.75 4.05 4.36 8.60 9.52 6.81 7.45 -
$N_{\epsilon=0.05}$ 0.12 10.16 4.18 4.04 8.91 10.17 6.52 7.36 -
Medical $N_{char-aug}$ 0.00 - 7.59 5.28 12.84 11.05 7.92 7.40 26.97
$N_{pj-aug}$ 0.00 - 10.31 8.49 15.67 14.90 9.18 10.58 28.59
$N_{char}$ 5.28 - 50.12 49.78 53.99 57.76 48.02 52.07 55.44
$N_{pj}$ 2.83 - 47.11 46.14 52.12 56.14 44.59 48.27 57.36
$N_{\epsilon=0.05}$ 8.68 - 51.60 50.31 55.67 58.52 50.10 53.65 59.76
Table 14: Verifiability of the robustly trained DNNs on the R-U-A-Robot and the Medical datasets, for a slection of semantic subspaces; for ERAN verifier.
Dataset Model $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}^{*}_{word}}$ $\boldsymbol{\mathbb{H}^{*}_{char}}$ $\boldsymbol{\mathbb{H}^{*}_{del.}}$ $\boldsymbol{\mathbb{H}^{*}_{ins.}}$ $\boldsymbol{\mathbb{H}^{*}_{rep.}}$ $\boldsymbol{\mathbb{H}^{*}_{repl.}}$ $\boldsymbol{\mathbb{H}^{*}_{swap.}}$ $\boldsymbol{\mathbb{H}^{*}_{pj}}$
RUAR $N_{char-aug}$ 0.72 13.90 8.49 7.92 13.67 15.50 9.56 11.88 -
$N_{word-aug}$ 0.24 11.30 3.87 4.05 8.27 8.84 5.71 7.72 -
$N_{char}$ 7.37 41.93 30.41 30.23 38.20 45.87 32.74 36.62 -
$N_{word}$ 12.17 45.12 25.82 25.39 33.85 37.45 26.87 30.99 -
$N_{\epsilon=0.05}$ 18.46 41.93 21.99 20.32 28.13 32.83 23.52 26.74 -
Medical $N_{char-aug}$ 1.14 - 37.05 35.29 41.50 42.47 34.89 37.94 49.65
$N_{pj-aug}$ 5.77 - 39.00 38.66 42.28 44.22 37.29 39.03 38.22
$N_{char}$ 51.70 - 77.59 77.25 77.50 77.98 77.92 78.67 76.58
$N_{pj}$ 57.45 - 81.94 81.47 82.31 83.48 82.47 82.72 82.24
$N_{\epsilon=0.05}$ 62.57 - 79.32 78.57 78.70 80.21 79.40 80.76 66.22
Table 15: Verifiability of the robustly trained DNNs on the R-U-A-Robot and the Medical datasets, for a selection of semantic subspaces; for Marabou verifier.
Model Adversarial training Train Accuracy RUAR Test Accuracy RUAR Train Accuracy Medical Test Accuracy Medical
$N_{sh}$
Yes
0.9339 ± 0.0022
0.9296 ± 0.0013
0.9614 ± 0.0012
0.9429 ± 0.0026
$N_{50}$
Yes
0.9432 ± 0.0014
0.9349 ± 0.0019
0.9556 ± 0.0020
0.9515 ± 0.0012
$N_{100}$
Yes
0.9488 ± 0.0004
0.9418 ± 0.0024
0.9571 ± 0.0011
0.9547 ± 0.0016
$N_{200}$
Yes
0.9509 ± 0.0009
0.9445 ± 0.0014
0.9585 ± 0.0005
0.9543 ± 0.0010
$N_{250}$
Yes
0.9522 ± 0.0008
0.9422 ± 0.0023
0.9607 ± 0.0013
0.9538 ± 0.0022
$N_{\epsilon=0.005}$
Yes
0.9348 ± 0.0021
0.9159 ± 0.0007
0.9624 ± 0.0004
0.9513 ± 0.0009
Table 16: Accuracy of the robustly (adversarially) trained DNNs on the R-U-A-Robot and the Medical datasets.
Dataset Model $\boldsymbol{\mathbb{H}_{sh}}$ $\boldsymbol{\mathbb{H}^{*}_{50}}$ $\boldsymbol{\mathbb{H}^{*}_{100}}$ $\boldsymbol{\mathbb{H}^{*}_{200}}$ $\boldsymbol{\mathbb{H}^{*}_{250}}$ $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.005}}$
RUAR $N_{sh}$ 0.00 0.00 1.33 0.52 0.41 88.62
$N_{50}$ 0.00 0.00 0.00 0.00 0.41 90.02
$N_{100}$ 0.00 0.00 0.00 0.00 0.41 92.74
$N_{200}$ 0.00 0.00 0.00 0.00 0.08 93.54
$N_{250}$ 0.00 0.00 0.00 0.00 0.33 93.86
$N_{\epsilon=0.005}$ 0.00 0.00 0.00 0.00 0.33 98.22
Medical $N_{sh}$ 0.00 0.00 0.00 2.50 4.40 97.47
$N_{50}$ 0.00 0.00 1.08 3.60 6.00 98.79
$N_{100}$ 0.00 0.00 1.08 3.00 5.04 99.09
$N_{200}$ 0.00 0.00 1.08 2.90 4.96 99.05
$N_{250}$ 0.00 0.00 0.00 2.90 4.40 98.73
$N_{\epsilon=0.005}$ 0.00 0.00 0.00 2.30 4.32 99.60
Table 17: Verifiability of the robustly (adversarially) trained DNNs on the R-U-A-Robot and the Medical datasets, for a selection of geometric subspaces; for ERAN verifier.
As a final note, we report results from robust training using the subspaces of Experiment 0.4.1 (Table 6). Table 16 reports the accuracy and the details of the robustly trained networks on those subspaces, while the verification results are presented in Table 17. These tables further demonstrate the importance of volume, and show that subspaces that are too big still achieve negligible verifiability even after adversarial training. Generalisability of the shapes used in Tables 13 - 17 remains the same, see Tables 812.
0.5 NLP Case Studies
The purpose of this section is two-fold:
  • Firstly, the case studies we present here apply the NLP Verification Piplene set out in Section 0.2.5 using more modern NLP tools. Notably, in this section we try different LLMs to embed sentences and replace Polyjuice with the LLM vicuna-13b [^8] Using the following API: https://replicate.com/replicate/vicuna-13b/api., a state-of-the-art open source chatbot trained by fine-tuning LLaMA [115] on user-shared conversations collected from ShareGPT  [^9] https://sharegpt.com/. For further details, please refer to Section 0.3.1.
    To support the work with varying different components of the NLP Verification pipeline, we released the tool ANTONIO [21] that implements different choices for each pipeline component, as shown in Figure 4.
    Figure 4: Tool ANTONIO that implements a modular approach to the NLP verification pipeline introduced in this paper.
  • And secondly, and perhaps more fundamentally, the purpose of this section is to draw attention to the importance of the purely NLP parts of the pipeline, namely the parts that have to do with generation, processing, perturbing, and embedding sentences. This part is largely ignored in NLP verification papers. We show in this section that taking NLP decisions for granted may compromise the practical value of the NLP verification pipelines.
    To show what we mean by these let us show some examples:
    Example 3 (Failure of the Embedding Function; Falsifiable Subspaces).
    As an example, imagine a scenario when a DNN was verified on subpaces of a class $c_{j}$. Then it is used on new, yet unseen sentences. Imagine also that the implicit assumption that the embedding function embeds only semantically similar sentences close in the embedding space, fails. Then, the new sentences that should belong to a class $c_{i}$ may be by mistake embedded into the verified subspace (of class $c_{j}$). It will be wrong to assure the user that the verified subspace guarantees that these new sentences belong to $c_{j}$ (even if the DNN is guaranteed to classify them as $c_{j}$)! In fact we will say that new sentences of class $c_{i}$ that fall inside the verified subspace of class $c_{j}$ falsify the verified subspace. Omitting to acknowledge and report the problem of falsifiable verified subspaces may have different implications, depending on the usage scenarios. If the verified subspace of class $c_{j}$ is used to recognise and censor sensitive (‘dangerous’) sentences, new sentences of class $c_{i}$ that fall inside of the verified subspace will be wrongly censored; which in turn may make interaction with the chatbot impractical. But if the verified subspace of class $c_{j}$ certifies safe sentences, then potentially dangerous sentences of class $c_{i}$ could be wrongly admitted as verifiably safe. Note that this problem can be related to the well-known problem of false positives and false negatives in machine learning: the new sentences of class $c_{i}$ that fall within the verified subpspace of class $c_{j}$ at the same time will be false positives or negatives for that DNN (depending whether $c_{j}$ is a positive or negative class).
    Example 4 (Failure of the Sentence Perturbation Algorithm).
    Another assumption that most NLP verification papers make is that we can algorithmically generate sentence perturbations in a way that is guaranteed to retain their original semantic meaning. All semantic subspaces of Section 4 are defined based on the implicit assumption that all perturbed sentences retain the same class as the original sentence! But if this assumption fails, we will once again end up drawing falsifiable semantic subspaces around embeddings of sentences belonging to different classes, with similar consequences as in the previous example.
As a consequence of the embedding gap, we obtain a seemingly paradoxical situation, when in principle, the same semantic subspace can be both formally verified and empirically falsified! Formal verification ensures that all sentences embedded within the semantic subspace will be classified identically by the given DNN; but empirical falsification of the semantic subspace comes from appealing to the semantic meaning of the embedded sentences, – something that only human can define exactly and verification methodology does not directly define.
In the light of this limitation, How can we measure and improve the quality of the purely NLP components of the pipeline, in a way that ensures that the NLP verification pipeline is sufficiently reliable and practically usable? – is the main question that we answer in this section.
0.5.1 Practicalities of NLP Verification for Critically Safe Applications
Role of false positives and false negatives.
Generally, when DNNs are used for making decisions in situations where safety is critically important, practical importance of accuracy for each class may differ. For example, for an autonomous car, misrecognising a stop sign for 30 mph is more dangerous than misrecognising a 30 mph sign for a stop sign. Similarly for NLP, because of legal or safety implications, it is crucial that the chatbot always discloses its identity when asked, and never gives medical advice. We want to avoid false negatives altogether, i.e. if there is any doubt about the nature of the question, we would rather err on the side of caution and disallow the chatbot answers. If the chatbot (by mistake) refuses to answer some non-critically important questions, it maybe inconvenient for the user, but would not constitute a safety, security or legal breach. Thus, false positives maybe tolerated.
On the technical level, this has two implications:
  1. 1.
    Firstly, if we use DNN on its own, without verification, we may want to report precision and recall  [^10] Recall that precision is defined as true positivestrue positives+false $\frac{\text{true positives}}{\text{true positives}+\text{false negatives}}$ and recall is defined as true positivestrue positives+false $\frac{\text{true positives}}{\text{true positives}+\text{false positives}}$. in addition to accuracy. Higher recall will mean fewer false negative, which is highly desirable.
  2. 2.
    In the literature and in this paper, it is assumed that verified DNNs serve as filters that allow the larger system to use machine learning in a safe manner. Verification covers only subspaces that characterise the critically important class or classes. I.e.: DNN verification guarantees that all meaning-preserving perturbations of a stop sign, or “Are you a Robot?” sentence will be classified by the given DNN in the same way as the original unperturbed sample. In practice, only DNN outputs on these safe regions are trusted, and DNN output bears no verification guarantees for all other inputs. This approach completely eliminates the effect of any false negatives. But the verified subspaces may still contain false positives (cf. Example 3) and in this sense be “falsifiable”, which may not be pragmatically desirable (but should not constitute a security threat).
Bearing these considerations in mind, throughout this section we report precision and recall of models, in addition to accuracy, to see whether one or another class is particularly challenging to recognise. Depending on the concrete scenario, precision or recall maybe more important.
We also introduce the falsifiability metric, that should be used together with verifiability and generalisability metrics in all NLP benchmarks.
0.5.2 Robustness of Large Language Models
We start first with gaining an intuition about robustness of modern LLMs. We will prompt an LLM to generate the desired output without additional fine-tuning, i.e. a zero-shot model. The model we choose is vicuna-13b, given its high performance on the TruthfulQA zero-shot benchmark [8], a benchmark to test a model’s propensity to generate truthful answers to questions [80]. Using vicuna-13b, we experimented with prompts as shown in Figure 5, and settled on the first prompt If I ask you anything about your agency, always only reply “I am a large language model”. If not, reply normally.“[Example]”? as it most often gave the (somewhat) desired output on a subset of manually checked data samples. Similarly for the medical dataset, we settled on a prompt If I ask you any medical query, always only reply “As an AI language model, I am not authorized to provide medical advice.”. If not, reply normally.“[Example]”.
Figure 5: Zero-shot prompts with 2 basic examples from the R-U-A-Robot dataset. Answers from vicuna-13b are given in italics. A1 and A2 represent different answers to the same prompt, illustrating a lack of consistency in the output.
For our zero-shot model, results are reported on the test set. Using the test sets of our chosen datasets, we use rule-based methods to filter through the answers [^11] Additionally omitting $\approx 40\%$ of answers which returned empty due to API errors. . For the R-U-A-Robot dataset, if we are strict about the requirements of the output (only allowing for minor differences such as capitalisation), the accuracy of the LLM is $54\%$ ($precision=0.51$, $recall=0.58$, and $F1=0.54$). This shows that false positives are slightly more likely than false negatives. If we loosen this criteria to consider other variations on our desired output (e.g. ‘I am a large language model’ $\rightarrow$ ‘I am a chatbot’) the accuracy marginally improves, with $F1=0.56$. For the medical safety dataset, the results are $precision=0.58$, $recall=0.70$, and $F1=0.64$, indicating comparatively less false negatives.
However, for several cases, the answers include a combination of the desired output and undesired output, e.g. ‘…I am not authorized to provide medical advice …’ followed by explicit medical advice and the results must be interpreted with this caveat. Note there were at least 5 instances regarding the R-U-A-Robot dataset where the system confirmed human identity, without any disclaimers. Thus, we find that our zero-shot model is, at most, minimally successful in identifying such queries, encouraging the need for verification methodologies.
0.5.3 Verification Pipeline Setup and the Role of Embedding Functions
For all experiments in this section, we set up the NLP verification pipelines as shown in Table 18; and implement them using the tool ANTONIO [21] that we implemented to support this paper.
In setting up the pipeline, we follow key conclusions made in Section 0.4 about successful verification strategies, namely:
  • semantic subspaces should be preferred over geometric subpspaces;
  • best semantic subspaces result from using stronger NLP perturbations;
  • adversarial training should be done on these “best” subpspaces;
  • Marabou performs best thanks to its precision.
Based on this, we further strengthen NLP perturbations: we substitute Polyjuice used in the previous section with Vicuna. Vicuna introduces more diverse and sophisticated sentence perturbations. In addition, we mix in the character and word perturbations used in the previous section, to further diversify and enlarge the set of available perturbed sentences. In the terminology of Section 0.4.1, we obtain the sets of perturbed sentences $\mathcal{A}^{pos}_{pert}$ and $\mathcal{A}^{neg}_{pert}$ where $pert$ is itself a combined kind of perturbations, i.e. $pert=\{char,word,vicuna\}$. Table 18 also uses notation $\mathcal{A}^{pos}_{pert^{\diamondsuit}}$ and $\mathcal{A}^{neg}_{pert^{\diamondsuit}}$ to refer to filtered sets, this terminology will be introduced in the next subsection. Having $\mathcal{A}^{pos}_{pert}$, the set of hyper-rectangles $\mathbb{H}^{*}_{pert}$ is obtained according to Definition 3. Accordingly, we set the adversarial training to explore hyper-rectangles in $\mathbb{H}^{*}_{pert}$, and obtain the adversarially robust DNN $N_{pert}$.
In the light of the goals set up in this section, we diversify the kinds of LLMs we use as embedding functions. We use the sentence transformers package from Hugging Face originally proposed in [101] (as our desired property is to give guarantees on entire sentences). Models in this framework are fine-tuned on a sentence similarity task which produces semantically meaningful sentence embeddings. We select 3 different encoders to experiment with the size of the model. For our smallest model, we choose all-MiniLM-L6-v2, an s-transformer based on MiniLMv2 [119], a compact version of the BERT architecture [32] that has comparable performance. Additionally we choose 2 GPT-based models, available in the S-GPT package [92]. We refer to these 3 models as s-bert 22M, s-gpt 1.3B, and s-gpt 2.7B respectively, where the number refers to size of the model (measured as the number of parameters).
Pipeline Component Section 0.5 Choices Additional Details
0. Choice of Datasets
RUAR, Medical
Same as in Section 0.4 experiments. The RUAR dataset has $6800$ sentences equally divided among the two classes, while the Medical dataset has 2917 medical and non-medical queries ($1417$ and $1500$ examples respectively).
1. Generating Sentence Perturbations
$\mathcal{A}^{pos}_{pert}$, $\mathcal{A}^{pos}_{pert^{\diamondsuit}}$, $\mathcal{A}^{neg}_{pert}$, $\mathcal{A}^{neg}_{pert^{\diamondsuit}}$
With $pert=\{char,word,vicuna\}$, the resulting set of sentences $\mathcal{A}_{pert}$ has $54400$ sentences for RUAR and $15824$ sentences for Medical. The superscript ${}^{\diamondsuit}$ refers to filtering that will be introduced in Section 0.5.4.
2. Embedding Sentences into Real Vector Space
s-bert 22M, s-gpt 1.3B, s-gpt 2.7B
In the experiments of Section 0.4 only s-bert 22M was used.
3. Defining Semantic Subspaces based on Sentence Perturbations
$\mathbb{H}^{*}_{pert}$, $\mathbb{H}^{*}_{pert^{\diamondsuit}}$
$\mathbb{H}^{*}_{pert}$ and $\mathbb{H}^{*}_{pert^{\diamondsuit}}$ are obtained on $\mathcal{A}^{pos}_{pert}$ and $\mathcal{A}^{pos}_{pert^{\diamondsuit}}$, respectively. Their cardinality is $3400$ for RUAR and $989$ for Medical.
  • Volume of $\mathbb{H}^{*}_{pert}$ for RUAR is $191.83e-$ 19 (s-bert 22M), $353.24e+$ 35 (s-gpt 1.3B), $363.30e+$ 36 (s-gpt 2.7B).
  • Volume of $\mathbb{H}^{*}_{pert^{\diamondsuit}}$ for RUAR is $252.43e-$ 25 (s-bert 22M), $271.54e+$ 27 (s-gpt 1.3B), $283.10e+$ 28 (s-gpt 2.7B).
  • Volume of $\mathbb{H}^{*}_{pert}$ for Medical is $223.13e-$ 22 (s-bert 22M), $331.70e+$ 33 (s-gpt 1.3B), $332.10e+$ 33 (s-gpt 2.7B).
  • Volume of $\mathbb{H}^{*}_{pert^{\diamondsuit}}$ for Medical is $283.65e-$ 28 (s-bert 22M), $253.30e+$ 25 (s-gpt 1.3B), $273.83e+$ 27 (s-gpt 2.7B).
4. Training Robust DNNs using Semantic Subspaces
$N_{base}$, $N_{pert}$, $N_{pert^{\diamondsuit}}$
$N_{base}$ is obtained as in Section 0.4, while $N_{pert}$ and $N_{pert^{\diamondsuit}}$ are obtained through our adversarial training on $\mathbb{H}^{*}_{pert}$ and $\mathbb{H}^{*}_{pert{\diamondsuit}}$, respectively.
5. Verifying resulting DNNs on the given semantic subpspaces
Marabou
Same settings as in Section 0.4
Table 18: Section 0.5 NLP verification pipeline setup, implemented using ANTONIO. Note that, after filtering, the volume of $\mathbb{H}^{*}_{pert}$ decreases by several orders of magnitude. Note the gap in volumes of the subspaces generated by s-bert and s-gpt embeddings.
For illustration as well as the initial sanity check, we report accuracy of the obtained models, per each of the chosen embedding functions, in Table 19. Overall the figures are as expected: the choice of the encoder influences accuracy. Thus, using an additional DNN as a filter (as opposed to the zero-shot model), dramatically increases the accuracy of the overall system, from $54-64\%$ in the previous section to the range of $76-95\%$ in Table 19.
Dataset Model Test set Perturbed test set
Precision Recall F1 Precision Recall F1
RUAR $N_{zero-shot}$ 0.5167 0.5835 0.5481 - - -
$N_{base}$ (s-bert 22M) 0.9568 0.9129 0.9344 0.9477 0.7186 0.8174
$N_{pert}$ (s-bert 22M) 0.8497 0.9863 0.9129 0.8125 0.9466 0.8745
$N_{base}$ (s-gpt 1.3B) 0.9620 0.8725 0.9151 0.9545 0.6738 0.7898
$N_{pert}$ (s-gpt 1.3B) 0.6303 0.9980 0.7724 0.6126 0.9860 0.7554
$N_{base}$ (s-gpt 2.7B) 0.9674 0.8729 0.9177 0.9549 0.6982 0.8066
$N_{pert}$ (s-gpt 2.7B) 0.6018 0.9980 0.7508 0.5846 0.9899 0.7350
Medical $N_{zero-shot}$ 0.5895 0.7022 0.6409 - - -
$N_{base}$ (s-bert 22M) 0.9523 0.9325 0.9423 0.9520 0.8964 0.9234
$N_{pert}$ (s-bert 22M) 0.9335 0.9736 0.9531 0.9238 0.9517 0.9376
$N_{base}$ (s-gpt 1.3B) 0.9193 0.8811 0.8998 0.9217 0.8417 0.8798
$N_{pert}$ (s-gpt 1.3B) 0.8441 0.9627 0.8938 0.8315 0.9470 0.8854
$N_{base}$ (s-gpt 2.7B) 0.9325 0.8929 0.9123 0.9289 0.8479 0.8866
$N_{pert}$ (s-gpt 2.7B) 0.8603 0.9656 0.9098 0.8488 0.9499 0.8964
Table 19: Accuracy of the models on the test/perturbation set. The average standard deviation is $0.0049$.
Dataset Model Test set Perturbed test set
Precision Recall F1 Precision Recall F1
RUAR $N_{base^{\text{ \ }}}$ (s-bert 22M) 0.9568 0.9129 0.9344 0.9477 0.7186 0.8174
$N_{pert^{\diamondsuit}}$ (s-bert 22M) 0.8507 0.9894 0.9148 0.8289 0.9412 0.8815
$N_{base^{\text{ \ }}}$ (s-gpt 1.3B) 0.9620 0.8725 0.9151 0.9545 0.6738 0.7898
$N_{pert^{\diamondsuit}}$ (s-gpt 1.3B) 0.6493 0.9965 0.7862 0.6356 0.9808 0.7713
$N_{base^{\text{ \ }}}$ (s-gpt 2.7B) 0.9674 0.8729 0.9177 0.9549 0.6982 0.8066
$N_{pert^{\diamondsuit}}$ (s-gpt 2.7B) 0.6305 0.9969 0.7724 0.6143 0.9852 0.7566
Medical $N_{base^{\text{ \ }}}$ (s-bert 22M) 0.9523 0.9325 0.9423 0.9520 0.8964 0.9234
$N_{pert^{\diamondsuit}}$ (s-bert 22M) 0.9313 0.9717 0.9511 0.9251 0.9493 0.9370
$N_{base^{\text{ \ }}}$ (s-gpt 1.3B) 0.9193 0.8811 0.8998 0.9217 0.8417 0.8798
$N_{pert^{\diamondsuit}}$ (s-gpt 1.3B) 0.8445 0.9571 0.8972 0.8426 0.9424 0.8896
$N_{base^{\text{ \ }}}$ (s-gpt 2.7B) 0.9325 0.8929 0.9123 0.9289 0.8479 0.8866
$N_{pert^{\diamondsuit}}$ (s-gpt 2.7B) 0.8682 0.9656 0.9143 0.8560 0.9474 0.8993
Table 20: Accuracy of the models on the test/perturbation set, after filtering. The average standard deviation is $0.0049$.
Looking into nuances, one can further notice the following:
  1. 1.
    Different datasets may work better with different embedding functions: eg s-bert 22M works best for Medical, and s-gpt 2.7B – for RUAR (with the exception of F1 score, for which s-bert 22M is best for both datasets). Smaller GPT model s-gpt 1.3B is systematically worse for both datasets;
  2. 2.
    Best accuracy may not be the best indicator of performance, depending on the scenario of use. Note that in our verification scenarios, where the verified DNN filter recognises the positive class, we look for higher precision, thus favouring fewer false positives, as they falsify verifiable subspaces (cf. Example 3). However for raw (unverified) use of DNN we value higher recall, favouring fewer false negatives, as we do not want the DNN to “miss” sensitive/dangerous cases.
    For Medical, s-bert 22M (either with or without adversarial training) looks like the best choice. However, for RUAR, the choice of the embedding function has a huge effect:
    • for general accuracy, s-bert 22M is the best choice (difference with the worst choice of the embedding function is $12-16$ percentage points),
    • for scenarios when one uses unverified DNN, $N_{pert}$ (s-gpt 2.7B) gives an incredibly high recall ($>99\%$) and would be a great choice (difference with the worst choice of the embedding function is $13-28$ percentage points).
    • However, if one wanted to use the same $N_{pert}$ (s-gpt 2.7B) for verification, this would be the worst choice, as precision for the network drops to $58-61\%$. For verification, $N_{base}$ (s-gpt 2.7B) $N_{base}$ or s-bert 22M would be better choices (with precision $>95\%$).
  3. 3.
    Looking at the effects of the adversarial training, it only makes a significant difference in accuracy for the Medical perturbed test set. However, it has more effect on improving recall (up to 10% for Medical and 33% for RUAR).
  4. 4.
    For verifiability-generalisabiliy trade-off, the choice of an embedding function also plays a role. Table 25 shows that s-gpt models exhibit lower verifiability compared to s-bert models. This observation also concurs with the findings in Section 0.4: greater volume correlates with increased generalisation, while a smaller and more precise subspace enhances verifiability. Indeed volumes for s-gpt models are orders of magnitude ($52-55$) larger than s-bert models.
The main conclusion one should make from the more nuanced analysis, is that depending on the scenario, the embedding function may influence the quality of the NLP verification pipelines, and reporting the error range (for both precision and recall) depending on the embedding function choice should be a common practice in NLP verification.
0.5.4 Analysis of Perturbations and Falsifiability
Recall that two problems were identified as potential causes of falsifiable semantic subspaces: the imprecise embedding functions and invalid perturbations (i.e. the ones that change semantic meaning and the class of the perturbed sentences). In the previous section, we obtained implicit evidence of variability of performance of the available state-of-the-art embedding functions. In this section, we turn our attention to analysis of perturbations. As outlined in [136], to be considered valid, the perturbations should be semantically similar to the original, grammatical and have label consistency, i.e. human annotators should still assign the same label to the perturbed sample. Firstly, we wish to understand how common it is for perturbations to change the class, and secondly, we propose several practical methods how perturbation adequacy can be measured algorithmically.
Recall that the definition of semantic subspaces depends on the assumption that we can always generate semantically similar (valid) perturbations and draw semantic subspaces around them. Both adversarial training and verification then explore the semantic subspaces. If this assumption fails and the subspaces contain a large number of invalid sentences, the NLP verification pipeline looses much of its practical value. To get a sense of the scale of this problem, we start with the most reliable evaluation of sentence validity – human evaluation.
Understanding the Scale of the Problem
For the human evaluation, we labelled a subset of the perturbed datasets considering all validity criteria. Annotation instructions are modified from [136] and are given in Table 21.
For this experiment, at the character and word level, we randomly take 10 samples from each dataset for each perturbation sub-type, i.e. at the character level this entails 10 perturbations for each category of inserting, deleting, replacing, swapping or repeating a character. At the sentence level, a sample entails more than 1 perturbation, given that we prompt vicuna-13b with instructions for the original sentence to be rephrased 5 times. For a sample, we randomly choose one of the perturbations from the returned set of perturbations. In this way, we take 50 samples from each dataset. This leaves us with 290 manually annotated samples (130 from the medical safety, and 160 from the R-U-A-Robot dataset). Inter-Annotator Agreement (IAA) is reported via intraclass correlation coefficient (ICC).
Human Evaluation.
ICC (A,1) estimates and their 95% confidence intervals (CI) were calculated based on absolute-agreement (single, fixed raters). Using cutoffs provided by [78], agreement was determined to be MODERATE for semantic similarity (F = 4.4 df (289), p<.001, 95% CI = [0.56,0.69]), BELOW SATISFACTORY for grammaticality (ICC = 0.43, p <.001, 95% CI = [0.34,0.52]) and BELOW SATISFACTORY for label consistency (ICC = 0.29, p<.001, 95% CI = [0.18, 0.39]). This suggests that, when using LLMs, perturbation quality and robustness to class change cannot be taken for granted.
Limitations. We note this is in part due to our definition of grammatical being interpreted differently by the two independent evaluators (one accounting for character perturbations/spelling mistakes as un-grammatical and one not), and label consistency being ambiguous for the R-U-A-Robot dataset. Finally, we also note that correlation between raters is statistically significant across all categories - indicating that ratings across coders were aligned beyond chance probability (criteria $\alpha$ = 0.05). Future replications are warranted.
Criteria Instructions
Semantic similarity
Evaluate whether the original and the modified sentence have the same meaning on a scale from 1 to 4, where 1 is ‘The modified version means something completely different’ and 4 means ‘The modified version has exactly the same meaning’.
Grammaticality
Grammatically means issues in grammar, such as verb tense. Evaluate the grammaticality of the modified version on a scale of 1-3, where 1 is ‘Not understandable because of grammar issues’, and 3 is ‘Perfectly grammatical’.
Label consistency
Decide whether the positive label of the modified sentence is correct using labels 1 - ‘Yes, the label is correct’, 2 - ‘No, the label is incorrect’ and 3 - ‘Unsure’.
Table 21: Annotation instructions for manual estimation of the perturbation validity.
Effective Ways to Measure and Report Perturbation Validity
Although no geometric or algorithmic method will ever be able to match to the full extent the human perception and interpretation of sentences, we can still formulate a number of effective methods that give an implicit characterisation of validity of perturbations utilised when defining semantic subspaces. We propose three:
  1. 1.
    Using cosine similarity of embedded sentences, we can implicitly characterise semantic similarity
  2. 2.
    Using the ROUGE-N method [79] – a standard technique to evaluate natural sentence overlap, we can measure lexical and syntactic validity
  3. 3.
    Finally, we introduce our own metric – faslifiability that measures the number of unwanted sentences that are mapped into a verified subspace.
We proceed to describe and evaluate each of them in order.
Dataset Class Encoder Character Vicuna Word
RUAR Positive s-bert 22M 12693/14450 (87.84%) 8190/12223 (67.00%) 17209/17340 (99.24%)
s-gpt 1.3B 14170/14450 (98.06%) 9677/12223 (79.17%) 17123/17340 (98.75%)
s-gpt 2.7B 14168/14450 (98.05%) 10024/12223 (82.01%) 17112/17340 (98.69%)
Negative s-bert 22M 11288/14450 (78.12%) 5008/8511 (58.84%) 2167/17309 (12.52%)
s-gpt 1.3B 13315/14450 (92.15%) 5943/8511 (69.83%) 2164/17309 (12.50%)
s-gpt 2.7B 13404/14450 (92.76%) 6377/8511 (74.93%) 2229/17309 (12.88%)
Medical Positive s-bert 22M 4753/4945 (96.12%) 4282/4651 (92.07%) 5908/5934 (99.56%)
s-gpt 1.3B 4914/4945 (99.37%) 4219/4651 (90.71%) 5909/5934 (99.58%)
s-gpt 2.7B 4910/4945 (99.29%) 4309/4651 (92.65%) 5917/5934 (99.71%)
Negative s-bert 22M 5037/5260 (95.76%) 947/1137 (83.29%) 6271/6312 (99.35%)
s-gpt 1.3B 5216/5260 (99.16%) 983/1137 (86.46%) 6258/6312 (99.14%)
s-gpt 2.7B 5220/5260 (99.24%) 1017/1137 (89.45%) 6280/6312 (99.49%)
Table 22: Number of perturbations kept for each model after filtering with cosine similarity > 0.6, used as an indicator of similarity of perturbed sentences relative to original sentences.
Cosine similarity.
Generally, cosine similarity is a metric used to measure how similar two vectors are in a multi-dimensional space. The cosine similarity between two vectors is calculated based on the cosine of the angle between them. Given two vectors $v_{1}$ and $v_{2}$, the cosine similarity is computed using the following formula:
$\text{CoS}(v_{1},v_{2})=\frac{v_{1}\cdot v_{2}}{\|v_{1}\|\|v_{2}\|}$
where $\cdot$ is the dot product and $\|v\|=\sqrt{v\cdot v}$. The resulting value ranges from 00 to $1$. A value of $1$ indicates that the vectors are parallel (highest similarity), while 00 means that the vectors are orthogonal (no similarity).
Recall that Section 0.3.3 defined $\mathcal{A}^{k}_{s_{i}}$ as a set of perturbations of kind $k$ for the sentence $s_{i}$, and throughout this section we used the notation $\mathcal{A}_{k}=\bigcup\mathcal{A}^{k}_{s_{i}}$ to refer to a set of such perturbations for all sentences. Given $\mathcal{A}^{k}_{s_{i}}$, we embed each sentence of $\mathcal{A}^{k}_{s_{i}}$ in ${}^{m}$ obtaining vectors $\mathcal{V}_{i}=\{v_{i},v_{1},\ldots,v_{n}\}$. Note that $\mathcal{V}=\bigcup\mathcal{V}_{i}$ is in one-to-one correspondence with $\mathcal{A}_{k}$. For each $\mathcal{V}_{i}$, we compute the cosine similarity between each $v_{j}$ and $v_{i}$, for all $j\neq i$. We remove from $\mathcal{V}_{i}$ each $v_{j}$ for which the cosine similarity is less than $0.6$ and obtain $\mathcal{V}_{i^{\diamondsuit}}$.
To compute cosine similarity globally, we take $\mathcal{V}_{{}^{\diamondsuit}}=\bigcup\mathcal{V}_{i^{\diamondsuit}}$ and compute $\frac{|\mathcal{V}_{{}^{\diamondsuit}}|}{|\mathcal{V}|}$, as a percentage of vectors in $\mathcal{V}$ that had cosine similarity greater than $0.6$. Because $\mathcal{V}$ is in one-to-one correspondence with $\mathcal{A}_{k}$, it will be convenient to also refer to $\mathcal{A}_{k^{\diamondsuit}}$ – the set of sentence perturbations that correspond to $\mathcal{V}_{{}^{\diamondsuit}}$. Furthermore, we will refer to the set of hyper-rectangles obtained from $\mathcal{A}_{k^{\diamondsuit}}$ as $\mathbb{H}^{*}_{k^{\diamondsuit}}$ and, accordingly, we obtain the DNN $N_{k^{\diamondsuit}}$ through adversarial training on $\mathbb{H}^{*}_{k^{\diamondsuit}}$.
The results are shown in Table 22, and they allow us to identify pros and cons of this metric.
  • Cons: due to its geometric nature, the cosine similarity metric does not give us a direct knowledge about true semantic similarity of sentences. Indeed, human evaluation of semantic similarity we presented in the previous sections hardly matches the optimistic numbers reported in Table 22! Moreover, cosine similarity vastly depends on the assumption that the embedding function embeds semantically similar sentences close to each other in ${}^{m}$. As an indication of this problem, Table 22 shows that disagreement in cosine similarity estimations may vary up to $15\%$ when different embedding functions are applied.
  • Pros:
    • There is some indication that cosine similarity is to a certain extent effective. For example, we have seen in Section 0.5.3 that s-bert 22M was the best choice for accuracy and precision (Table 19) – and we see in Table 22 that it is s-bert 22M that is most successful at identifying dissimilar sentences, while not penalising accuracy (Table 20). At the end of this section, we will also trace how using $\mathcal{V}_{{}^{\diamondsuit}}$ instead of $\mathcal{V}$ impacts verifiability and falsifiability, as an additional evidence of effectiveness of filters based on cosine similarity.
    • Cosine similarity metric is general (i.e. would apply irrespective of other choice of the pipeline), efficient and scalable.
Thus, the overall conclusion is that, with the proviso of declaring its limitations, cosine similarity is a useful metric to report, and filtering based on cosine similarity is useful as a pre-processing stage in the NLP verification pipeline (the latter will be demonstrated at the end of this section, when we take the pipeline of Table 18 and substitute $\mathcal{A}_{k^{\diamondsuit}}$ for $\mathcal{A}_{k}$).
Dataset ROUGE-N Precision Recall
No filtering Filtering No filtering Filtering
s-bert 22M s-gpt 1.3B s-gpt 2.7B s-bert 22M s-gpt 1.3B s-gpt 2.7B
RUAR ROUGE-1 0.500 0.568 0.545 0.537 0.281 0.635 0.612 0.604
ROUGE-2 0.557 0.342 0.320 0.314 0.312 0.382 0.358 0.352
ROUGE-3 0.511 0.208 0.190 0.185 0.285 0.230 0.210 0.205
Medical ROUGE-1 0.451 0.466 0.469 0.465 0.230 0.553 0.555 0.551
ROUGE-2 0.529 0.242 0.246 0.243 0.268 0.285 0.288 0.285
ROUGE-3 0.471 0.131 0.135 0.133 0.238 0.156 0.159 0.157
Table 23: ROUGE-N scores comparing the original samples with Vicuna perturbations (of the positive class) for lexical overlap.
Dataset ROUGE-N Precision Recall
No filtering Filtering No filtering Filtering
s-bert 22M s-gpt 1.3B s-gpt 2.7B s-bert 22M s-gpt 1.3B s-gpt 2.7B
RUAR ROUGE-1 0.731 0.748 0.747 0.743 0.501 0.767 0.769 0.765
ROUGE-2 0.738 0.524 0.521 0.514 0.504 0.532 0.532 0.525
ROUGE-3 0.710 0.350 0.347 0.340 0.483 0.349 0.346 0.339
Medical ROUGE-1 0.670 0.674 0.678 0.676 0.410 0.710 0.714 0.712
ROUGE-2 0.694 0.415 0.422 0.419 0.422 0.434 0.441 0.438
ROUGE-3 0.657 0.247 0.254 0.252 0.399 0.258 0.263 0.260
Table 24: ROUGE-N scores comparing the original samples with Vicuna perturbations (of the positive class) for syntax overlap.
ROUGE-N
We additionally calculate lexical and syntactic variability of the generated vicuna-13b output by reporting ROUGE-N $precision$ and $recall$ scores (i.e. which measures $ngram$ overlap) [79], where $n\in[1,2,3]$. Intuitively if $s_{i}$ is a sentence from the dataset and $s_{j}$ a perturbation of $s_{i}$, ROUGE-N is an overlap measure, which measures:
  • $precision$, i.e. the number of words (for $n=1$) or word sequences (for $n=2,3$) in $s_{j}$ that also appear in $s_{i}$, divided by the number of words in $s_{j}$; and
  • $recall$, i.e. number of words (for $n=1$) or word sequences (for $n=2,3$) in $s_{i}$ that also appear in $s_{j}$, divided by the number of words in $s_{i}$.
For lexical ROUGE-N, we compare the strings of the original sample to the perturbations, while for syntax we follow the same procedure, but using the corresponding POS tags [116]. Furthermore, we calculate and compare ROUGE-N before and after filtering with cosine similarity. Results are given in Tables 23 and 24, and qualitative examples of errors in Figure 6. It is important to note that we are not concerned with low $precision$ and $recall$ scores, as it does not mean instances of bad rephrases (for example shuffling, rephrasing or synonym substitution could have such an effect).
Example 5 (Validity of Perturbations).
Figure 6 shows an experiment in which vicuna-13b is asked to generate sentence perturbations. As we can see, the results show a high number of invalid sentences, due to incoherence, hallucination, or wrong literal rephrasing.
Figure 6: Analysis of some common issues found in the vicuna-13b generated perturbations.
  1. 1.
    Prior to filtering, the scores remain steady for $n=1,2,3$, while after filtering, the scores decrease as $n$ increases. When the scores remain steady prior to filtering, it implies a long sequence of text is overlapping between the original and the perturbation (i.e. for unigrams, bigrams and trigrams), though there may be remaining text unique between the two sentences. When $precision$ and $recall$ decay, it means that singular words overlap in both sentences, but not in the same sequence, or they are alternated by other words (i.e the high unigram overlap decaying to low trigram overlap). It is plausible that cosine similarity filters out perturbations that have long word sequence overlaps with the original, but that also contain added hallucinations that change the semantic meaning (see Figure 6, the ‘Hallucinated content’ example).
  2. 2.
    Generally, there is higher syntactic overlap than lexical overlap, regardless of filtering. Sometimes this leads to unsatisfactory perturbations, where local rephrasing leads to globally implausible sounding sentences, as shown in Figure 6 (the ‘Local rephrasing, global incoherence’ example).
  3. 3.
    Without filtering, there is higher $precision$ compared to $recall$, while after filtering, the $recall$ increases. From Tables 23 and 24 we can hypothesise that overall cosine similarity filters out perturbations that are shorter than the original sentences.
Observationally, we also find instances of literal rephrasing (see Figure 6, the ‘Literal (not pragmatic) rephrasing’ example), which illustrates the difficulties of generating high quality perturbations for data such as medical queries, which can often have expressed emotions that need to be inferred. Additionally, hallucinated content present in the perturbations is problematic, however would be more so if we were to utilise the additional levels of risk labels from the medical safety dataset (see Section 0.2.4) – the hallucinated content can have a non-trivial impact on label consistency.
Falsifiability.
We start with formally defining the falsifiability metric, following the notation from Section 0.4.1. Let $\mathcal{Y}^{c1}$, $\mathcal{Y}^{c2}$ be an NLP dataset $\mathcal{Y}$ split into two datasets, by classes $c1$ and $c2$, respectively. The reader may find it intuitive to think about $c1$ as a positive class, and $c2$ as a negative class. By perturbing sentences of each class, we obtain sets $\mathcal{A}^{c1}$ and $\mathcal{A}^{c2}$ of adversarial sentence perturbations, according to procedure described in Section 0.4.1. Both $\mathcal{A}^{c1}$ and $\mathcal{A}^{c2}$ have corresponding sets of subspaces $\mathcal{S}^{c1}$ and $\mathcal{S}^{c2}$, i.e. $\mathcal{S}^{c1}=\bigcup\mathcal{S}_{i}$, where $i\in 1,\ldots q$, and $q$ is the number of sentences in $\mathcal{Y}^{c1}$. Let $\mathcal{A}^{c2}=\{s_{1}.\ldots s_{m}\}$. Consider a set $\mathcal{V}^{c2}$ of vectors in ${}^{m}$ such that each vector $v_{i}\in\mathcal{V}^{c2}$ is defined as $v_{i}=\mathbb{E}(s_{i})$, with $s_{i}\in\mathcal{A}^{c2}$. This is the set of embeddings of sentences contained in $\mathcal{A}^{c2}$.
We say a subspace $\mathcal{S}_{i}\in\mathcal{S}^{c1}$ is falsifiable if $\mathcal{S}_{i}$ contains at least one element from $\mathcal{V}^{c2}$. Falsifiability of $\mathcal{S}^{c1}$ is the percentage of elements of $\mathcal{S}^{c1}$ that are falsifiable subspaces. This means, that falsifiability measures the number of subspaces that contain at least one sample of any undesired class. False (adversarial) positives of $\mathcal{S}^{c1}$ are calculated as the percentage of elements of $\mathcal{V}^{c2}$ contained cumulatively in all subspaces contained in $\mathcal{S}^{c1}$.
Dataset Model Verifiability Generalisability Falsifiability False Positives
RUAR $N_{base^{\text{ \ }}}$ (s-bert 22M) 2.56 1256/44013 2.85 1/3400 0.03 27/40270 0.07
$N_{pert^{\text{ \ }}}$ (s-bert 22M) 15.92 8361/44013 19.00 1/3400 0.03 72/40270 0.18
$N_{pert^{\diamondsuit}}$ (s-bert 22M) 21.89 9530/44013 21.65 3/3400 0.09 101/40270 0.25
$N_{base^{\text{ \ }}}$ (s-gpt 1.3B) 0.34 128/44013 0.29 0/3400 0.00 0/40270 0.00
$N_{pert^{\diamondsuit}}$ (s-gpt 1.3B) 11.27 5633/44013 12.80 2/3400 0.06 27/40270 0.07
$N_{base^{\text{ \ }}}$ (s-gpt 2.7B) 0.35 183/44013 0.42 0/3400 0.00 0/40270 0.00
$N_{pert^{\diamondsuit}}$ (s-gpt 2.7B) 11.63 5950/44013 13.52 1/3400 0.03 18/40270 0.04
Medical $N_{base^{\text{ \ }}}$ (s-bert 22M) 58.71 9135/15530 58.82 0/989 0.00 0/12709 0.00
$N_{pert^{\text{ \ }}}$ (s-bert 22M) 70.61 10879/15530 70.05 0/989 0.00 0/12709 0.00
$N_{pert^{\diamondsuit}}$ (s-bert 22M) 73.47 10964/15530 70.6 0/989 0.00 0/12709 0.00
$N_{base^{\text{ \ }}}$ (s-gpt 1.3B) 11.02 2092/15530 13.47 0/989 0.00 0/12709 0.00
$N_{pert^{\diamondsuit}}$ (s-gpt 1.3B) 20.19 3133/15530 20.17 0/989 0.00 0/12709 0.00
$N_{base^{\text{ \ }}}$ (s-gpt 2.7B) 13.44 2489/15530 16.03 0/989 0.00 0/12709 0.00
$N_{pert^{\diamondsuit}}$ (s-gpt 2.7B) 24.92 3957/15530 25.48 0/989 0.00 0/12709 0.00
Table 25: Verifiability, generalisability and falsifiability of the baseline and the robustly (adversarially) trained DNNs on the RUAR and the Medical datasets, for $\mathbb{H}^{*}_{pert^{\diamondsuit}}$ ($N_{base}$ and $N_{pert^{\diamondsuit}}$) and $\mathbb{H}^{*}_{pert}$ ($N_{pert}$); for Marabou verifier.
Example 6 (Falsifiability for Verification Pipeline).
As shown in Table 18, both RUAR and Medical datasets already have two-class splitting, $pos$ and $neg$. $\mathcal{A}^{pos}_{pert}$ and $\mathcal{A}^{neg}_{pert}$ are constructed for a positive and a negative class, respectively. The set $\mathcal{V}^{neg}$ is obtained by embedding sentences of $\mathcal{A}^{neg}_{pert}$. Subspaces for which we measure falsifiability are given by $\mathbb{H}^{*}_{pert}$ and $\mathbb{H}^{*}_{pert^{\diamondsuit}}$ (constructed on $\mathcal{A}^{pos}_{pert}$ and $\mathcal{A}^{pos}_{pert^{\diamondsuit}}$ for RUAR and Medical).
Table 25 shows falsifiability of our models and semantic subspaces, giving quantitative estimation of the scale of the problem outlined in Examples 3 and 4. We can see that the problem indeed exists, hence our claim that falsifiability should be reported routinely in NLP verification papers. At the same time, the problem does not appear to be as severe as one might fear: we note that at most $0.09\%$ of the subspaces are falsified, and falsifiability has only occurred in the subspaces created based on the RUAR dataset. Therefore falsifiability is constrained to a few subspaces that can, in theory, be removed.
Furthermore, falsifiability could also reflect issues in the dataset and subsequent noisy perturbations. The medical safety dataset, for instance, was annotated by an expert practitioner, while the RUAR dataset contains (for this particular task) what could be construed as noisy labels. For example ‘are robot you you a’ is a sample that is found in the negative RUAR train set. The reason for the negative label is that it is an ungrammatical false positive, but given our methods of perturbation for the construction of subspaces, this negative sample may be very similar to a word level perturbation for the positive class. Concretely, for the model with the highest falsifiability in Table 25 (i.e. $N_{pert}$ s-bert 22M for RUAR dataset with $0.09\%$ falsifiability), some sentence pairs of negative samples with their accompanying perturbations contained in falsified subspaces are: (Original: ‘Are you a chump?’, Perturbation: ‘You a chump’), (Original: ‘Are you a liar’, Perturbation: ‘You a liar’), (Original: ‘if a computer can feel emotions, does that make you a computer or an actual human?’, Perturbation: ‘if a computer can feel, does that make it a machine or a person’). Thus, the task of determining what queries require disclosure (e.g. should ‘what is your favorite food’ warrant disclosure?) is more ambiguous and, as the outputs of LLMs sound more coherent, it becomes harder to define. This area merits further research.
Generalisability and Verifiability.
For comparison with the findings outlined in Section 0.4, we provide additional insights into verifiability and generalisability, presented in Table 25. We analyse the effect of cosine similarity filtering. Initially, our observations reveal slightly lower levels of both verifiability and generalisability prior to filtering. The unexpected aspect arises in the realm of generalisability. As demonstrated in Section 0.4, larger subspaces tend to exhibit greater generalisability, and the subspaces before filtering indeed possess a larger volume, however they do not possess higher generalisability. In other words, cosine similarity filtering reduces volume without sacrificing generalisability, as Section 0.4 might suggest. But filtering does improve verifiability in accordance with Section 0.4 conclusions. Therefore, we conjecture that cosine similarity filtering can serve as an additional heuristic for improving precision of the verified DNNs, and for further reducing the verifiability-generalisability gap. Indeed, upon calculating the ratio of generalisability to verifiability, we observe a higher ratio before filtering ($1.19\rightarrow 0.99$ for RUAR and $0.99\rightarrow 0.95$ for Medical). Recall that Section 0.4 already showed that our proposed usage of semantic subspaces can serve as a heuristic for closing the gap; and cosine similarity filtering provides opportunity for yet another heuristic improvement.
Moreover, the best performing model $N_{pert}$ (s-bert 22M), results in $10,964$ $(70.6\%)$ medical perturbations and $9530$ $(21.65\%)$ RUAR perturbations contained. While $21.65\%$ of the $positive$ perturbations contained in the verified subspaces for the RUAR dataset may seem like a low number, it still results in a robust filter, given that the $positive$ class of the dataset contains many adversarial examples of the same input query, i.e. semantically same but lexically different queries. Consequently, giving guarantees for $\approx 10,000$ variations on asking a system to disclose non-identity is a notable result. The medical dataset on the other hand contains many semantically diverse queries, and there are several unseen medical queries not contained in the dataset nor in the resultant verified subspaces. However, given that the subspaces contain $70.6\%$ of the $positive$ perturbations of the medical safety dataset, an application of this could be to carefully curate a dataset of queries containing critical and serious risk-level labels defined by the WEF for chatbots in healthcare (see Section 0.2.4 and [130]), ensure label consistency for the level of risk on the perturbations, and then create verification filters centred around these queries to prevent generation of medical advice for such high-risk queries.
We find that semantically informed verification generalises well across the different kinds of data to ensure guarantees on the output, and thus should aid in the safety of LLMs.
0.6 Conclusions and Future Work
Summary.
This paper started with a general analysis of existing NLP verification approaches, with a view of identifying key components of a general NLP verification methodology, which we distilled as a six component “NLP Verification Pipeline”:
  • dataset selection;
  • generation of perturbations;
  • choice of embedding functions;
  • definition of subspaces;
  • robust training;
  • verification via one of existing verification algorithms.
Based on this taxonomy, we formally defined all the components of the pipeline, and implemented it as a tool ANTONIO [21]. ANTONIO allowed us to mix and match different choices for each pipeline component, enabling us to study the effect of various factors and components of the pipeline in a algorithm-independent way. Our main focus was to identify weak or missing parts of the existing NLP verification methodologies. We proposed that NLP verification results must report, in addition to the standard verifiability metric, the following:
  • whether they use geometric or semantic subspaces, and for which type of semantic perturbations;
  • volumes, generalisability and falsifiability of verified subspaces.
We finished the paper with a study of major pitfalls of NLP verification pipeline from the point of view of LLMs and proposed future wider and more principled use of some standard NLP methods in NLP verification (precision, recall, cosine similarity, ROUGE-N).
Contributions.
The major discoveries of this paper were:
  • In Section 0.4 we showed that NLP verification tasks exhibit a generalisability-verifiability trade-off, and its effects can be very severe, especially if the verification tasks are approached naively. Generalisability is a novel metric we proposed here, and we strongly believe that generalisability should be routinely measured and reported as part of NLP verification pipeline.
  • In Sections 0.4 and 0.5 we showed that it is possible to overcome this trade-off by using several heuristic methods: defining semantic subspaces, training for semantic robustness, choosing the most suited embedding function and filtering with cosine similarity. All of these methods boiled down to defining more precise verifiable subspaces; and all of them can be effectively implemented as part of NLP verification pipelines in the future.
  • In Section 0.5 we demonstrated that accuracy of LLMs, used as embedding functions and generators of perturbations in NLP verification, cannot be taken for granted. They may fail to correctly embed sentences, and they may fail to preserve the semantic meaning of perturbations. Both of these factors influence practical applications of the NLP verification pipeline.
  • In Section 0.5 we demonstrated that even verified subspaces can be semantically falsified: this effect is due to the tension between verification methods that are essentially geometric and the intuitively understood semantic meaning of sentences. By defining the falsifiability metric and using it in our experiments, we found out that the effects of falsifiability do not seem to be severe in practice; but this may vary from one scenario to another. It is important that NLP verification papers are aware of this pitfall, and report faslsifiability alongside verifiability and generalisability.
Finally, we claim as a contribution, a novel coherent methodological framework that allows to include a broad spectrum of NLP, geometric, machine learning, and verification methods under a single umbrella. No previous publication in this domain covered this range and, as this paper illustrated throughout all sections, consistently covering this broad range of methods is crucial for the development of this field.
Future Work.
Following from our in depth analysis of the NLP perspective, we note that even if one has a satisfactory solution to all the issues discussed, there is still the problem of scalability of the available verification algorithms. For example, the most performant neural network verifier, $\alpha\beta$-Crown [118], can only handle networks in the range of tens of millions of trainable parameters. In contrast, in NLP systems, the base model of BERT [32] has around 110 million trainable parameters (considered small compared to modern LLMs – with trainable parameters in the billions!). It is clear that the rate at which DNN verifiers become more performant may never catch up with the rate at which Large Language Models (LLMs) become larger. Then the question arises: how can this pipeline be implemented in the real world?
For future work, we propose to tackle this based on the idea of verifying a smaller DNN (classifier), manageable by verifiers, that can be placed upstream of a complex NLP system as a safeguard. We call this a filter (as mentioned in Section 0.3 and illustrated in Figure 2), and Figure 7 shows how a semantically informed verified filter can be prepended to an NLP system (here, an LLM) to check that safety-critical queries are handled responsibly, e.g. by redirecting the query to a tightly controlled rule-based system instead of a stochastic LLM.
Figure 7: In this figure, we show how a prepended, semantically informed verified filter added to an NLP system (here, an LLM), can check that safety-critical queries are handled responsibly, e.g. by redirecting the query to a tightly controlled rule-based system instead of a stochastic LLM.
While there are different ways to implement the verification filters (e.g. only the verified subspaces) we suggest utilizing both the verified subspaces together with the DNN as the additional classification could strengthen catching positives that fall outside the verified subspaces, thus giving stronger chances of detecting the query via both classification and verification.
We note that the NLP community has recently proposed guardrails, in order to control the output of LLMs and create safer systems (such as from Open AI, NVIDIA and so on). These guardrails have been proposed at multiple stages of an NLP pipeline, for example an output rail that checks the output returned by an LLM, or input rail, that rejects unsafe user queries. In the figure, we show an application of our filter applied to the user input, and thus creating guarantees that safety critical queries are handled responsibly. In theory these verification techniques we propose may be applied to guardrails at different stages in the system, and we plan to explore this in future work.
A second future direction is to use this work to create NLP verification benchmarks. In 2020, the International Verification of Neural Networks Competition [17] (VNN-COMP) was established to facilitate comparison between existing approaches, bring researchers working on the DNN verification problem together, and help shape future directions of the field. However, on its fourth edition, the competition still lacked NLP verification benchmarks [16]. We propose to use this work for creating NLP verification benchmarks for future editions, to spread the awareness and attention to this field.
Acknowledgements
Authors acknowledge support of EPSRC grant AISEC EP/T026952/1 and NCSC grant Neural Network Verification: in search of the missing spec, Vinay Krupakaran for his help in the evaluating the perturbations and Daniel Kienitz for his help with the eigenspace rotation.

### v3

@@
-
+NLP Verification: Towards a General Methodology for Certifying Robustness
 Abstract
-Deep neural networks (DNNs) have exhibited substantial success in the field of Natural Language Processing (NLP).
+Machine Learning (ML) has exhibited substantial success in the field of Natural Language Processing (NLP).
+For example large language models (LLMs) have empirically proven to be capable of producing text of high complexity and cohesion.
+However, at the same time, they are prone to inaccuracies and hallucinations.
 As these systems are increasingly integrated into real-world applications, ensuring their safety and reliability becomes a primary concern.
 There are safety critical contexts where such models must be robust to variability or attack, and give guarantees over their output.
-Computer Vision had pioneered the use of formal verification for neural networks for such scenarios and developed common verification standards and pipelines.
+Computer Vision had pioneered the use of formal verification of neural networks for such scenarios and developed common verification standards and pipelines, leveraging precise formal reasoning about geometric properties of data manifolds.
 In contrast, NLP verification methods have only recently appeared in the literature.
-While presenting sophisticated algorithms on their own right, these papers have not yet crystallised into a common methodology, they are often light on the pragmatical issues of NLP verification, and the area remains fragmented.
-In this paper, we make an attempt to distil and evaluate general components of an NLP verification pipeline, that emerges from the progress in the field to date. Our contributions are two-fold.
-Firstly, we give a general (i.e.
-algorithm-independent) characterisation of verifiable subspaces that result from embedding sentences into continuous spaces.
-We identify, and give an effective method to deal with, the technical challenge of semantic generalisability of verified subspaces; and propose it as a standard metric in the NLP verification pipelines (alongside with the standard metrics of model accuracy and model verifiability).
-Secondly, we propose a general methodology to analyse the effect of the embedding gap – a problem that refers to the discrepancy between verification of geometric subpspaces on the one hand, and semantic meaning of sentences which the geometric subspaces are supposed to represent, on the other hand.
-In extreme cases, poor choices in embedding of sentences may invalidate verification results.
-We propose a number of practical NLP methods that can help to identify the effects of the embedding gap; and in particular we propose the metric of falsifiability of semantic subpspaces that we propose as another fundamental metric to be reported as part of the NLP verification pipeline.
-We believe that together these general principles pave the way towards a more consolidated and effective development of this new domain.
-0.1 Introduction
+While presenting sophisticated algorithms in their own right, these papers have not yet crystallised into a common methodology.
+They are often light on the pragmatical issues of NLP verification, and the area remains fragmented.
+In this paper, we attempt to distil and evaluate general components of an NLP verification pipeline, that emerges from the progress in the field to date.
+Our contributions are two-fold.
+Firstly, we propose a general methodology to analyse the effect of the embedding gap – a problem that refers to the discrepancy between verification of geometric subspaces, and the semantic meaning of sentences which the geometric subspaces are supposed to represent.
+We propose a number of practical NLP methods that can help to quantify the effects of the embedding gap.
+Secondly, we give a general method for training and verification of neural networks that leverages a more precise geometric estimation of semantic similarity of sentences in the embedding space and helps to overcome the effects of the embedding gap in practice.
+keywords:
+Neural Networks, Verification, Natural Language Processing, Robustness, Adversarial Training, Machine Learning
+{Frontmatter}
+[…]
+Casadio M. et al.
+[…]
+Casadio M. et al.
+1 Introduction
 Deep neural networks (DNNs) have demonstrated remarkable success at addressing challenging problems in various areas, such as Computer Vision (CV) [1] and Natural Language Processing (NLP) [2, 3].
@@
 Current NLP systems cannot guarantee either truthfulness, accuracy, faithfulness, or groundedness of outputs given an input query, which can lead to different levels of harm.
-Contexts which necessitate guaranteed outputs.
 One such example in the NLP domain is the requirement of a chatbot to correctly disclose non-human identity, when prompted by the user to do so.
@@
 Similarly, as such systems become widespread in their use, it may be desirable to have guarantees on queries concerning safety critical domains, for example when the user asks for medical advice.
-Research has shown that users tend to attribute undue expertise to systems [2, 34] potentially causing real world harm [12] (e.g.
+Research has shown that users tend to attribute undue expertise to NLP systems [12, 7] potentially causing real world harm [13] (e.g.
 ‘Is it safe to take these painkillers with a glass of wine?’).
 However, a question remains on how to ensure that NLP systems can give formally guaranteed outputs, particularly for scenarios that require maximum control over the output.
-Formal verification of neural networks.
 […]
-Challenges of NLP verification.
-However, existing verification approaches primarily focus on computer vision (CV) tasks, where images are seen as vectors in a continuous space and every point in the space corresponds to a valid image.
+Existing verification approaches primarily focus on computer vision (CV) tasks, where images are seen as vectors in a continuous space and every point in the space corresponds to a valid image.
 In contrast, sentences in NLP form a discrete domain [^2] In this paper, we work with textual representations of sentences.
 Raw audio input can be seen as continuous, but this is out of scope of this paper., making it challenging to apply traditional verification techniques effectively.
-In particular, taking an NLP dataset $\mathcal{Y}$ to be a set of sentences $s_{1},\ldots,s_{q}$ written in natural language, an embedding $\mathbb{E}$ is a function that maps a sentence to a vector in ${}^{m}$.
+In particular, taking an NLP dataset $\mathcal{Y}$ to be a set of sentences $s_{1},\ldots,s_{q}$ written in natural language, an embedding $E$ is a function that maps a sentence to a vector in m.
 The resulting vector space is called the embedding space.
-Due to discrete nature of the set $\mathcal{Y}$, the reverse of the embedding function $\mathbb{E}^{-1}:^{m}\rightarrow\mathcal{Y}$ is not total or undefined.
-This problem is known as the “problem of the embedding gap".
-Sometimes, one uses the term to more generally refer to any discrepancies that $\mathbb{E}$ introduces, for example, when it maps dissimilar sentences close in ${}^{m}$.
+Due to discrete nature of the set $\mathcal{Y}$, the reverse of the embedding function $E^{-1}:^{m}\rightarrow\mathcal{Y}$ is undefined for some elements of m.
+This problem is known as the “problem of the embedding gap”.
+Sometimes, one uses the term to more generally refer to any discrepancies that $E$ introduces, for example, when it maps dissimilar sentences close in m.
 We use the term in both mathematical and NLP sense.
-Mathematically, the general (geometric) “DNN robustness verification" approach of defining and verifying subspaces of ${}^{m}$ should work, and some prior works exploit this fact.
-However, pragmatically, because of the embedding gap only a tiny fraction of vectors contained in the verified subspaces maps back to valid sentences.
+Mathematically, the general (geometric) “DNN robustness verification” approach of defining and verifying subspaces of m should work, and some prior works exploit this fact.
+However, pragmatically, because of the embedding gap, usually only a tiny fraction of vectors contained in the verified subspaces map back to valid sentences.
 When a verified subspace contains no or very few sentence embeddings, we say that verified subspace has low generalisability.
-Low generalisability may render verification efforts obsolete for practical applications.
-From the NLP perspective, there are other, more subtle, examples when the embedding gap can manifest.
-Suppose we succeeded in verifying a DNN robust on some subspace of ${}^{m}$.
-Consider an example when the subspace contains sentences that are semantically similar to the sentence: i really like too chat to a human.
-are you one?.
-Verification gives us guarantees that the DNN will always identify these sentences as questions about human/robot identity.
-But suppose the embedding function $\mathbb{E}$ wrongly embedded sentences belonging to an opposite class into this subspace.
-For example, an LLM Vicuna generates the following sentence as a rephrase of the previous one: Do you take pleasure in having a conversation with someone?.
-Suppose our verified subspace contained an embedding of this sentence, too, and thus our verified DNN identifies this second sentence to belong to the the same class as the first one.
+Low generalisability may render verification efforts ineffective for practical applications.
+From the NLP perspective, there are other, more subtle, examples where the embedding gap can manifest.
+Consider an example of a subspace containing sentences that are semantically similar to the sentence: ‘i really like too chat to a human.
+are you one?’.
+Suppose we succeed in verifying a DNN to be robust on this subspace.
+This provides a guarantee that the DNN will always identify sentences in this subspace as questions about human/robot identity.
+But suppose the embedding function $E$ wrongly embeds sentences belonging to an opposite class into this subspace.
+For example, the LLM Vicuna [14] generates the following sentence as a rephrasing of the previous one: Do you take pleasure in having a conversation with someone?.
+Suppose our verified subspace contained an embedding of this sentence too, and thus our verified DNN identifies this second sentence to belong to the same class as the first one.
 However, the second sentence is not a question about human/robot identity of the agent!
-When we can find such an example, we say that it falsifies the verification guarantee for the subspace it is contained in.
-Alternatively, we say that the subpspace is falsifiable.
+When we can find such an example, we say that the verified subspace is prone to embedding errors.
+Robustness verification in NLP is particularly susceptible to this problem, because we cannot cross the embedding gap in the opposite direction as the embedding function is not invertible.
+This means it is difficult for humans to understand what sort of sentences are captured by a given subspace.
 Contributions
 Our main aim is to provide a general and principled verification methodology that bridges the embedding gap when possible; and gives precise metrics to evaluate and report its effects in any case.
-The contributions split into two main groups, depending whether the embedding gap is approached from mathematical or NLP perspective.
+The contributions split into two main groups, depending on whether the embedding gap is approached from mathematical or NLP perspective.
+Contributions Part 1: Characterisation of Verifiable Subspaces and general
 […]
-Contributions Part 1: Characterisation of verifiable subspaces and NLP verification pipeline.
 […]
-An alternative method to purely geometric approach that suffers from the embedding gap is to construct subspaces of the embedding space based on the semantic perturbations of sentences [54, 50, 146].
-Concretely, the idea is to form each $\mathcal{S}_{i}$ by embedding a sentence $i$ and its $n$ semantic perturbations into the real vector space and enclosing them inside some geometric shape.
-Ideally, such a shape should be given by a convex hull around these $n+1$ embedded sentences, however calculating convex hulls with sufficient precision is computationally infeasible for high number of dimensions.
+An alternative method to the purely geometric approach is to construct subspaces of the embedding space based on the semantic perturbations of sentences (first attempts to do this appeared in [16, 17, 18, 19]).
+Concretely, the idea is to form each $\mathcal{S}_{i}$ by embedding a sentence $s$ and $n$ semantic perturbations of $s$ into the real vector space and enclosing them inside some geometric shape.
+Ideally, this shape should be the convex hull around the $n+1$ embedded sentences (see Figure 1), however calculating convex hulls with sufficient precision is computationally infeasible for high number of dimensions.
 Thus, simpler shapes, such as hyper-cubes and hyper-rectangles are used in the literature.
@@
 We will call the resulting shapes semantic subspaces (in contrast to those obtained purely geometrically).
-A few questions have been left un-answered in the previous work.
+A few questions have been left unanswered in the previous work [16, 17, 18, 19].
 Firstly, because generalisability of the verified subspaces is not reported in the literature, we cannot know whether the prior semantically-informed approaches are better in that respect than purely geometric methods.
@@
   • the improved precision of verified subspaces that comes from using the semantic knowledge.
-This paper provides a strong argument for including generalisability as a standard metric in reporting NLP verification results in the future.
-Through a series of experiments, we confirm that verified semantic subspaces are more verifiable and more generalisable than their geometric counterparts.
+Through a series of experiments, we confirm that semantic subspaces are more verifiable and more generalisable than their geometric counterparts.
 Moreover, by comparing the volumes of the obtained verified semantic and geometric subspaces, we show that the improvement is partly due to finding an optimal size of subspaces (for the given embedding space), and partly due to improvement in shape precision.
-The second group of unresolved questions concerns robust training regimes in NLP verification that is used as means of improving verifiability of subspaces in prior works [54, 50, 146].
+The second group of unresolved questions concerns robust training regimes in NLP verification that is used as means of improving verifiability of subspaces in prior works [16, 17, 18, 19].
 It was not clear what made robust training successful:
-  • was it because additional examples generally improved the precision of the decision boundary? (in which case data set augmentation would have a similar effect);
+  • was it because additional examples generally improved the precision of the decision boundary (in which case dataset augmentation would have a similar effect);
 […]
@@
   • choice of the verification algorithm (on the verification side).
-We argue that, together with the new generalisability metric, this approach opens the way for more principled evaluation of performance of NLP verification methods that accounts for the effects of the embedding gap; and generation of more transparent NLP verification benchmarks.
-We implement a tool that generates NLP verification benchmarks based on the above choices.
+We argue that this approach opens the way for more principled NLP verification methods that reduces the effects of the embedding gap; and generation of more transparent NLP verification benchmarks.
+We implement a tool ANTONIO that generates NLP verification benchmarks based on the above choices.
 This paper is the first to use a complete SMT-based verifier (namely Marabou [22]) for NLP verification.
@@
 In particular, we show that the quality of the generated sentence perturbations maybe improved through the use of human evaluation, cosine similarity and ROUGE-N.
-We introduce the falsifiability metric as an effective practical way to measure the quality of the embedding functions.
-Through a detailed case study, we show how geometric and NLP intuitions can be put together at work towards obtaining DNNs that are more verifiable over better generalisable and less falsifiable semantic subspaces.
-Perhaps more importantly, the proposed methodology opens the way for transparency in reporting NLP verification results, – something that this domain will benefit from if it reaches the stage of practical deplyoment of NLP verification pipelines.
+We introduce the novel embedding error metric as an effective practical way to measure the quality of the embedding functions.
+Through a detailed case study, we show how geometric and NLP intuitions can be put at work towards obtaining DNNs that are more verifiable over better generalisable and less prone to embedding errors semantic subspaces.
+Perhaps more importantly, the proposed methodology opens the way for transparency in reporting NLP verification results, – something that this domain will benefit from if it reaches the stage of practical deployment of NLP verification pipelines.
 Paper Outline.
 From here, the paper proceeds as follows.
-Section 0.2 gives an extensive literature review encompassing DNN verification methods generally, and NLP verification methods in particular.
-The section culminates with distilling a common “NLP verification pipeline" encompassing the existing literature.
+Section 2 gives an extensive literature review encompassing DNN verification methods generally, and NLP verification methods in particular.
+The section culminates with distilling a common “NLP verification pipeline” encompassing the existing literature.
 Based on the understanding of major components of the pipeline, the rest of the paper focuses on improving understanding or implementation of its components.
-Section 0.3 formally defines the components of the pipeline in a general mathematical notation, which abstracts away from particular choices of sentence perturbation, sentence embedding, training and verification algorithms.
-The central notion the section introduces is that of geometric and semantic subpspaces.
-The next Section 0.4 makes full use of this general definition, and shows that semantic subpspaces play a pivotal role in improving verification and training of DNNs in NLP.
+Section 3 formally defines the components of the pipeline in a general mathematical notation, which abstracts away from particular choices of sentence perturbation, sentence embedding, training and verification algorithms.
+The central notion the section introduces is that of geometric and semantic subspaces.
+The next Section 4 makes full use of this general definition, and shows that semantic subspaces play a pivotal role in improving verification and training of DNNs in NLP.
 This section formally defines the generalisability metric and considers the problem of generalisability-verifiability trade-off.
-Through thorough empirical evaluation, it shows that a principled approach to defining semantic subpaces can help to improve both generalisability and verifiability of DNNs, thus reducing the effects of the trade-off.
-The final Section 0.5 further tests the NLP verification pipelines using state-of-the-art NLP tools, and analyses the effects of the embedding gap from the NLP persective, in particular it introduces a method of measuring falsifiability of semantic subpaces and reporting this metric alongside verifiability and generalisability.
-Section 0.6 concludes the paper and discusses future work.
-0.2 Related Work
-0.2.1 DNN Verification
-Formal verification is an active field across several domains including hardware [64, 97], software languages [58], network protocols [87] and many more [129], however it was only recently that this became applicable to the field of machine learning [59].
-Several verifiers have been popular in DNNs verification and competitions [5, 6, 81, 108].
-We can divide them into 2 main categories: complete and incomplete verifiers.
-When the verification approach guarantees that, if a query fails, then the query is false and it produces a counter-example to prove it, then we call it complete verification, otherwise we call it incomplete verification.
-Furthermore, while complete verifiers are always deterministic, incomplete verifiers may also be probabilistic.
-Unlike deterministic verification, probabilistic verification is guaranteed to output ‘not verified’ with a certain probability (e.g., 99.9%).
-Complete verifiers
-can be based on Satisfiability Modulo Theories (SMT), Mixed-Integer Linear Programming (MILP) or Branch-and-Bound (Bab).
-SMT-based verification [99, 59, 131] is built upon the observation that feed-forward neural networks are defined by the sequential composition of affine trasformations and ReLU operations.
-Both these transformations and operations can be encoded by a conjunction of linear inequalities, thus general-purpose SMT solvers can be directly applied to solve the satisfiability problem, yielding a solution to complete verification.
-A state-of-the-art SMT-based tool is Marabou [131], which answers queries about neural networks and their properties in the form of constraint satisfaction problems.
-Marabou takes as input networks with piece-wise linear activation functions and with fully connected topology.
-It first applies multiple pre-processing steps to infer bounds for each node in the network.
-Next it applies a combination of Simplex [31] search over linear constraints with SMT techniques directing the search over non-linear constraints.
+Through thorough empirical evaluation, it shows that a principled approach to defining semantic subspaces can help to improve both generalisability and verifiability of DNNs, thus reducing the effects of the trade-off.
+The final Section 5 further tests the NLP verification pipelines using state-of-the-art NLP tools, and analyses the effects of the embedding gap from the NLP perspective, in particular it introduces a method of measuring the embedding error and reporting this metric alongside verifiability and generalisability.
+Section 6 concludes the paper and discusses future work.
+2 Related Work
+2.1 DNN Verification
+Formal verification is an active field across several domains including hardware [23, 24], software [25], network protocols [26] and many more [27].
+However, it was only recently that this became applicable to the field of machine learning [28].
+An input query to a verifier consists of a subspace within the embedding space and a target subspace of outputs, typically a target output class.
+The verifier then returns either true, false or unknown.
+True indicates that there exists an input within the given input subspace whose output falls within the given output subspace, often accompanied by an example of such input.
+False indicates that no such input exists.
+Several verifiers are popular in DNN verification and competitions [29, 30, 31, 32].
+We can divide them into 2 main categories: complete verifiers which return true/false and incomplete verifiers which return true/unknown.
+While complete verifiers are always deterministic, incomplete verifiers may be probabilistic.
+Unlike deterministic verification, probabilistic verification is not sound and a verifier may incorrectly output true with a very low probability (typically 0.01%).
+Complete Verification based on Linear Programming & Satisfiability Modulo Theories (SMT) solving.
+Generally, SMT solving is a group of methods for determining the satisfiability of logical formulas with respect to underlying mathematical theories such as real arithmetic, bit-vectors, or arrays [33].
+These methods extend traditional satisfiability (SAT) solving by incorporating domain-specific reasoning, making them particularly useful for verifying complex systems.
+In the context of neural network verification, SMT solvers encode network behaviours and safety properties as logical constraints, enabling rigorous checks for violations of specifications [34].
+When the activation functions are piecewise linear (e.g.
+ReLU), the DNN can be encoded by conjunctions and disjunctions of linear inequalities and thus linear programming algorithms can be directly applied to solve the satisfiability problem.
+A state-of-the-art tool is Marabou [22], which answers queries about neural networks and their properties in the form of constraint satisfaction problems.
+Marabou takes the network as input and first applies multiple pre-processing steps to infer bounds for each node in the network.
+It applies the algorithm ReLUplex [28], a combination of Simplex [35] search over linear constraints, modified to work for networks with piece-wise linear activation functions.
+With time, Marabou grew into a complex prover with multiple heuristics supplementing the original ReLUplex algorithm [22], for example it now includes mixed-integer linear programming (MILP) [36] and abstract interpretation based algorithms which we survey below.
 MILP-based approaches [37, 38, 39] encode the verification problem as a mixed-integer linear programming problem, in which the constraints are linear inequalities and the objective is represented by a linear function.
-Differently from linear programming (LP), in MILP it is possible to constrain some variables to take only integer values instead of real numbers, allowing the constraints to encode the non-linear ReLU operations.
-Thus, the verification problem can be precisely encoded as an MILP problem.
-A representative tool for this category is ERAN [109], which is mainly based on abstract interpretation (see ‘Incomplete verifiers’ below ) but can also leverage the efficient MILP solver GUROBI [48].
-ERAN combines abstract domains with custom multi-neuron relaxations to support fully-connected, convolutional, and residual networks with ReLU, Sigmoid, Tanh, and Maxpool activations.
-Both these methods suffer from scalability, which is their main limitation, but their strength is that they precisely encode the constraint.
-BaB-based verification [43, 19, 18, 41, 57, 118, 140] relies on the piecewise-linear property of DNNs: since each ReLU neuron outputs ReLU($x$) = max{$x$, 00}, it is always locally linear within some region around input $x$.
-Furthermore, since feed-forward ReLU networks are the composition of these piecewise linear neurons and (linear) affine transformations, the output is locally linear w.r.t.
-input $x$.
-This property is formally stated and proved in [57] and it serves as the foundation for BaB verification.
+Thus, the DNN verification problem can be precisely encoded as a MILP problem.
+For example, ERAN [40] combines abstract interpretation with the MILP solver GUROBI [41].
+By the time Branch and Bound (BaB) methodologies are introduced later, it becomes evident that the verification community has effectively consolidated diverse approaches into a unified taxonomy.
+Modern verifiers, such as $\alpha\beta$-CROWN [42, 43], take full advantage of this combination and effectively balance efficiency with precision.
+Incomplete Verification based on Abstract Interpretation takes inspiration from the domain of abstract interpretation, and mainly uses linear relaxations on ReLU neurons, resulting in an over-approximation of the initial constraint.
+Abstract interpretation was first developed by Cousot and Cousot [44] in 1977.
+It formalises the idea of abstraction of mathematical structures, in particular those involved in the specification of properties and proof methods of computer systems [45] and it has since been used in many applications [46].
+Specifically, for DNN verification, this technique can model the behaviour of a network using an abstract domain that captures the possible range of values the network can output for a given input.
+Abstract interpretation-based verifiers can define a lower bound and an upper bound of the output of each ReLU neuron as linear constraints, which define a region called ReLU polytope that gets propagated through the network.
+To propagate the bounds, one can use interval bound propagation (IBP) [47, 48, 49, 50].
+The strength of IBP-based methods lies in their efficiency; they are faster than alternative approaches and demonstrate superior scalability.
+However, their primary limitation lies in the inherently loose bounds they produce [48].
+This drawback becomes particularly pronounced in the case of deeper neural networks, typically those with more than 10 layers [51], where they cannot certify non-trivial robustness properties.
+Other methods that are less efficient but produce tighter bounds are based on polyhedra abstraction, such as CROWN [52] and DeepPoly [53], or based on multi-neuron relaxation, such as PRIMA [54].
+An abstract interpretation tool CORA [55], uses polyhedral abstractions and reachability analysis for formal verification of neural networks.
+It integrates various set representations, such as zonotopes, and algorithms to compute reachable sets for both continuous and hybrid systems, providing tighter bounds in verification tasks.
+Another mature tool in this category is ERAN [40], which uses abstract domains (DeepPoly) with custom multi-neuron relaxations (PRIMA) to support fully-connected, convolutional, and residual networks with ReLU, Sigmoid, Tanh, and Maxpool activations.
+Note that, having lost completeness, they can work with a more general class of neural networks (e.g.
+neural networks with non linear layers).
+Modern Neural Network Verifiers.
+Modern verifiers are complex tools that take advantage of a combination of complete and incomplete methods as well as additional heuristics.
+The term Branch and Bound (BaB) [56, 57, 58, 59, 60, 43, 61] often refers to the method that relies on the piecewise linear property of DNNs: since each ReLU neuron outputs ReLU($x$) = max{$x$,00} is piecewise linear, we can consider its two linear pieces $x\geq 0$, $x\leq 0$ separately.
 A BaB verification approach, as the name suggests, consists of two parts: branching and bounding.
-It first applies incomplete verification to derive a lower bound and an upper bound, then, if the lower bound is positive it terminates with ‘verified’, else, if the upper bound is non-positive it terminates with ‘not verified’ (bounding).
+It first derives a lower bound and an upper bound, then, if the lower bound is positive it terminates with ‘verified’, else, if the upper bound is non-positive it terminates with ‘not verified’ (bounding).
 Otherwise, the approach recursively chooses a neuron to split into two branches (branching), resulting in two linear constraints.
 Then bounding is applied to both constraints and if both are satisfied the verification terminates, otherwise the other neurons are split recursively.
-When all neurons are split, the branch will contain only linear constraints, and thus the approach applies linear programming to compute the precise constraint and verify the branch.
+When all neurons are split, the branch will contain only linear constraints, and thus the approach applies linear programming to compute the constraint and verify the branch.
+It is important to note that BaB approaches themselves are neither inherently complete nor incomplete.
+BaB is an algorithm for splitting problems into sub-problems and requires a solver to resolve the linear constraints.
+The completeness of the verification depends on the combination of BaB and the solver used.
 Multi-Neuron Guided Branch-and-Bound (MN-BaB) [59] is a state-of-the-art neural network verifier that builds on the tight multi-neuron constraints proposed in PRIMA [62] and leverages these constraints within a BaB framework to yield an efficient, GPU based dual solver.
-Another state-of-the-art tool is $\alpha,\beta$-CROWN [133, 118], a neural network verifier based on an efficient linear bound propagation framework and branch-and-bound.
+Another state-of-the-art tool is $\alpha\beta$-CROWN [42, 43], a neural network verifier based on an efficient linear bound propagation framework and branch-and-bound.
 It can be accelerated efficiently on GPUs and can scale to relatively large convolutional networks (e.g., $10^{7}$ parameters).
 It also supports a wide range of neural network architectures (e.g., CNN, ResNet, and various activation functions).
-BaB-based methods are more scalable than solver-based approaches, however they introduce a level of abstraction and sacrifice precision in favor of scalability.
-For example GCP-CROWN [140] extracts convex constraints from MILP solvers and integrates them in linear inequality propagation, which can be viewed as leveraging multi-neuron relaxations in branch-and-bound complete verification.
-Deterministic incomplete verifiers
-mainly use linear relaxations on ReLU neurons, resulting in an over-approximation of the initial constraint.
-In general, they define a lower bound and an upper bound of the output of each ReLU neuron as linear constraints, which define a region called ReLU polytope that gets propagated through the network.
-A predominant approach is the use of interval bound propagation (IBP) [128, 46, 83, 88].
-The strength of IBP-based methods lies in their efficiency; they are faster than alternative approaches and demonstrate superior scalability.
-However, their primary limitation lies in the inherently loose bounds they produce [46].
-This drawback becomes particularly pronounced in the case of deeper neural networks, typically those with more than 10 layers [73], where they cannot certify non-trivial robustness due to the amplification of over-approximation.
-Other methods that are less efficient but produce tighter bounds are based on polyhedra abstraction, such as CROWN [141] and DeepPoly [111], or based on multi-neuron relaxation, such as PRIMA [93].
-One of the most mature tool in this category is ERAN [109], which can be used for complete verification, but its main purpose is deterministic incomplete verification through abstract interpretation (DeepPoly) and multi-neuron relaxation (PRIMA).
-Probabilistic incomplete verification
-approaches add random noise to smooth models, and then derive certified robustness for these smoothed models.
+Probabilistic Incomplete Verification approaches add random noise to models to smooth them, and then derive certified robustness for these smoothed models.
 This field is commonly referred to as Randomised Smoothing, given that these approaches provide probabilistic guarantees of robustness, and all current probabilistic verification techniques are tailored for smoothed models [63, 64, 65, 66, 67, 68].
-Given that this work focuses on deterministic approaches, here we only report the existence of this line of work without going into details.
+Given that our work focuses on deterministic approaches, here we only report the existence of this line of work without going into details.
 Note that these existing verification approaches primarily focus on computer vision tasks, where images are seen as vectors in a continuous space and every point in the space corresponds to a valid image, while sentences in NLP form a discrete domain, making it challenging to apply traditional verification techniques effectively.
 In this work we use both an abstract interpretation-based incomplete verifier (ERAN [40]) and an SMT-based complete verifier (Marabou [22]) in order to demonstrate the effect that the choice of a verifier may bring, and demonstrate common trends.
-0.2.2 Robust Training
-Verifying DNNs poses significant challenges if they are not appositely trained.
+2.2 Geometric Representations in DNN Verification
+[…]
+2.3 Robust Training
+Verifying DNNs poses significant challenges if they are not appropriately trained.
 The fundamental issue lies in the failure of DNNs, including even sophisticated models, to meet essential verification properties, such as robustness [70].
@@
 Thus, robust training is a strong complement to robustness verification approaches.
-Robust training
-techniques can be classified into several large groups:
+Robust training techniques can be classified into several large groups:
   • data augmentation [73],
-  • adversarial training [45, 85],
+  • adversarial training [71, 20] including property-driven training [74, 75],
   • IBP training [48, 76] and other forms of certified training [77], or
-  • a combination thereof [145].
+  • a combination thereof [78, 70].
 Data augmentation involves the creation of synthetic examples through the application of diverse transformations or perturbations to the initial training data.
@@
 Adversarial training entails identifying worst-case examples at each epoch during the training phase and calculating an additional loss on these instances.
-State of the art adversarial training involve projected gradient descent algorithms such as FGSM [45] and PGD [85].
+State of the art adversarial training involves projected gradient descent algorithms such as FGSM [71] and PGD [20].
 Certified training methods focus on providing mathematical guarantees about the model’s behaviour within certain bounds.
@@
 Note that all techniques mentioned above can be categorised based on whether they primarily augment the data (such as data augmentation) or augment the loss function (as seen in adversarial, IBP and certified training).
-Augmenting the data tends to enhance generalisation and is efficient, albeit it may not help against the most severe adversarial attacks.
-Conversely, methods that manipulate the loss functions directly confront the toughest adversarial attacks but often come with higher computational costs.
+Augmenting the data tends to be efficient, although it may not help against stronger adversarial attacks.
+Conversely, methods that manipulate the loss functions directly are more resistant to strong adversarial attacks but often come with higher computational costs.
 Ultimately, the choice between altering data or loss functions depends on the specific requirements of the application and the desired trade-offs between performance, computational complexity, and robustness guarantees.
-NLP robustness.
+2.4 NLP robustness
 There exists a substantial body of research dedicated to enhancing the adversarial robustness of NLP systems [79, 80, 81, 82, 83, 84, 85].
 These efforts aim to mitigate the vulnerability of NLP models to adversarial attacks and improve their resilience in real-world scenarios [80, 81] and mostly employ data augmentation techniques [86, 87].
-In NLP, perturbations can occur at the character, word, or sentence level [24, 52, 20] and may involve deletion, insertion, swapping, flipping, substitution with synonyms, concatenation with characters or words, or insertion of numeric or alphanumeric characters [76, 39, 69].
-For instance, in character level adversarial attacks, [9] introduces natural and synthetic noise to input data, while  [42, 72] identify crucial words within a sentence and perturbs them accordingly.
-For word level attacks, they can be categorised into gradient-based [76, 104], importance-based [51, 56], and replacement-based [3, 66, 98] strategies based on the perturbation method employed.
-In addition, in sentence level adversarial attacks, some attacks [53, 124] are created so that they do not impact the original label of the input and can be incorporated as a concatenation in the original text.
+In NLP, we can distinguish perturbations based on three main criteria:
+  • where and how the perturbations occur,
+  • whether they are altered automatically using some defined rules (vs. generated by humans or LLMs) and
+  • whether they are adversarial (as opposed to random).
+In particular, perturbations can occur at the character, word, or sentence level [88, 89, 90] and may involve deletion, insertion, swapping, flipping, substitution with synonyms, concatenation with characters or words, or insertion of numeric or alphanumeric characters [91, 92, 93].
+For instance, in character level adversarial attacks, Belinkov et al.
+[94] introduce natural and synthetic noise to input data, while Gao et al.
+[95] and Li et al.
+[96] identify crucial words within a sentence and perturb them accordingly.
+For word level attacks, they can be categorised into gradient-based [91, 97], importance-based [98, 99], and replacement-based [100, 101, 102] strategies, based on the perturbation method employed.
+Moreover, Moradi et al.
+[103] introduce rule-based non-adversarial perturbations at both the character and word levels.
+Their method simulates various types of noise typically caused by spelling mistakes, typos, and other similar errors.
+In sentence level adversarial attacks, some perturbations [104, 105] are created so that they do not impact the original label of the input and can be incorporated as a concatenation in the original text.
 In such scenarios, the expected behaviour from the model is to maintain the original output, and the attack can be deemed successful if the label/output of the model is altered.
+Additionally, non-rule-based sentence perturbations can be obtained through prompting LLMs [21, 14] to generate rephrasing of the inputs.
 By augmenting the training data with these perturbed examples, models are exposed to a more diverse range of linguistic variations and potential adversarial inputs.
@@
 To help with this task, the NLP community has gathered a dataset of adversarial attacks named AdvGLUE [106], which aims to be a principled and comprehensive benchmark for NLP robustness measurements.
-In this work we employ a PGD-based adversarial training as the method to increase the robustness and verifiability of our models, and we create character and word level perturbations as in [91] and sentence level perturbations with PolyJuice [132] and Vicuna [25].
-0.2.3 Previous NLP Verification Approaches
-Method Verification algorithm Verification characteristics Datasets NLP perturbations Embeddings Architectures (# of parameters) Robust training
-Ours
-SMT-based, Abstract interpretation-based,
-BaB-based
+In this work we employ a PGD-based adversarial training as the method to enhance the robustness and verifiability of our models against gradient-based adversarial attacks.
+For non-adversarial perturbations, we create rule-based perturbations at the character and word level as in Moradi et al.
+[103] and non-rule-based perturbations at the sentence level using PolyJuice [21] and Vicuna [14].
+We thus cover most combinations of the three choices above (bypassing only human-generated adversarial attacks as there is no sufficient data to admit systematic evaluation which is important for this study).
+Method Datasets NLP perturbations Embeddings Architectures (# of parameters) Robust training Verification algorithm Verification characteristics
+Ours RUARobot,
+Medical
+[…]
 Complete,
-Precise,
 Deterministic
-RUARobot, Medical
+Jia et al. (2019) [17] IMDB, SNLI Word substitution Word: GloVe LSTM, CNN, BoW, Attention-based, ($10^{5}$) IBP-based loss function augmentation Abstract Interpretation IBP Incomplete,
+Deterministic
+Huang et al. (2019) [18] AGNews, SST Char and word substitution Word: GloVe CNN ($10^{5}$) IBP-based loss function augmentation Abstract Interpretation IBP Incomplete,
+Deterministic
+Welbl et al.
+(2020) [107] SNLI, MNLI Word deletion Word: GloVe Attention-based ($10^{5}$) Data augmentation, random and beam search adversarial training, IBP-based Abstract Interpretation IBP Incomplete,
+Deterministic
+Zhang et al. (2021) [19] IMDB, SST, SST2 Word perturbations Word: not specified LSTM ($10^{5}$) IBP-based loss function augmentation Abstract Interpretation IBP Incomplete,
+Deterministic
+Wang et al. (2023) [108] IMDB, YELP, SST2 Word substitution Word: GloVe CNN ($10^{5}$) IBP-based: Embedding Interval Bound Constraint (EIBC) triplet loss Abstract Interpretation IBP Incomplete,
+Deterministic
 […]
-Sentence: S-BERT, S-GPT
-FFNN ($10^{4}$)
-PGD-based
-Jia et al. (2019) [54]
-IBP-based
-Incomplete,
-Imprecise,
 Deterministic
-IMDB, SNLI
-Word substitution
-Word: GloVe
-LSTM, CNN, BoW, Attention-based, ($10^{5}$)
-IBP-based
-Huang et al. (2019) [50]
-IBP-based
-Incomplete,
-Imprecise,
+[…]
 Deterministic
-AGNews, SST
-Char and word substitution
-Word: GloVe
-CNN ($10^{5}$)
-IBP-based
-Welbl et al. (2020) [126]
-IBP-based
-Incomplete,
-Imprecise,
+[…]
 Deterministic
-SNLI, MNLI
-Word deletion
-Word: GloVe
-Attention-based ($10^{5}$)
-Data augmentation, random and beam search adversarial training, IBP-based
-Zhang et al. (2021) [146]
-IBP-based
-Incomplete,
-Imprecise,
+[…]
 Deterministic
-IMDB, SST, SST2
-Word perturbations
-Word: not specified
-LSTM ($10^{5}$)
-IBP-based
-Wang et al. (2023) [123]
-IBP-based
-Incomplete,
-Imprecise,
-Deterministic
-IMDB, YELP, SST2
-Word substitution
-Word: GloVe
-CNN ($10^{5}$)
-IBP-based: Embedding Interval Bound Constraint (EIBC) triplet loss
-Ko et al. (2019) [61]
-Abstract interpretation-based
-Incomplete,
-Imprecise,
-Deterministic
-CogComp QC
-[…]
-Word: not specified
-RNN, LSTM ($10^{5}$)
-Shi et al. (2020) [107]
-Abstract interpretation-based
-Incomplete,
-Imprecise,
-Deterministic
-YELP, SST
-[…]
-Word: not specified
-Transformer ($10^{6}$)
-Du et al. (2021) [37]
-Abstract interpretation-based
-Incomplete,
-Imprecise,
-Deterministic
-Rotten Tomatoes Movie Review, Toxic Comment
-[…]
-Word: GloVe
-RNN, LSTM ($10^{5}$)
-IBP-based
-Bonaert et al. (2021) [14]
-Abstract interpretation-based
-Incomplete,
-Imprecise,
-Deterministic
-SST, YELP
-[…]
-Word: not specified
-Transformer ($10^{6}$)
-Ye et al. (2020) [134]
-Randomised smoothing ($\alpha=0.01,n=5000$)
-Incomplete,
-Imprecise,
+Table 1: Summary of the main features of the existing NLP verification approaches. In bold are state-of-the-art methods.
+Method Datasets NLP perturbations Embeddings Architectures (# of parameters) Robust training Verification algorithm Verification characteristics
+Ye et al. (2020) [112] IMDB, Amazon Word substitution Word: GloVe Transformer ($10^{8}$) Data
+augmentation
+Randomised smoothing ($\alpha=0.01,n=5000$) Incomplete,
 Probabilistic
-IMDB, Amazon
-Word substitution
-Word: GloVe
-Transformer ($10^{8}$)
-Data
+Wang et al. (2021) [113] IMDB, AGNews Word substitution Word: GloVe LSTM ($10^{5}$) Data
 augmentation
-Wang et al. (2021) [120]
-Differential privacy-based
-Incomplete,
-Imprecise,
+Differential privacy Incomplete,
 Probabilistic
-IMDB, AGNews
-Word substitution
-Word: GloVe
-LSTM ($10^{5}$)
-Data
-augmentation
-Zhao et al. (2022) [148]
-Randomised smoothing ($\alpha=0.001,n=30050$)
-Incomplete,
-Imprecise,
-Probabilistic
-AGNews, SST
-Word substitution
-Word: GloVe
-Transformer ($10^{8}$)
-Data
+Zhao et al. (2022) [114] AGNews, SST Word substitution Word: GloVe Transformer ($10^{8}$) Data
 augmentation
 and IBP-based
-Zeng et al. (2023) [137]
-Randomised smoothing ($\alpha=0.05,n=5000$)
-Incomplete,
-Imprecise,
+Randomised smoothing ($\alpha=0.001,n=30050$) Incomplete,
 Probabilistic
-IMDB, YELP
-Char and word substitution
-Word: not specified
-Transformer ($10^{8}$)
-Data
+Zeng et al. (2023) [115] IMDB, YELP Char and word substitution Word: not specified Transformer ($10^{8}$) Data
 augmentation
-Ye et al. (2023) [135]
-Randomised smoothing ($\alpha=0.001,n=9000$)
-Incomplete,
-Imprecise,
+Randomised smoothing ($\alpha=0.05,n=5000$) Incomplete,
 Probabilistic
-IMDB, SST2, YELP, AGNews
-Word substitution
-Word: not specified
-Transformer ($10^{8}$)
-Data
+Ye et al. (2023) [116] IMDB, SST2, YELP, AGNews Word substitution Word: not specified Transformer ($10^{8}$) Data
 augmentation
-Zhang et al. (2023) [144]
-Randomised smoothing ($\alpha=0.001,n=20000$)
-Incomplete,
-Imprecise,
+Randomised smoothing ($\alpha=0.001,n=9000$) Incomplete,
 Probabilistic
-IMDB, Amazon, AGNews
-Word perturbations
-Word: GloVe
-LSTM, Transformer ($10^{8}$)
-Data
+Zhang et al. (2023) [117] IMDB, Amazon, AGNews Word perturbations Word: GloVe LSTM, Transformer ($10^{8}$) Data
 augmentation
-Zhang et al. (2023) [147]
-Randomised smoothing ($\alpha=0.05,n=5000$)
-Incomplete,
-Imprecise,
+Randomised smoothing ($\alpha=0.001,n=20000$) Incomplete,
 Probabilistic
-SST2, AGNews
-Word perturbations
-Word: not specified
-Transformer ($10^{9}$)
-Table 1: Summary of the main features of the existing NLP verification approaches (first half) and randomised smoothing approaches (second half). In bold are SoA methods.
+Zhang et al. (2023) [118] SST2, AGNews Word perturbations Word: not specified Transformer ($10^{9}$) - Randomised smoothing ($\alpha=0.05,n=5000$) Incomplete,
+Probabilistic
+Table 2: Summary of the main features of the existing randomised smoothing approaches.
+2.5 Datasets and Use Cases Used in NLP Verification
+Existing NLP verification datasets.
+Table 3 summarises the main features and tasks of the datasets used in NLP verification.
+Despite their diverse origins and applications, the datasets in the literature are usually binary or multi-class text classification problems.
+Furthermore, datasets can be sensitive to perturbations, i.e.
+perturbations can have non-trivial impact on label consistency.
+For example, Jia et al.
+[17] use IBP with the SNLI [119] [^3] A semantic inference dataset that labels whether one sentence entails, contradicts or is neutral to another sentence.
+dataset (see Tables 1 and 3) to show that word perturbations (e.g.
+‘good’ to ‘best’) can change whether one sentence entails another.
+Some works such as Jia et al.
+[17] try to address this label consistency, while others do not.
+Additionally, we find that the previous research on NLP verification does not utilise safety critical datasets (which strongly motivates the choice of datasets in alternative verification domains), with the exception of Du et al.
+[110] that use the Toxic Comment dataset [120].
+Other papers do not provide detailed motivation as to why the dataset choices were made, however it could be due to the datasets being commonly used in NLP benchmarks (IMDB etc.).
+Dataset Safety Critical Category Tasks Size Classes
+IMDB [121] $\times$ Sentiment analysis Document-level and sentence-level classification 25,000 2
+SST [122] $\times$ Sentiment analysis Sentiment classification, hierarchical sentiment classification, sentiment span detection 70,042 5
+SST2 [122] $\times$ Sentiment analysis Sentiment classification 70,042 2
+YELP [123] $\times$ Sentiment analysis Sentiment classification 570,771 2
+Rotten Tomatoes Movie Review [124] $\times$ Sentiment analysis Sentiment classification 48,869 3/4
+Amazon [125] $\times$ Sentiment analysis Sentiment classification, aspect-based sentiment analysis 34,686,770 5
+SNLI [119] $\times$ Semantic inference Natural language inference, semantic similarity 570,152 3
+MNLI [126] $\times$ Semantic inference Natural language inference, semantic similarity, generalisation 432,702 3
+AGNews [127] $\times$ Text analysis Text classification, sentiment classification 127,600 4
+CogComp QC [128] $\times$ Text analysis Question classification, semantic understanding 15,000 6/50
+Toxic Comment [120] $\sqrt{}$ Text analysis Toxic comment classification, fine-grained toxicity analysis, bias analysis 18,560 6
+Table 3: Summary of the main features of the datasets used in NLP verification.
+2.5.1 Datasets Proposed in This Paper
+In this paper, we focus on two existing datasets that model safety-critical scenarios.
+These two datasets have not previously been applied or explored in the context of NLP verification.
+Both are driven by real-world use cases of safety-critical NLP applications, i.e.
+applications for which law enforcement and safety demand formal guarantees of “good” DNN behaviour.
+Chatbot Disclosure (R-U-A-Robot Dataset [129]).
+The first case study is motivated by new legislation which states that a chatbot must not mislead people about its artificial identity [11, 10].
+Given that the regulatory landscape surrounding NLP models (particularly LLMs and generative AI) is rapidly evolving, similar legislation could be widespread in the future – with recent calls for the US Congress to formalise such disclosure requirements [130].
+The prohibition on deceptive conduct act may apply to the outputs generated by NLP systems if used commercially [131], and at minimum a system must guarantee a truthful response when asked about its agency [129, 132].
+Furthermore, the burden of this should be placed on the designers of NLP systems, and not on the consumers.
+Our first safety critical case is the R-U-A-Robot dataset [129], a written English dataset consisting of 6800 variations on queries relating to the intent of ‘Are you a robot?’, such as ‘I’m a man, what about you?’.
+The dataset was created via a context-free grammar template, crowd-sourcing and pre-existing data sources.
+It consists of 2,720 positive examples (where given the query, it is appropriate for the system to state its non-human identity), 3,400 negative examples and 680 ‘ambiguous-if-clarify’ examples (where it is unclear whether the system is required to state its identity).
+The dataset was created to promote transparency which may be required when the user receives unsolicited phone calls from artificial systems.
+Given systems like Google Duplex [133], and the criticism it received for human-sounding outputs [134], it is also highly plausible for the user to be deceived regarding the outputs generated by other NLP-based systems [131].
+Thus we choose this dataset to understand how to enforce such disclosure requirements.
+We collapse the positive and ambiguous examples into one label, following the principle of ‘better be safe than sorry’, i.e.
+prioritising a high recall system.
+Medical Safety Dataset.
+Another scenario one might consider is that inappropriate outputs of NLP systems have the potential to cause harm to human users [13].
+For example, a system may give a user false impressions of its ‘expertise’ and generate harmful advice in response to medically related user queries [7].
+In practice it may be desirable for the system to avoid answering such queries.
+Thus we choose the Medical safety dataset [12], a dataset consisting of 2,917 risk-graded medical and non-medical queries (1,417 and 1,500 examples respectively).
+The dataset was constructed via collecting questions posted on reddit, such as r/AskDocs.
+The medical queries have been labelled by experts and crowd annotators for both relevance and levels of risk (i.e.
+non-serious, serious to critical) following established World Economic Forum (WEF) risk levels designated for chatbots in healthcare [135].
+We merge the medical queries of different risk-levels into one class, given the high scarcity of the latter two labels to create an in-domain/out-of-domain classification task for medical queries.
+Additionally, we consider only the medical queries that were labelled as such by expert medical practitioners.
+Thus this dataset will facilitate discussion on how to guarantee a system recognises medical queries, in order to avoid generating medical output.
+An additional benefit of these two datasets is that they are distinct semantically, i.e.
+the R-U-A-Robot dataset contains several semantically similar, but lexically different queries, while the medical safety dataset contains semantically diverse queries.
+For both datasets, we utilise the same data splits as given in the original papers, and refer to the final binary labels as positive and negative.
+The positive label in the R-U-A-Robot dataset implies a sample where it is appropriate to disclose non-human identity, while in the medical safety dataset it implies an in-domain medical query.
+2.6 Previous NLP Verification Approaches
 Although DNN verification studies have predominantly focused on computer vision, there is a growing body of research exploring the verification of NLP.
-This research can be categorised into three main approaches: IBP, abstract interpretation, and randomised smoothing.
-Table 1 shows a comparison of these approaches.
+This research can be categorised into three main approaches: using IBP, zonotopes, and randomised smoothing.
+Tables 1 and 2 show a comparison of these approaches.
 To the best of our knowledge, this paper is the first one to use an SMT-based verifier for this purpose, and compare it with an abstract interpretation-based verifier on the same benchmarks.
-Verification via Interval Bound Propagation.
-The first technique successfully adapted from the computer vision domain for verifying NLP models was the IBP.
-In the NLP approaches, IBP is used for both training and verification.
-Its aim is to minimise the upper bound on the maximum difference between the classification boundary and the input perturbation region by augmenting the loss function.
-This facilitates the minimisation of the perturbation region in the last layer, ensuring it remains on one side of the classification boundary.
-As a result, the adversarial region becomes tighter and can be considered certified robust.
+NLP Verification via Interval Bound Propagation.
+The first technique successfully adopted from the computer vision domain for verifying NLP models was the IBP.
+IBP was used for both training and verification with the aim to minimise the upper bound on the maximum difference between the classification boundary and the input perturbation region.
+It was achieved by augmenting the loss function with a term that penalises large perturbations.
+Specifically, IBP incorporates interval bounds during the forward propagation phase, adding a regularisation term to the loss function that minimises the distance between the perturbed and unperturbed outputs.
+This facilitated the minimisation of the perturbation region in the last layer, ensuring it remained on one side of the classification boundary.
+As a result, the adversarial region becomes tighter and can be considered certifiably robust.
 Notably, Jia et al.
@@
 The authors employed IBP to optimise the upper bound over perturbations, providing an upper bound over the discrete set of perturbations in the word vector space.
+Similarly, POPQORN[109] introduced robustness guarantees for RNN-based networks by handling the non-linear activation functions of complicated RNN structures (like LSTMs and GRUs) using linear bounds.
+Later, Shi et al.[15] developed a verification algorithm for transformers with self-attention layers.
+This algorithm provides a lower bound to ensure the probability of the correct label remains consistently higher than that of the incorrect labels.
 Furthermore, Huang et al.
-[50] introduced a verification and verifiable training method for neural networks in NLP, proposing a tighter over-approximation in the form of a ‘simplex’ in the embedding space for input perturbations.
+[18] introduced a verification and verifiable training method with a tighter over-approximation in style of the Simplex algorithm [28].
 To make the network verifiable, they defined the convex hull of all the original unperturbed inputs as a space of perturbations.
@@
 The strength of IBP-based methods is their efficiency and speed, while their main limitation is the bounds’ looseness, further accentuated if the neural network is deep.
-Verification via Abstract Interpretation.
-Another popular verification technique applied to various NLP models is based on abstract interpretation.
-Abstract interpretation was first developed by Cousot and Cousot [29] in 1977.
-It formalises the idea of abstraction of mathematical structures, in particular those involved in the specification of properties and proof methods of computer systems [28] and it has since been used in many applications [30].
-Specifically, for DNN verification, this technique can model the behaviour of a network using an abstract domain that captures the possible range of values the network can output for a given input.
-This abstract domain can then be used to reason about the network’s behaviour under different conditions, such as when the network receives inputs that are adversarially perturbed.
-One notable contribution in this area is POPQORN [61], which is the first work that gives robustness guarantees for RNN-based networks.
-They handle the challenging non-linear activation functions of complicated RNN structures (like LSTMs and GRUs) by bounding them with linear functions.
-Later on, Du et al.
-improve on POPQORN by introducing Cert-RNN [37], a robust certification framework for RNNs that overcomes the limitations of POPQORN.
+NLP Verification via Propagating Zonotopes.
+Another popular verification technique applied to various NLP models is based on propagating zonotopes, which produces tighter bounds then IBP methods.
+One notable contribution in this area is Cert-RNN [110], a robust certification framework for RNNs that overcomes the limitations of POPQORN.
 The framework maintains inter-variable correlation and accelerates the non-linearities of RNNs for practical uses.
@@
 This results in improved precision and tighter bounds, leading to a significant speedup compared to POPQORN.
-Differently, Shi et al.
-[107] focus on transformers with self-attention layers.
-They developed a verification algorithm that can provide a lower bound to ensure the probability of the correct label is consistently higher than that of the incorrect labels.
 Analogously, Bonaert et al.
@@
 It is specifically designed to verify the robustness of transformers against synonym replacement-based attacks.
-DeepT employes multi-norm Zonotopes to achieve larger robustness radii in the certification and can work with networks much larger than Shi et al.
-[…]
-Verification via Randomised Smoothing.
+DeepT employs multi-norm zonotopes to achieve larger robustness radii in the certification and can work with networks much larger than Shi et al.
+Methods that propagate zonotopes produce much tighter bounds than IBP-based methods, which can be used with deeper networks.
+However, they use geometric methods and do not take into account semantic considerations (e.g.
+do not use semantic perturbations).
+NLP Verification via Randomised Smoothing.
 Randomised smoothing [137] is another technique for verifying the robustness of deep language models that has recently grown in popularity due to its scalability [112, 113, 114, 115, 116, 117, 118].
-Its basic idea is to leverage randomness during inference to create a smoothed classifier that is more robust to small perturbations in the input.
+The idea is to leverage randomness during inference to create a smoothed classifier that is more robust to small perturbations in the input.
 This technique can also be used to give certified guarantees against adversarial perturbations within a certain radius.
@@
 The main advantage of randomised smoothing-based methods is their scalability, indeed recent approaches are tested on larger transformer such as BERT and Alpaca.
-However, their main issue is that they are probabilistic approaches, meaning they give certifications up to a certain probability (e.g., 99.9%).
+However, their main issue is that they are probabilistic approaches, meaning they give certifications up to a certain probability.
 In this work we focus on deterministic approaches, hence we only report these works in Table 2 for completeness without delving deeper into each paper here.
 All randomised smoothing-based approaches use data augmentation obtained by semantic perturbations.
-0.2.4 Data Sets and Use Cases used in NLP Verification
-Existing NLP verification data sets.
-Table 2 summarises the main features and tasks of the datasets used in NLP verification.
-Despite their diverse origins and applications, the use of these datasets in the NLP verification literature converge on text classification, with a predominant focus on binary or multi-class categorisation.
-Furthermore, datasets can be sensitive to perturbations, i.e.
-perturbations can have non-trivial impact on label consistency.
-For example, Jia et al.
-[54] use IBP with the SNLI [15] [^3] A semantic inference dataset that labels whether one sentence entails, contradicts or is neutral to another sentence.
-dataset (see Tables 1 and 2) to show that word perturbations (e.g.
-‘good’ to ‘best’) can change whether one sentence entails another.
-Some works such as [54] try to address this label consistency, while others do not.
-Additionally, we find that the previous research on NLP verification does not utilise safety critical datasets (which strongly motivates the choice of datasets in alternative verification domains), with the exception of [37].
-The papers do not provide detailed motivation as to why the dataset choices were made, however it could be due to the datasets being commonly used in NLP benchmarks (IMDB …).
-Instead, we focus on datasets motivated by safety critical applications.
-Dataset Safety Critical Category Tasks Size Classes
-IMDB [84]
-$\times$
-Sentiment analysis
-Document-level and sentence-level classification
-25,000
-2
-SST [112]
-$\times$
-Sentiment analysis
-Sentiment classification, hierarchical sentiment classification, sentiment span detection
-70,042
-5
-SST2 [112]
-$\times$
-Sentiment analysis
-Sentiment classification
-70,042
-2
-YELP [106]
-$\times$
-Sentiment analysis
-Sentiment classification
-570,771
-2
-Rotten Tomatoes Movie Review [96]
-$\times$
-Sentiment analysis
-Sentiment classification
-48,869
-3/4
-Amazon [86]
-$\times$
-Sentiment analysis
-Sentiment classification, aspect-based sentiment analysis
-34,686,770
-5
-SNLI [15]
-$\times$
-Semantic inference
-Natural language inference, semantic similarity
-570,152
-3
-MNLI [127]
-$\times$
-Semantic inference
-Natural language inference, semantic similarity, generalisation
-432,702
-3
-AGNews [143]
-$\times$
-Text analysis
-Text classification, sentiment classification
-127,600
-4
-CogComp QC [74]
-$\times$
-Text analysis
-Question classification, semantic understanding
-15,000
-6/50
-Toxic Comment [26]
-Text analysis
-Toxic comment classification, fine-grained toxicity analysis, bias analysis
-18,560
-6
-Table 2: Summary of the main features of the datasets used in NLP verification.
-Data Sets Proposed in This Paper
-In this paper we propose to use two datasets that have not been used in NLP verification literature before.
-Both are driven by real-world use cases of safety-critical NLP applications, i.e.
-applications for which law enforcement and safety demand formal guarantees of “good" DNN behaviour.
-Chatbot Disclosure Dataset.
-First case concerns new legislation which states that a chatbot must not mislead people about its artificial identity [68, 63].
-Given that the regulatory landscape surrounding NLP models (particularly LLMs and generative AI) is rapidly evolving, similar legislation could be widespread in the future – with recent calls for the US Congress to formalise such disclosure requirements [90].
-The prohibition on deceptive conduct act may apply to the outputs generated by NLP systems if used commercially [4], and at minimum a system must guarantee a truthful response when asked about its agency [47, 1].
-Furthermore, the burden of this should be placed on the designers of NLP systems, and not on the consumers.
-Our first safety critical case is the R-U-A-Robot dataset [47], a written English dataset consisting of 6800 variations on queries relating to the intent of ‘Are you a robot?’, such as ‘I’m a man, what about you?’.
-The dataset was created via a context-free grammar template, crowd-sourcing and pre-existing data sources.
-It consists of 2,720 positive examples (where given the query, it is appropriate for the system to state its non-human identity), 3,400 negative/adversarial examples and 680 ‘ambiguous-if-clarify’ examples (where it is unclear whether the system is required to state its identity).
-The dataset was created to promote transparency which may be required when the user receives unsolicited phone calls from artificial systems.
-Given systems like Google Duplex [70], and the criticism it received for human-sounding outputs [77], it is also highly plausible for the user to be deceived regarding the outputs generated by other NLP-based systems [4].
-Thus we choose this dataset to understand how to enforce such disclosure requirements.
-We collapse the positive and ambiguous examples into one label, following the principle of ‘better be safe than sorry’, i.e.
-prioritising a high recall system.
-Medical Safety Dataset.
-Another scenario one might consider is that inappropriate outputs of NLP systems have the potential to cause harm to human users [12].
-For example, a system may give a user false impressions of its ‘expertise’ and generate harmful advice in response to medically related user queries [34].
-In practice it may be desirable for the system to avoid answering such queries.
-Thus we choose the Medical safety dataset [2], a written English dataset consisting of 2,917 risk-graded medical and non medical queries (1,417 and 1,500 examples respectively).
-The dataset was constructed via collecting questions posted on reddit, such as r/AskDocs.
-The medical queries have been labelled by experts and crowd annotators for both relevance and levels of risk (i.e.
-non-serious, serious to critical) following established World Economic Forum (WEF) risk levels designated for chatbots in healthcare [130].
-We merge the medical queries of different risk-levels into one class, given the high scarcity of the latter 2 labels to create an in-domain/out-of-domain classification task for medical queries.
-Additionally, we consider only the medical queries that were labelled as such by expert medical practitioners.
-Thus this dataset will facilitate discussion on how to guarantee a system recognises medical queries, in order to avoid generating medical output.
-An additional benefit of these two datasets is that they are distinct semantically, i.e.
-the R-U-A-Robot dataset contains several semantically similar, but lexically different queries, while the medical safety dataset contains semantically diverse queries.
-For both datasets, we utilise the same data splits as given in the original papers, and refer to the final binary labels as positive and negative.
-The positive label in the R-U-A-Robot dataset implies a sample where it is appropriate to disclose non-human identity, while in the medical safety dataset it implies an in-domain medical query.
-0.2.5 Our Work (NLP Verification Pipelines)
-To show relation of our work to the body of already existing work, we distill an “NLP verification pipeline” that is common across many related papers.
-Figure 2 shows the pipeline diagrammatically.
+To systematically compare the existing body of research, we distil an “NLP verification pipeline” that is common across many related papers.
+This pipeline is outlined diagrammatically in Figure 2, while Tables 1 and 2 provide a detailed breakdown, with columns corresponding to each stage of the pipeline.
 It proceeds in stages:
@@
 IBP and randomised smoothing use word and character perturbations, abstract interpretation papers usually do not use any semantic perturbations.
-Our method allows to use all existing semantic perturbations, in particular, we implement character and word level perturbations as in [91], sentence level perturbations with PolyJuice [132] and Vicuna.
+Tables 1 and 2 give the exact mapping of perturbation methods to papers.
+Our method allows to use all existing semantic perturbations, in particular, we implement character and word level perturbations as in Moradi et al.
+[103], sentence level perturbations with PolyJuice [21] and Vicuna.
   2. 2.
@@
 […]
-    Given a notion of geometric/semantic subspaces, one can use it in two different ways:
-    • 4.
+  4. 4.
 […]
-    • 5.
-      $\ldots$ verify the classifier’s behaviour within those subspaces.
-IBP papers use IBP algorithms which are incomplete and imprecise, abstract interpretation gives incomplete and imprecise methods, and we use SMT-based tool Marabou (complete and precise) and abstract-interpretation tool ERAN (incomplete and imprecise).
-Table 1 summarises differences and similarities of the above NLP verification approaches against ours.
-To the best of our knowledge, we are the first to use complete and precise methods in NLP verification.
-This paper is the first to employ SMT-based verifiers and to show how they out-perform abstract interpretation-based verification, which produce tighter bounds than IBP-based methods.
+  5. 5.
+    Use the geometric/semantic subspaces to verify the classifier’s behaviour within those subspaces.
+The papers [17, 18, 107, 19, 108] use IBP algorithms and the papers  [109, 15, 110, 111] use abstract interpretation; in both cases it is incomplete and deterministic verification.
+See ‘Verification algorithm’ and ‘Verification characteristics’ columns of Tables 1 and 2.
+We use SMT-based tool Marabou (complete and deterministic) and abstract-interpretation tool ERAN (incomplete and deterministic).
+Tables 1 and 2 summarise differences and similarities of the above NLP verification approaches against ours.
+To the best of our knowledge, we are the first to use SMT-based complete methods in NLP verification and we show how they achieve higher verifiability than abstract interpretation verification approaches (ERAN and CORA) or IBP and BaB ($\alpha\beta$-CROWN), thanks to the increased precision of the ReLUplex algorithm that underlies Marabou.
 Furthermore, our study is the first to demonstrate that the construction of semantic subspaces can happen independently of the choice of the training and verification algorithms.
-Likewise, although training and verification build upon the defined (semantic) subpaces, the actual choice of the training and verification algorithms can be made independently of the method used to define the semantic subspaces.
+Likewise, although training and verification build upon the defined (semantic) subspaces, the actual choice of the training and verification algorithms can be made independently of the method used to define the semantic subspaces.
 This separation, and the general modularity of our approach, facilitates a comprehensive examination and comparison of the two key components involved in any NLP verification process:
   • effects of the verifiability-generalisability trade-off for verification with geometric and semantic subspaces;
-  • relation between the volume/shape of semantic subpaces and verifiability of neural networks obtained via semantic training with these subpaces.
+  • relation between the volume/shape of semantic subspaces and verifiability of neural networks obtained via semantic training with these subspaces.
 These two aspects have not been considered in the literature before.
-0.3 The Parametric NLP Verification Pipeline
+3 The Parametric NLP Verification Pipeline
+This section presents a parametric NLP verification pipeline, shown in Figure 2 diagrammatically.
+We call it “parametric” because each component within the pipeline is defined independently of the others and can be taken as a parameter when studying other components.
+The parametric nature of the pipeline allows for the seamless integration of state-of-the-art methods at every stage, and for more sophisticated experiments with those methods.
+The following section provides a detailed exposition of the methodological choices made at each step of the pipeline.
 Figure 2: Visualisation of the NLP verification pipeline followed in our approach.
-This section presents a parametric NLP verification pipeline, shown in Figure 2 diagrammatically.
-We call it “parametric” because each component within the pipeline operates independently of the others and can be taken as a parameter when studying other components.
-The parametric nature of the pipeline allows for the seamless integration of state-of-the-art methods at every stage, and for more sophisticated experiments with those methods.
-Note that the outlined pipeline can be seen as a filter which can be applied on top of an NLP system or LLM (such as S-BERT and S-GPT) to certify intended DNN behavior for safety-critical input queries.
-The following section provides a detailed exposition of the methodological choices made at each step of the pipeline.
-0.3.1 Semantic Perturbations
-As discussed in Section 0.2.5, we require semantic perturbations for creating semantic subspaces.
+3.1 Semantic Perturbations
+As discussed in Section 2.6, we require semantic perturbations for creating semantic subspaces.
 To do so, we consider three kinds of perturbations – i.e.
@@
 This systematically accounts for different variations of the samples.
-Character and word level perturbations
-are created via a rule-based method proposed in [91], to simulate different kinds of noise one could expect from spelling mistakes, typos etc.
+Character and word level perturbations are created via a rule-based method proposed by Moradi et al.
+[103] to simulate different kinds of noise one could expect from spelling mistakes, typos etc.
 These perturbations are non-adversarial and can be generated automatically.
-The authors found that NLP models are sensitive to such small errors, while in practice this should not be the case.
+Moradi et al.
+[103] found that NLP models are sensitive to such small errors, while in practice this should not be the case.
 Character level perturbations types include randomly inserting, deleting, replacing, swapping or repeating a character of the data sample.
-At the character level, we do not apply letter case changing, given it does not change the sentence-level representation of the sample, nor do we apply perturbations to commonly misspelled words, given only a small overlap of those words occur in our datasets.
+At the character level, we do not apply letter case changing, given it does not change the sentence-level representation of the sample.
+Nor do we apply perturbations to commonly misspelled words, given only a small percentage of the most commonly misspelled words occur in our datasets.
 Perturbations types at the word level include randomly repeating or deleting a word, changing the ordering of the words, the verb tense, singular verbs to plural verbs or adding negation to the data sample.
@@
 please help’).
-Further examples of character and word rule-based perturbations can be found in Tables 3 and 4.
-Method Description Original sentence Altered sentence
-Insertion
-A character is randomly selected and inserted in a random position.
-Are you a robot?
-Are yovu a robot?
-Deletion
-A character is randomly selected and deleted.
-Are you a robot?
-Are you a robt?
-Replacement
-A character is randomly selected and replaced by an adjacent character on the keyboard.
-Are you a robot?
-Are you a ronot?
-Swapping
-A character is randomly selected and swapped with the adjacent right or left character in the word.
-Are you a robot?
-Are you a rboot?
-Repetition
-A character in a random position is selected and duplicated.
-Are you a robot?
-Arre you a robot?
+Note that the Medical dataset contains several sentences without a verb (like the one above) for which it is impossible to pluralise or change the tense of the verb.
+Further examples of character and word rule-based perturbations can be found in Tables 4 and 5.
+Method Description Altered sentence
+(Are you a robot?)
+Insertion A character is randomly selected and inserted in a random position. Are yovu a robot?
+Deletion A character is randomly selected and deleted. Are you a robt?
+Replacement A character is randomly selected and replaced by an adjacent character on the keyboard. Are you a ronot?
+Swapping A character is randomly selected and swapped with the adjacent right or left character in the word. Are you a rboot?
+Repetition A character in a random position is selected and duplicated. Arre you a robot?
 Table 4: Character-level perturbations: their types and examples of how each type acts on a given sentence from the R-U-A-Robot dataset [129].
 Perturbations are selected from random words that have 3 or more characters, first and last characters of a word are never perturbed.
-Method Description Original sentence Altered sentence
-Deletion
-Randomly selects a word and removes it.
-Can u tell me if you are a chatbot?
-Can u tell if you are a chatbot?
-Repetition
-Randomly selects a word and duplicates it.
-Can u tell me if you are a chatbot?
-Can can u tell me if you are a chatbot?
-Negation
-Identifies verbs then flips them (negative/positive).
-Can u tell me if you are a chatbot?
-Can u tell me if you are not a chatbot?
-Singular/ plural verbs
-Changes verbs to singular form, and conversely.
-Can u tell me if you are a chatbot?
-Can u tell me if you is a chatbot?
-Word order
-Randomly selects consecutive words and changes the order in which they appear.
-Can u tell me if you are a chatbot?
-Can u tell me if you are chatbot a?
-Verb tense
-Converts present simple or continuous verbs to their corresponding past simple or continuous form.
-Can u tell me if you are a chatbot?
-Can u tell me if you were a chatbot?
+Method Description Altered sentence
+(Can u tell me if you are a chatbot?)
+Deletion Randomly selects a word and removes it. Can u tell if you are a chatbot?
+Repetition Randomly selects a word and duplicates it. Can can u tell me if you are a chatbot?
+Negation Identifies verbs then flips them (negative/positive). Can u tell me if you are not a chatbot?
+Singular/ plural verbs Changes verbs to singular form, and conversely. Can u tell me if you is a chatbot?
+Word order Randomly selects consecutive words and changes the order in which they appear. Can u tell me if you are chatbot a?
+Verb tense Converts present simple or continuous verbs to their corresponding past simple or continuous form. Can u tell me if you were a chatbot?
 Table 5: Word-level perturbations: their types and examples of how each type acts on a given sentence from the R-U-A-Robot dataset [129].
@@
 Vicuna is a state-of-the-art open source chatbot trained by fine-tuning LLaMA [138] on user-shared conversations collected from ShareGPT  [^5] https://sharegpt.com/.
-For Vicuna, we use the following prompt to generate variations on our data samples ‘Rephrase this sentence 5 times: “[Example]”.’ For example, from the sentence “How long will I be contagious?", we can obtain “How many years will I be contagious?" or “Will I be contagious for long?" and so on.
+For Vicuna, we use the following prompt to generate variations on our data samples ‘Rephrase this sentence 5 times: “[Example]”.’ For example, from the sentence “How long will I be contagious?”, we can obtain “How many years will I be contagious?” or “Will I be contagious for long?” and so on.
 We will use notation $\mathcal{P}$ to refer to a perturbation algorithm abstractly.
-0.3.2 NLP Embeddings
-Next component of the pipeline are embeddings.
+Semantic similarity of perturbations.
+In later sections we will make an assumption that the perturbations that we use produce sentences that are semantically similar to the originals.
+However, precisely defining or measuring semantic similarity is a challenge in its own right, as semantic meaning of sentences can be subjective, context-dependent, which makes evaluating their similarity intractable.
+Nevertheless, Subsection 5.5.2 will discuss and use several metrics for calculating semantic similarity of sentences, modulo some simplifying assumptions.
+3.2 NLP Embeddings
+The next component of the pipeline is the embeddings.
 Embeddings play a crucial role in NLP verification as they map textual data into continuous vector spaces, in a way that should capture semantic relationships and contextual information.
-Given an NLP dataset $\mathcal{Y}$ as a set of sentences $s_{1},\ldots,s_{q}$ written in natural language, an embedding $\mathbb{E}$ is a function that maps a sentence to a vector in ${}^{m}$.
-The vector space ${}^{m}$ is called the embedding space.
-Ideally, $\mathbb{E}$ should reflect the semantic similarities between sentences in $\mathcal{Y}$, i.e.
-the more semantically similar two sentences $s_{i}$ and $s_{j}$ are, the closer the distance between $\mathbb{E}(s_{i})$ and $\mathbb{E}(s_{j})$ should be in ${}^{m}$.
-Of course, defining semantic similarity in precise terms may not be tractable (the number of unseen sentences maybe infinite, the similarity maybe understood subjectively and/or depending on the context).
+Given the set of all strings, $\mathbb{S}$, an NLP dataset $\mathcal{Y}\subset\mathbb{S}$ is a set of sentences $s_{1},\ldots,s_{q}$ written in natural language.
+The embedding $E$ is a function that maps a string in $\mathbb{S}$ to a vector in m.
+The vector space m is called the embedding space.
+Ideally, $E$ should reflect the semantic similarities between sentences in $\mathbb{S}$, i.e.
+the more semantically similar two sentences $s_{i}$ and $s_{j}$ are, the closer the distance between $E(s_{i})$ and $E(s_{j})$ should be in m.
+Of course, defining semantic similarity in precise terms may not be tractable (the number of unseen sentences may be infinite, the similarity may be subjective and/or depend on the context).
 This is why, the state-of-the-art NLP relies on machine learning methods to capture the notion of semantic similarity approximately.
-Currently, the most common approach to obtain an embedding function $\mathbb{E}$ is by training transformers [32, 101].
+Currently, the most common approach to obtain an embedding function $E$ is by training transformers [139, 140].
 Transformers are a type of DNNs that can be trained to map sequential data into real vector spaces and are capable of handling variable-length input sequences.
 They can also be used for other tasks, such as classification or sentence generation, but in those cases, too, training happens at the level of embedding spaces.
-In this work, a transformer is trained as a function $\mathbb{E}:\mathcal{Y}\rightarrow^{m}$ for some given $m$.
-The key feature of the transformer is the “self-attention mechanism", which allows the network to weigh the importance of different elements in the input sequence when making predictions, rather than relying solely on the order of elements in the sequence.
+In this work, a transformer is trained as a function $E:\mathbb{S}\rightarrow^{m}$ for some given $m$.
+The key feature of the transformer is the “self-attention mechanism”, which allows the network to weigh the importance of different elements in the input sequence when making predictions, rather than relying solely on the order of elements in the sequence.
 This makes them good at learning to associate semantically similar words or sentences.
@@
 Given an NLP dataset $\mathcal{Y}$ that should be classified into $n$ classes, the standard approach is to construct a function $N:^{m}\rightarrow^{n}$ that maps the embedded inputs to the classes.
-In order to do that, a domain specific classifier $N$ is trained on the embeddings $\mathbb{E}(\mathcal{Y})$ and the final system will then be the composition of the two subsystems, i.e.
-$N\circ\mathbb{E}$.
-0.3.3 Geometric Analysis of Embedding Spaces
-We now formally define geometric and semantic subpaces of the embedding space. Our goal is to define subpaces on the embedding space ${}^{m}$ by using an effective algorithmic procedure.
-We start with an observation that, given an NLP dataset $\mathcal{Y}$ that contains (a finite number of sentences) $s_{1},\ldots,s_{q}$, and an embedding function $\mathbb{E}:\mathcal{Y}\rightarrow^{m}$, we can define an embedding matrix $\mathcal{X}\in^{q\times m}$, where each row $j$ is given by $\mathbb{E}(s_{j})$.
+In order to do that, a domain specific classifier $N$ is trained on the embeddings $E(\mathcal{Y})$ and the final system will then be the composition of the two subsystems, i.e.
+$N\circ E$.
+3.3 Geometric Analysis of Embedding Spaces
+In the recent years, the study of manifold subspaces has gained significant attention in the context of machine learning verification [55], where the geometry of data regions plays an important role.
+In this section, we formally define most common subspaces used in verification: convex sets, convex hulls, zonotopes, and hyper-rectangles (also known as multi-dimensional intervals), following closely [55].
+Definition 1 (Convex Set)
+A set $\mathbb{Z}\subseteq^{m}$ is said to be convex if, for any two points $x_{1},x_{2}\in\mathbb{Z}$, the line segment joining them is entirely contained within $\mathbb{Z}$.
+Formally, this means that for all $x_{1},x_{2}\in\mathbb{Z}$ and $\lambda\in[0,1]$, the points
+$\lambda x_{1}+(1-\lambda)x_{2}\in\mathbb{Z}.$
+In other words, a set is convex if, for any pair of points in the set, the entire segment connecting them lies within the set.
+The convex hull of a set is the smallest convex set that contains all the points of the set. It can be seen as the “tightest” boundary enclosing the points.
+Definition 2 (Convex Hull)
+The convex hull of a set $\mathbb{Z}\subseteq^{m}$, denoted by $\text{conv}(\mathbb{Z})$, is the smallest convex set containing $\mathbb{Z}$. Formally, it can be defined as:
+$\text{conv}(\mathbb{Z}):=\left\{\sum_{i=1}^{p}\lambda_{i}x_{i}\;\middle|\;x_{i}\in\mathbb{Z},\;\lambda_{i}\geq 0,\;\sum_{i=1}^{p}\lambda_{i}=1,\;p\in{\rm Nature}\right\}.$
+In other words, $\text{conv}(\mathbb{Z})$ consists of all finite convex combinations of points in $\mathbb{Z}$.
+The construction of the convex hull has a complexity of $O(p^{m/2})$, where $p$ is the number of points and $m$ is the number of dimensions.
+A zonotope is a geometric shape formed by the Minkowski sum of line segments.
+Definition 3 (Zonotope)
+Given a center $c\in^{m}$ and generators $g_{1},\ldots,g_{p}$, a zonotope is
+$\mathcal{Z}:=\left\{c+\sum_{i=1}^{p}\lambda_{i}g_{i}\;\middle|\;\lambda_{i}\in[-1,1],\forall i\in[1,\ldots,p]\right\}$
+Zonotopes are computationally more efficient than convex hulls, with a construction complexity of $O(m\cdot p)$, where $m$ is the dimensionality and $p$ is the number of generators.
+Finally, an interval is a simple shape defined by lower and upper bounds for each dimension, and it is equivalent to a multi-dimensional rectangle (or hyper-rectangle).
+Intervals are easy to construct with a complexity of $O(m)$, where $m$ is the dimensionality, and are often used in verification.
+Definition 4 (Interval (aka Hyper-Rectangle))
+Given a lower and upper bound $\underline{x},\overline{x}\in^{m}$ such that $\underline{x}_{(i)}\leq\overline{x}_{(i)}\forall i\in{1,\ldots,m}$, a multi-dimensional interval $\mathcal{I}\subset^{m}$ is
+:=$\mathcal{I}:=\left\{x\in^{m}\;\middle|\;\underline{x}_{(i)}\leq x_{(i)}\leq\overline{x}_{(i)},\forall i\in[1,\ldots,m]\right\}$
+Table 6 summarises the construction complexities of these different shapes.
+Ideally, convex hulls would be the preferred choice due to their precise and detailed representations of subspaces.
+However, their computational complexity renders them infeasible in high dimensions.
+Zonotopes provide a promising alternative, as they are more precise than hyper-rectangles while remaining computationally tractable.
+Despite their theoretical compatibility with complete verifiers, practical limitations arise because most state-of-the-art verifiers do not support zonotopes.
+Hyper-rectangles, or intervals, are the simplest to construct and are supported by all verifiers, making them the default choice in many verification pipelines.
+Shape Construction Complexity (Big-O)
+Convex Hull $O(p^{m/2})$
+Zonotope $O(m\cdot p)$
+Interval $O(m)$
+Table 6: Construction complexity for different geometric shapes, where $m$ is the number of dimensions and $p$ is the number of points or generators.
+3.4 Working with Embedding Spaces: Our Approach
+We now formally define geometric and semantic subspaces of the embedding space.
+Our goal is to define subspaces on the embedding space m by using an effective algorithmic procedure.
+We will use notation $\mathcal{S}$ to refer to a subspace of the embedding space.
+Recall that an hyper-rectangle of dimension $m$ is a list of points $(a_{1},b_{1}),\ldots,(a_{m},b_{m})$ such that a point $x\in^{m}$ is a member if for every dimension $j$ we have $a_{j}\leq x_{j}\leq b_{j}$.
+We start with an observation that, given an NLP dataset $\mathcal{Y}$ that contains a finite set of sentences $s_{1},\ldots,s_{q}$ belonging to the same class, and an embedding function $E:\mathbb{S}\rightarrow^{m}$, we can define an embedding matrix $\mathcal{X}\in^{q\times m}$, where each row $j$ is given by $E(s_{j})$.
+We will use the notation $x_{i}$ to refer to the $i$th element of the vector $x$, and $\mathcal{X}^{ij}$ to refer to the element in the $i$th row and $j$th column of $\mathcal{X}$.
 Treating embedded sentences as matrices, rather than as points in the real vector space, makes many computations easier.
-This motivates the following definition of a hyper-rectangle for $\mathcal{X}$.
-Definition 1 (Hyper-rectangle for an Embedding Matrix).
+We can therefore define a hyper-rectangle for $\mathcal{X}$ as follows.
+Definition 5 (Hyper-rectangle for an Embedding Matrix)
 Given an embedding matrix $\mathcal{X}\in^{q\times m}$, the $m$-dimensional hyper-rectangle for $\mathcal{X}$ is defined as:
-$\mathbb{H}_{\mathcal{X}}\coloneqq\{x\in^{m}:\forall j\in[1,...,m].\>x^{j}\in[\min\nolimits_{i}(\mathcal{X}^{ij}),\max\nolimits_{i}(\mathcal{X}^{ij})]\}.$
-where $\min\nolimits_{i}(\mathcal{X}^{ij})$ stands for the minimum value of the $j$-th element $x^{j}$ across all rows $i\in\{0,\ldots,q\}$ in $\mathcal{X}$; similarly for $\max\nolimits_{i}(\mathcal{X}^{ij})$.
-Definition 2 (Subspace of the Embedding Space).
-Given a NLP dataset $\mathcal{Y}$, an embedding function $\mathbb{E}:\mathcal{Y}\rightarrow^{m}$, and $\mathcal{Y}^{\prime}\subseteq\mathcal{Y}$ (with $\mathcal{Y}^{\prime}=\{s_{1},\ldots,s_{q}\}$), we say a hyper-rectangle $\mathbb{H}_{\mathcal{X}^{\prime}}$ is a subspace of the embedding space ${}^{m}$ around $\mathcal{Y}^{\prime}$ if the matrix $\mathcal{X}^{\prime}\in^{q\times m}$ satisfies the following condition: each $j$th row is given by $\mathbb{E}(s_{j})$.
+$\mathbb{H}(\mathcal{X})\coloneqq\{(\min\nolimits_{i=0}^{q}\mathcal{X}^{ij},~{}\max\nolimits_{i=0}^{q}\mathcal{X}^{ij})\mid j\in[1,...,m]\}$
+Therefore given an embedding function $E:\mathbb{S}\rightarrow^{m}$, and a set of sentences $\mathcal{Y}=\{s_{1},\ldots,s_{q}\}$, we can form a subspace $\mathbb{H}(E(\mathcal{Y}))$ by constructing the embedding matrix, as described above, and forming the corresponding hyper-rectangle.
+To simplify the notation, we will omit the application of $E$ and from here on simply write $\mathbb{H}(\mathcal{Y})$.
 […]
@@
 $\mathbb{B}(\hat{x},\epsilon)\coloneqq\{x\in^{m}:||\hat{x}-x||\leq\epsilon\}.$
-[…]
 […]
@@
 Approaches that use under-approximations to speed up the algorithms [144, 145] do not work well in NLP scenarios, as under-approximated subspaces are so small that they contain near zero sentence embeddings.
-A computationally efficient way of making hype-rectangles more precise is to rotate them to align to the position of the given points in the embedding space.
+3.4.1 Exclusion of Unwanted Sentences Via Shrinking
+Another concern is that the generated hyper-rectangles may contain sentences from a different class.
+This would make it unsuitable for verification.
+In order to exclude all samples from the wrong class, we define a shrinking algorithm $SH(\mathcal{X},\mathcal{Y},c)$ that calculates a new subspace that is a subset of the original hyper-rectangle around $\mathcal{X}$, that only contains embeddings of sentences in $\mathcal{Y}$ that are of class $c$.
+Of course, to ensure this, the algorithm may have to exclude some sentences of class $c$.
+The second graph of Figure 3 gives a visual intuition of how this is done.
+Figure 3: An example of hyper-rectangle drawn around all points of the same class (a), shrunk hyper-rectangle $\mathbb{H}_{sh}$ that is obtained by excluding all points from the opposite class (b) and clustered hyper-rectangles (c) in 2-dimensions.
+The red dots represent sentences in the embedding space of one class, while the blue dots are embedded sentences that do not belong to that class.
+Formally, for each sentence $s$ in $\mathcal{Y}$ that is not of class $c$, the algorithm performs the following procedure.
+If $E(s)$ lies in the current hyper-rectangle $(a_{1},b_{1}),\dots,(a_{m},b_{m})$, then for each dimension ${j\in[1,\dots,m]}$ we compute the distance whether $E(s)_{j}$ is closer to $a_{j}$ or $b_{j}$.
+Without loss of generality, assume $a_{j}$ is closer.
+We then compute the number of sentences of class $c$ that would be excluded by replacing $a_{j}$ with $E(s)_{j}+\delta$ in the hyper-rectangle where $\delta$ is a small positive number (we use $e^{-100}$).
+This gives us a penalty for each dimension $j$, and we exclude $s$ by updating the hyper-rectangle in the dimension that minimises this penalty.
+The idea is to shrink the hyper-rectangle in the dimensions that exclude as few embedded sentences from the desired class $c$ as possible [^6] Note that this algorithm shrinks exactly one dimension by a minimal amount to exclude the unwanted embedded sentence.
+This choice keeps the algorithm fast while guaranteeing the subspace to retain the highest number of wanted inputs.
+However, it is not necessarily the best choice for verification: there might be cases where perturbations of the unwanted input are left inside after shrinking and, if the network classifies them correctly, the subspace can never be verified.
+For large subspaces, our algorithm might render verification unachievable and more clever algorithms should be explored and discussed..
+3.4.2 Exclusion of Unwanted Sentences Via Clustering
+An alternative approach to excluding unwanted sentences, is to split the dataset up by clustering semantically similar sentences in the embedding space, and then compute the hyper-rectangles around each cluster individually, as shown in the last graph of Figure 3.
+In this paper we will use the k-means algorithm for clustering.
+We will use the notation $CL(\mathcal{Y},k)$ to refer to the $k$-clusters formed by applying it to dataset $\mathcal{Y}$.
+While in our experiments we have found this is often sufficient to exclude unwanted sentences, it is not guaranteed to do so.
+Therefore, this method is combined with the shrinking algorithm in our experiments.
+3.4.3 Eigenspace Rotation
+A final alternative and computationally efficient way of reducing the likelihood that the hyper-rectangles will contain embedded sentences of an unwanted class, is to rotate them to better align to the distribution of the embedded sentences of the desired class in the embedding space.
 This motivates us to introduce the Eigenspace rotation.
-Eigenspace Rotation.
 To construct the tightest possible hyper-rectangle, we define a specific method of eigenspace rotation.
-As shown in Figure 1 (C and D), our approach is to calculate a rotation matrix $A$ such that the rotated matrix $\mathcal{X}_{\text{rot}}=\mathcal{X}A$ is better aligned with the axes than $\mathcal{X}$, and therefore $\mathbb{H}_{\mathcal{X}_{\text{rot}}}$ has a smaller volume.
-By a slight abuse of terminology, we will refer to $\mathbb{H}_{\mathcal{X}_{\text{rot}}}$ as the rotated hyper-rectangle, even though strictly speaking, we are rotating the data, not the hyper-rectangle itself.
+As shown in Figure 1 (C and D), our approach is to calculate a rotation matrix $A$ such that the rotated matrix $\mathcal{X}_{\text{rot}}=\mathcal{X}A$ is better aligned with the axes than $\mathcal{X}$, and therefore $\mathbb{H}(\mathcal{X}_{\text{rot}})$ has a smaller volume.
+By a slight abuse of terminology, we will refer to $\mathbb{H}(\mathcal{X}_{\text{rot}})$ as the rotated hyper-rectangle, even though strictly speaking, we are rotating the data, not the hyper-rectangle itself.
 In order to calculate the rotation matrix $A$, we use singular value decomposition [146].
@@
 Formally, using $V^{*}$, we can compute the rotation (or change-of-basis) matrix $A$ that rotates the right-singular vectors onto the canonical standard basis vectors $I$, where $I$ is the identity matrix.
-To do this, we observe that $V^{*}A=I\rightarrow V^{*}=IA^{-1}\rightarrow V^{-1}=A^{-1}\rightarrow V=A$.
+To do this, we observe that $V^{*}A=I$ implies $V^{*}=IA^{-1}$, which implies $V^{-1}=A^{-1}$, and thus $V=A$.
 We thus obtain $\mathcal{X}_{rot}=\mathcal{X}A$ as desired.
-All hyper-rectangles constructed and used in this paper are rotated.
-Geometric and Semantic Subspaces.
+All hyper-rectangles constructed in this paper are rotated.
+3.4.4 Geometric and Semantic Subspaces
 […]
 We will finish this section with defining semantic subspaces formally.
-Given a data set $\mathcal{Y}$, and some defined algorithm for sentence perturbation $\mathcal{P}$, we can form $\mathcal{Y}^{\prime}=\{s^{*}_{1},\ldots,s^{*}_{q}\}$, such that each $s^{*}_{i}=\mathcal{P}(s_{i})$.
-Intuitively, we want to construct a subspace for $\mathcal{Y}^{\prime}$ as described in Definition 2.
-However, in the later sections, we will need to refer to different kinds of perturbations (e.g.
-character, word or sentence level as explained in Section 0.3.1) or different types of perturbations (e.g.
-insertion, deletion, replacement, etc.) as illustrated in Tables 3 and 4).
+We will use $\mathcal{P}_{t}(s)$ to denote an algorithm for generating sentence perturbations of type $t$, applied to an input sentence $s$ in a random position.
+In the later sections, we will use $t$ to refer to the different types of perturbations illustrated in Tables 4 and 5, e.g.
+character-level insertion, deletion, replacement.
+Intuitively, given a single sentence we want to generate a set of semantically similar perturbations and then construct a hyper-rectangle around them, as described in Definition 5.
 This motivates the following definitions.
-Notation $\mathcal{P}_{k}^{t}(s,a)$ will denote a perturbation of kind $k$ and type $t$, applied in the sentence $s$, in a position $a$.
-As all algorithms we use choose $a$ randomly, we will omit $a$ and write just $\mathcal{P}_{k}^{t}(s)$.
-When we say that a sentence $s$ has $b$ semantic perturbations of kind $k$, we refer to a set $\mathcal{A}^{k}_{s}=\{\mathcal{P}_{k}^{t_{1}}(s),\mathcal{P}_{k}^{t_{2}}(s),\ldots,\mathcal{P}_{k}^{t_{b}}(s)\}$, for some given choice of $b$ perturbations of types $t_{1},\ldots,t_{b}$.
-When the choice is unimportant or clear from the context, we may omit mentioning the chosen types of perturbation.
-Definition 3 (Semantic Subspace for a Sentence).
-Suppose we are given an NLP dataset $\mathcal{Y}=\{s_{1},\ldots,s_{q}\}$, and an embedding function $\mathbb{E}:\mathcal{Y}\rightarrow^{m}$.
-Given a sentence $s_{i}$ and a corresponding set $\mathcal{A}^{k}_{s_{i}}$ of its semantic perturbations (of kind $k$), a semantic subspace for $s_{i}$ is a subspace around $\mathcal{Y}^{\prime}=\mathcal{A}^{k}_{s_{i}}$ constructed according to Definition 2.
-It will be useful in the later sections to use notation $\mathbb{H}_{\mathcal{X}_{s_{i}}}$ to refer to the hyper-rectangle used in constructing the semantic subspace for $s_{i}$.
-We will refer to a set of such hyper-rectangles $\mathbb{H}^{*}_{k}=\bigcup_{i=1}^{q}\mathbb{H}^{*}_{\mathcal{X}_{s_{i}}}$, where $s_{i}\in\{s_{1},\ldots,s_{q}\}$ and $k$ is a chosen perturbation kind.
-If $\mathbb{H}^{*}_{k}$ is based on perturbations of only one type, we will sometimes use that type as an index to $\mathbb{H}^{*}_{\_}$.
-Example 2 (Construction of Semantic Subspaces).
-To illustrate this construction, let us consider the sentence $s_{i}=$“Can u tell me if you are a chatbot?”.
+Given a sentence $s$, a number $b$, and a type $t$, the set $\mathcal{A}^{b}_{t}(s)=\{\mathcal{P}_{t}(s)\mid i\in[1,b]\}$ is the set of $b$ semantic perturbations of type $t$ generated from $s$.
+We will use the notation $\mathcal{A}^{b}_{t}(\mathcal{Y})=\bigcup_{s\in\mathcal{Y}}\mathcal{A}^{b}_{t}(s)$ to denote the new dataset generated by creating $b$ semantic perturbations of type $t$ around each sentence.
+Definition 6 (Semantic Subspace for a Sentence)
+Given an embedding function $E:\mathbb{S}\rightarrow^{m}$, the semantic subspace for a sentence $s$ is the subspace ($\mathbb{H}(\{s\}\cup\mathcal{A}^{b}_{t}(s))$.
+We will refer to a set of such semantic hyper-rectangles over an entire dataset $\mathcal{Y}$ as $\mathbb{H}_{t}^{b}(\mathcal{Y})=\bigcup_{s\in\mathcal{Y}}\mathbb{H}(\{s\}\cup\mathcal{A}^{b}_{t}(s))$.
+Example 2 (Construction of Semantic Subspaces)
+To illustrate this construction, let us consider the sentence $s$: “Can u tell me if you are a chatbot?”.
 This sentence is one of $3400$ original sentences of the positive class in the dataset.
-From this single sentence, we can create six sentences (word-level perturbations), see Table 4.
-Once these seven sentences are embedded into the vector space, they form the hyper-rectangle $\mathbb{H}_{\mathcal{X}_{s_{i}}}$.
-By repeating this construction for the remaining $3399$ sentences, we obtain the set of hyper-rectangles $\mathbb{H}^{*}_{word}$ for the dataset.
-0.3.4 Training
-As outlined in Section 0.2.2, robust training is essential for bolstering the robustness of DNNs; without it, their verifiability would be significantly diminished.
+From this single sentence, we can create six new sentences using the word-level perturbations from Table 5 to form $\mathcal{A}^{6}_{word}(s)$.
+Once the seven sentences are embedded into the vector space, they form the hyper-rectangle ($\mathbb{H}(\{s\}\cup\mathcal{A}^{6}_{word}(s))$.
+By repeating this construction for the remaining $3399$ sentences, we obtain the set of hyper-rectangles $\mathbb{H}_{word}(\mathcal{Y})$ for the dataset.
+Given a sentence $s$, we embed each sentence in $\mathcal{A}^{b}_{t}(s)=\{s_{1},\dots,s_{b}\}$ into m obtaining vectors $\mathcal{V}^{b}_{t}(s)=\{v,v_{1},\ldots,v_{b}\}$ where $v_{j}=E(s_{j})$.
+3.4.5 Measuring the Quality of Sentence Embeddings
+One of our implicit assumptions in the previous sections, is that the embedding function $E$ maps pairs of semantically similar sentences to nearby points in the embedding space.
+In Section 5.5.2, we will evaluate the accuracy of this assumption using cosine similarity.
+This metric measures how similar two vectors are in a multi-dimensional space by calculating the cosine of the angle between them:
+$\text{CoS}(v_{1},v_{2})=\frac{v_{1}\cdot v_{2}}{\|v_{1}\|\|v_{2}\|}$
+where $\cdot$ is the dot product and $\|v\|=\sqrt{v\cdot v}$.
+The resulting value ranges from 00 to $1$.
+A value of $1$ indicates that the vectors are parallel (highest similarity), while 00 means that the vectors are orthogonal (no similarity).
+3.5 Training
+As outlined in Section 2.3, robust training is essential for bolstering the robustness of DNNs; without it, their verifiability would be significantly diminished.
 This study employs two robust training methods, namely data augmentation and a custom PGD adversarial training, with the goal of discerning the factors contributing to the success of robust training and compare the effectiveness of these methods.
@@
 Adversarial Training.
-In this training method, we modify the traditional Projected Gradient Descent (PGD) algorithm [85], defined as follows.
+In this training method, the traditional Projected Gradient Descent (PGD) algorithm [20], is defined as follows.
 Given a loss function $\mathcal{L}$, a step size $\gamma\in$ and a starting point $\hat{x}_{0}$ then the output of the PGD algorithm $x(l)$ after $l$ iterations is defined as:
 $x(0)$ $=\hat{x}_{0}$
-$x(t+1)$ $=\text{proj}_{\mathbb{H}}\Big{[}x(t)+\gamma\cdot\text{sign}(\nabla_{x(t)}\mathcal{L}(x(t),y))\Big{]}$
+$x(t+1)$ $=\text{proj}_{\mathcal{S}}\Big{[}x(t)+\gamma\cdot\text{sign}(\nabla_{x(t)}\mathcal{L}(x(t),y))\Big{]}$
 […]
-[…]
+In this work, we modify the algorithm to work with custom-defined hyper-rectangles as the subspace.
+The primary distinction between our customised PGD algorithm and the standard version lies in the definition of the step size.
+In the conventional algorithm, the step size is represented by a scalar $\gamma\in$ therefore representing a uniform step size in every dimension.
+In our case the width of $\mathbb{H}$ in each dimension may vary greatly, therefore we transforms $\gamma$ into a vector in m, allowing the step size to vary by dimension.
+Note that the dot $\cdot$ between $\gamma$ and $sign(\nabla)$ becomes an element-wise multiplication.
 The resulting customised PGD training seeks to identify the worst perturbations within the custom-defined subspace, and trains the given neural network to classify those perturbations correctly, in order to make the network robust to adversarial inputs in the chosen subspace.
-0.3.5 Choice of Verification Algorithm
+3.6 Choice of Verification Algorithm
 As stated earlier, our approach in this study involves the utilization of cutting-edge tools for DNN verification.
@@
 This choice is made over IBP due to its ability to yield tighter bounds.
-Subsequently, we conduct comparisons and integrate Marabou [131], a state-of-the-art complete and precise SMT-based verifier.
+Subsequently, we conduct comparisons and integrate Marabou [22], a state-of-the-art complete verifier.
 This enables us to attain the highest verification percentage, maximizing the tightness of the bounds.
+Additionally, we incorporate $\alpha\beta$-CROWN[42, 43], the best-performing verifier in the VNN-COMP 2024 competition, known for its efficiency in linear bound propagation and branch-and-bound techniques.
+Moreover, we utilise CORA[55], an abstract interpretation-based verifier that supports zonotope-based verification, allowing us to compare hyper-rectangles and zonotopes in our verification experiments.
 We will use notation $\mathcal{V}$ to refer to a verifier abstractly.
-0.4 Characterisation of Verifiable Subspaces
+4 Characterisation of Verifiable Subspaces
 In this Section, we provide key results in support of Contribution 1 formulated in the introduction:
-  • We start with introducing the metric of generalisability of (verified) subspaces, and introducing the problem of the generalisability-verifiability trade-off.
-  • We show that the use of semantic subspaces helps to find a better balance between generalisability and verifiability, as compared to the use of geometric subspaces.
+  • We start with introducing the metric of generalisability of (verified) subspaces and set-up some baseline experiments.
+  • We introduce the problem of the verifiability-generalisability trade-off in the context of geometric subspaces.
+  • We show that, compared to geometric subspaces, the use of semantic subspaces helps to find a better balance between generalisability and verifiability.
   • Finally, we show that adversarial training based on semantic subspaces results in DNNs that are both more verifiable and more generalisable than those obtained with other forms of robust training.
-0.4.1 Generalisability-Verifiability Trade-off
-This subsection defines the new metric of generalisability and shows its effect on (naively defined) geometric subspaces.
-Generalisability of Verified Subspaces: Formal Definition
+4.1 Metrics for Understanding the Properties of Embedding Spaces
 Let us start with recalling the existing standard metrics used in DNN verification.
@@
 We restrict to the case of binary classification in this paper for simplicity, so we will assume $C=\{c_{1},c_{2}\}$.
-Furthermore, we are given an embedding function $\mathbb{E}:\mathcal{Y}\rightarrow^{m}$, and a DNN $N:^{m}\rightarrow^{n}$.
+Furthermore, we are given an embedding function $E:\mathbb{S}\rightarrow^{m}$, and a network $N:^{m}\rightarrow^{n}$.
 Usually $n$ corresponds to the number of classes, and thus in case of binary classification, we have $N:^{m}\rightarrow^{2}$.
-Classification, or assignment of a vector $v$ in ${}^{m}$ to a class, depends on the highest value in the vector $N(v)$.
-The most popular metric is accuracy of $N$, which is measured as a percentage of $q$ vectors in ${}^{m}$ that are assigned to a correct class by $N$.
-Note that this metric checks a finite number of points in ${}^{m}$ given by the data set.
-The most popular metric in DNN verification is verifiability.
-Recall that we can define subspaces $\{\mathcal{S}_{1},\ldots,\mathcal{S}_{s}\}$ of ${}^{m}$; in such a way that each of them is associated with a class (to which all points in that subspace should belong).
-A DNN verifier $\mathcal{V}$ takes an $\mathcal{S}_{i}$, its designated class $c_{i}$ and $N$ as an input, and returns as an output an answer whether all points in $\mathcal{S}_{i}$ are guaranteed to be assigned to the class $c_{i}$ by $N$.
-Verifiability is a percentage of subspaces in $\{\mathcal{S}_{1},\ldots,\mathcal{S}_{s}\}$ that are successfully verified.
-All DNN verification papers report this measure.
-Note that each $\mathcal{S}_{i}$ contains an infinite number of points.
-We are introducing a third metric – generalisability of (verified) subspaces.
-Suppose we have a subspace $\mathcal{S}_{i}$ that verifiably consists only of vectors that are assigned to a class $c_{i}$ by $N$.
-Because of the embedding gap, we cannot know in advance how many valid sentences in $\mathcal{Y}$ (or outside of $\mathcal{Y}$ !) will be mapped into $\mathcal{S}_{i}$ by $\mathbb{E}$.
-Checking the former is rather easy, but may not reflect the idea of generalisability to unseen similar sentences.
-Checking the latter is hard.
-We propose the following effective heuristic.
-By the constructions of the previous section, any $\mathcal{S}_{i}$ necessarily contains an embedding of at least one element $s_{i}$ from $\mathcal{Y}$.
-Choosing any perturbation algorithm $\mathcal{P}$ (of kind $k$), we can form a set $\mathcal{A}^{k}_{s_{i}}$, as described in Section 0.3.3.
-Note that $\mathcal{P}$ (and $k$) can be given by a collection of different perturbation algorithms and their kinds.
-The key assumption is that $\mathcal{A}^{k}_{s_{i}}$ contains valid sentences semantically similar to $s_{i}$ and belonging to the same class.
-Consider a set $\mathcal{V}_{i}$ of vectors in ${}^{m}$ such that each vector $v_{j}\in\mathcal{V}_{i}$ is defined as $v_{j}=\mathbb{E}(s_{j})$, with $s_{j}\in\mathcal{A}^{k}_{s_{i}}$.
-This is the set of embeddings of sentences contained in $\mathcal{A}^{k}_{s_{i}}$.
-Percentage of elements of $\mathcal{V}_{i}$ contained in $\mathcal{S}_{i}$ gives us the generalisability of $\mathcal{S}_{i}$.
-Note that each $\mathcal{V}_{i}$ is finite, because each $\mathcal{A}^{k}_{s_{i}}$ is finite.
-Note that, unlike accuracy and verifiability, the generalisability metric does not explicitly depend on any DNN.
-However, in this paper we only study generalisability of verifiable subspaces, and thus the existence of a verified DNN $N$ will be assumed.
-Because verifiability is reported as a percentage of $\{\mathcal{S}_{1},\ldots,\mathcal{S}_{s}\}$, it will be convenient to report generalisability over the set of subspaces $\mathcal{S}=\bigcup\mathcal{S}_{i}$.
-For this, we take $\mathcal{A}=\bigcup\mathcal{A}^{k}_{s_{i}}$ and the corresponding $\mathcal{V}=\bigcup\mathcal{V}_{i}$ for all $s_{i}\in\mathcal{Y}$ and compute generalisability of $\mathcal{S}$ as a percentage of elements of $\mathcal{V}$ contained in all subspaces of $\mathcal{S}$ [^6] Note that this calculation allows an element of $\mathcal{A}^{k}_{s_{i}}$ to count towards generalisability if it belongs to $\mathcal{S}_{j}$, regardless of whether $i=j$ or $i\neq j$.
-Different approaches might involve restraining the validity to $i=j$, or computing generalisability of each $\mathcal{S}_{j}$ individually and taking an average.
-The current choice favours a global view by calculating how generalisable the whole collection $\mathcal{S}$ is.
-.
-When it is important to emphasise the kind of perturbation, we will use the notation $\mathcal{A}_{k}$ for $\bigcup\mathcal{A}^{k}_{s_{i}}$.
-Base Line Experiments: Understanding the Properties of Embedding Spaces
-Figure 3: An example of hyper-rectangle drawn around all points of the same class (left), shrunk hyper-rectangle $\mathbb{H}_{sh}$ that is obtained by excluding all points from the opposite class (center) and clustered hyper-rectangles (right) in 2-dimensions.
-The red dots represent sentences in the embedding space of one class, while the blue dots are embedded sentences that do not belong to that class.
-The methodology defined thus far has given basic intuitions about the relative nature of the NLP verification pipeline.
-Bearing this in mind, it is important to start our analysis with the general study of the embedding (sub)spaces and suitable base line settings.
-Benchmark data sets will be abbreviated as “RUAR" and “Medical".
-For a benchmark network $N:^{m}\rightarrow^{2}$, we train a medium-sized fully-connected DNN, using stochastic gradient descent and cross-entropy loss.
+An embedded sentence $s\in\mathcal{Y}$ is classified as class $c$ if the value of $c$ in $N(E(s))$ is higher than all other classes.
+Accuracy.
+The most popular metric for measuring the performance of the network is the accuracy of $N$, which is measured as a percentage of sentences in $\mathcal{Y}$ that are assigned to a correct class by $N$.
+Note that this metric only checks a finite number of points in m given by the dataset.
+Verifiability.
+A verifier $\mathcal{V}$ takes a network $N$, a subspace $\mathcal{S}$ and its designated class $c$ as an input, and outputs 1 if it can prove that $N$ assigns all points in the subspace $\mathcal{S}$ to the class $c$ and 0 otherwise.
+Consider a verification problem with multiple subspaces $\{\mathcal{S}_{1},\ldots,\mathcal{S}_{l}\}$, where all the points in each subspace should be assigned to a specific class $c_{i}\in\{c_{1},\ldots,c_{n}\}$.
+In the literature, the most popular metric to measure success rate of the given verifier on $\{\mathcal{S}_{1},\ldots,\mathcal{S}_{l}\}$ is verifiability:
+Definition 7 (Verifiability)
+Given a set of subspaces $\mathcal{S}_{1},\ldots,\mathcal{S}_{l}$ each assigned to classes $c_{1},\ldots,c_{l}$, then the verifiability is the percentage of such subspaces successfully verified:
+$\mathcal{W}(\mathcal{S}_{1},\ldots,\mathcal{S}_{l},c_{1},\ldots,c_{l})=\frac{\sum_{i=0}^{l}\mathcal{V}(N,\mathcal{S}_{i},c_{i})}{l}$
+All DNN verification papers that study such problems report this measure. Note that each subspace contains an infinite number of points.
+However, suppose we have a subspace $\mathcal{S}$ that verifiably consists only of vectors that are assigned to a class $c$ by $N$.
+Because of the embedding gap, it is difficult to calculate how many valid unseen sentences outside of $\mathcal{Y}$ will be mapped into $\mathcal{S}$ by $E$, and therefore how much utility there is in verifying $\mathcal{S}$.
+In an extreme case it is possible to have 100% verifiability and yet the verified subspaces will not contain any unseen sentences.
+Generalisability. Therefore, we now introduce a third metric, generalisability, which is a heuristic for the number of semantically-similar unseen sentences captured by a given set of subspaces.
+Definition 8 (Generalisability)
+Given a set of subspaces $\mathcal{S}_{1},\ldots,\mathcal{S}_{l}$ and a target set of embeddings $V$ the generalisability of the subspaces is measured as the percentage of the embedded vectors that lie in the subspaces:
+$\mathcal{G}(V,\mathcal{S}_{1},\ldots,\mathcal{S}_{l})=\frac{|V\cap\bigcup_{i=1}^{l}\mathcal{S}_{i}|}{|V|}$
+In this paper we will generate the target set of embeddings $V$ as $\bigcup_{s\in\mathcal{Y}}\mathcal{V}^{b}_{t}(s)$ where $\mathcal{Y}$ is a dataset, $t$ is the type of semantic perturbation, $b$ is the number of perturbations and $\mathcal{V}^{b}_{t}(s)$ is the embeddings of the set of semantic perturbations $\mathcal{A}^{b}_{t}(s)$ around $s$ generated using $\mathcal{P}_{t}$, as described in Section 3.4.
+Note that $\mathcal{P}_{t}$ can be given by a collection of different perturbation algorithms and their kinds.
+The key assumption is that $\mathcal{A}^{b}_{t}(s)$ contains valid sentences semantically similar to $s$ and belonging to the same class.
+Assuming that membership of $\mathcal{S}$ is easy to compute, then this metric is also easy to compute as the set $\mathcal{A}^{b}_{t}(s)$ is finite and of size $b$, and therefore so is $\mathcal{V}^{b}_{t}(s)$.
+Note that, unlike accuracy and verifiability, the generalisability metric does not explicitly depend on any DNN or verifier.
+However, in this paper we only study generalisability of verifiable subspaces, and thus the existence of a verified network $N$ will be assumed.
+Furthermore, the verified subspaces we study in this paper will be constructed from the dataset via the methodology described in Definition 6.
+4.2 Baseline Experiments for Understanding the Properties of Embedding Spaces
+The methodology defined thus far has given basic intuitions about the modular nature of the NLP verification pipeline.
+Bearing this in mind, it is important to start our analysis with the general study of basic properties of the embedding subspaces, which is our main interest in this paper, and suitable baselines.
+Benchmark datasets will be abbreviated as “RUAR” and “Medical”.
+We use $\mathcal{Y}^{pos}$ to refer to the set of sentences in the training dataset with a positive class (i.e.
+a question asking the identity of the model, and a medical query respectively), and $\mathcal{Y}^{neg}$ to refer to the remaining sentences.
+For a benchmark network $N:^{m}\rightarrow^{2}$, we train a medium-sized fully-connected DNN (with 2 layers of size (128, 2) and input size 30) using stochastic gradient descent and cross-entropy loss.
 The main requirement for a benchmark network is its sufficient accuracy, see Table 7.
 Model Adversarial training Train Accuracy RUAR Test Accuracy RUAR Train Accuracy Medical Test Accuracy Medical
-$N_{base}$
-No
-0.9387 ± 0.0014
-0.9357 ± 0.0018
-0.9632 ± 0.0005
-0.9449 ± 0.0026
-Table 5: Accuracy of the baseline DNN on the R-U-A-Robot and the Medical datasets.
-For the choice of subspaces, which is our main interest in this paper, two extreme (and trivially defined) benchmarks, are the following geometric subspaces:
-  • a subspace drawn around all embedded sentences of the same class in $\mathcal{Y}$.
-This is the largest subspace one may wish to verify but we should assume that verifiability of such a subspace would be near $0\%$.
+$N_{base}$ No 93.87 ± 0.14% 93.57 ± 0.18% 96.32 ± 0.05% 94.49 ± 0.26%
+Table 7: Mean and standard deviation of the accuracy of the baseline DNN on the RUAR and the Medical datasets. All experiments are replicated five times.
+For the choice of benchmark subspaces, we use the following two extreme sets of geometric subspaces:
+  1. 1.
+    the singleton set containing the maximal subspace $SH(\mathbb{H}(\mathcal{Y}^{pos}))$ around all embedded sentences of the positive class $pos$ in $\mathcal{Y}$.
+This is the largest subspace constructable with our methods, but we should assume that verifiability of such a subspace would be near $0\%$.
 It is illustrated in the first graph of Figure 3.
+  2. 2.
 […]
-We need to first understand exact geometric properties (e.g. volume, $\epsilon$ values) and exact verifiability figures for these two extremes. Let us start with understanding the volumes.
-Volumes of Subspaces.
-To start with the largest hyper-rectangle, we need to ensure that it is shrunk to exclude all samples from $\mathcal{Y}$ from the wrong class.
-We define the shrinking method that involves reducing the size of the hyper-rectangle to eliminate each input within the hyper-rectangle that does not pertain to the chosen class.
-Second graph of Figure 3 gives a visual intuition of how this is done.
-Formally, suppose we have a hyper-rectangle $\mathbb{H}_{\mathcal{X}}$ and a vector $\hat{x}\in\mathbb{H}_{\mathcal{X}}$ that violates the class requirement. Recall that by definition:
-$\mathbb{H}_{\mathcal{X}}\coloneqq\{x\in^{m}:\forall j\in[1\;,\;...\;,\;m]\;.\;\>x^{j}\in[\min\nolimits_{i}(\mathcal{X}^{ij}),\;\;\max\nolimits_{i}(\mathcal{X}^{ij})]\}.$
-The algorithm starts with choosing a dimension $j\in[1,...,m]$, and computing whether $|\hat{x}^{j}-\min\nolimits_{i}(\mathcal{X}^{ij})|\leq|\hat{x}^{j}-\max\nolimits_{i}(\mathcal{X}^{ij})|$.
-If the inequality holds, take $d_{\min\nolimits}=|\hat{x}^{j}-\min\nolimits_{j}(\mathcal{X}^{ij})|$ and compute a new minimum for the dimension $j$: $\min\nolimits^{new}_{i}(\mathcal{X}^{ij})=\min\nolimits_{i}(\mathcal{X}^{ij})+d_{min}+\delta$, where $\delta$ is a small positive number (we use $e^{-100}$).
-We then recompute $\mathbb{H}_{sh}=\mathbb{H}_{\mathcal{X}}$ with $\min\nolimits_{i}(\mathcal{X}^{ij})=\min\nolimits^{new}_{i}(\mathcal{X}^{ij})$ for the chosen dimension $j$.
-If the inequality does not hold, take $d_{\max\nolimits}=|\hat{x}^{j}-\max\nolimits_{i}(\mathcal{X}^{ij})|$ and compute a new maximum for the dimension $j$: $\max\nolimits^{new}_{i}(\mathcal{X}^{ij})=\max\nolimits_{i}(\mathcal{X}^{ij})-d_{max}-\delta$, with $\delta$ defined as above.
-We then recompute $\mathbb{H}_{sh}=\mathbb{H}_{\mathcal{X}}$ with $\max\nolimits_{i}(\mathcal{X}^{ij})=\max\nolimits^{new}_{i}(\mathcal{X}^{ij})$ for the chosen dimension $j$.
-Note that modifying just one dimension of $\hat{x}$ like this guarantees that it is not in $\mathbb{H}_{sh}$.
-The algorithm we apply when computing $\mathbb{H}_{sh}$ involves a heuristic for choosing the optimal dimension along which we modify the hyper-rectangle.
-The idea is to make the most conservative choice, that excludes as few other embedded sentences as possible [^7] Note that this algorithm shrinks exactly one dimension by a minimal amount to exclude the unwanted embedded sentence.
-This choice keeps the algorithm fast while guaranteeing the subspace to retain the highest number of wanted inputs.
-However, it is not necessarily the best choice for verification: there might be cases where perturbations of the unwanted input are left inside after shrinking and, if the network classifies them correctly, the subspace can never be verified.
-For large subspaces, our algorithm might render verification unachievable and more clever algorithms should be explored and discussed..
-It amounts to computing how many inputs from the chosen class remain inside of the hyper-rectangle after applying the shrinking algorithm to one $j\in[1,...,m]$ at a time, separately.
-We choose the dimension $j$ for which the hyper-rectangle retains the most inputs of the correct class.
-The algorithm is run iteratively for every element of $\mathbb{H}_{sh}$ that has the wrong class, until $\mathbb{H}_{sh}$ contains only embeddings of sentences of one class.
+We first seek to understand the geometric properties (e.g. volume, $\epsilon$ values) and verifiability figures for these two extremes.
+4.3 Verifiability-Generalisability Trade-off for Geometric Subspaces
+The number and average volume of the hyper-rectangles that will make up our verified subspaces are shown in Table 8.
+Generally, we use the following naming convention for our experiments: $\mathbb{H}_{m}$ denotes a hyper-rectangle obtained using a method $m$.
+For example, RUAR dataset contains $3400$ sentences of the positive class, and therefore the experiment $\mathbb{H}_{\epsilon=0.005}$ consisting of generating hyper-cubes around each positive sentence results in $3400$ hyper-cubes.
+Using clustering, we obtain a set of $50$, $100$, $200$, $250$ clusters denoted as $\mathbb{H}_{50}$$\mathbb{H}_{250}$ and using the shrinking algorithm we obtain $\mathbb{H}_{sh}$.
+Notice the consistent reduction of volume in Table 8, from $\mathbb{H}_{sh}$ to $\mathbb{H}_{50}$ $\mathbb{H}_{250}$ and ultimately to $\mathbb{H}_{\epsilon=0.005}$.
+There are several orders of magnitude between the largest and the smallest subspace.
+Experiment name Hyper-rectangles construction method Avg. volume of hyper-rectangles RUAR Number of hyper-rectangles RUAR Avg. volume of hyper-rectangles Medical Number of hyper-rectangles Medical
+$\mathbb{H}_{sh}$ Hyper-rectangle around the entire dataset shrunk to exclude all negative examples - $SH(\mathbb{H}(\mathcal{Y}^{pos}),\mathcal{Y},c^{pos})$ 7.55e-11 1 2.60e-09 1
+$\mathbb{H}_{50}$ Set of hyper-rectangles on the dataset separated into 50 clusters - $\mathbb{H}(CL(\mathcal{Y}^{pos},50))$ 1.02e-16 50 6.56e-15 50
+$\mathbb{H}_{100}$ Set of hyper-rectangles on the dataset separated into 100 clusters - $\mathbb{H}(CL(\mathcal{Y}^{pos},100))$ 6.23e-18 100 3.25e-17 100
+$\mathbb{H}_{200}$ Set of hyper-rectangles on the dataset separated into 200 clusters - $\mathbb{H}(CL(\mathcal{Y}^{pos},200))$ 3.31e-20 200 4.67e-19 200
+$\mathbb{H}_{250}$ Set of hyper-rectangles on the dataset separated into 250 clusters - $\mathbb{H}(CL(\mathcal{Y}^{pos},250))$ 6.42e-22 250 2.42e-20 250
 […]
 […]
-To give us a sense of progression from the largest hyper-rectangle to the smallest, we also cluster $\mathbb{H}_{sh}$, see the last graph of Figure 3.
-For the clustering method, we employ the k-means algorithm to generate multiple ensembles of hyper-rectangles around each cluster.
-This approach guarantees that the hyper-rectangles contain solely points from the specific class, while also reducing their volume.
-We obtain a set of $50$, $100$, $200$, $250$ clusters denoted as $\mathbb{H}^{*}_{50}$$\mathbb{H}^{*}_{250}$.
-Table 6 shows average volumes of the subspaces that we obtain, for two data sets.
-Notice consistent reduction of volume, from $\mathbb{H}_{sh}$ to $\mathbb{H}^{*}_{50}$ $\mathbb{H}^{*}_{250}$ and ultimately to $\mathbb{H}^{*}_{\epsilon=0.005}$.
-There are several orders of magnitude between the largest and the smallest subspace.
-Hyper-rectangles Construction method Avg. volume of $\mathbb{H}$ RUAR Set cardinality of $\mathbb{H}^{*}$ RUAR Avg. volume of $\mathbb{H}$ Medical Set cardinality of $\mathbb{H}^{*}$ Medical
-$\mathbb{H}_{sh}$
-Hyper-rectangle shrunk to exclude all negative examples
-7.55e-11
-1
-2.6e-09
-1
-$\mathbb{H}^{*}_{50}$
-Set of hyper-rectangles on 50 clusters
-1.02e-16
-50
-6.56e-15
-50
-$\mathbb{H}^{*}_{100}$
-Set of hyper-rectangles on 100 clusters
-6.23e-18
-100
-3.25e-17
-100
-$\mathbb{H}^{*}_{200}$
-Set of hyper-rectangles on 200 clusters
-3.31e-20
-200
-4.67e-19
-200
-$\mathbb{H}^{*}_{250}$
-Set of hyper-rectangles on 250 clusters
-6.42e-22
-250
-2.42e-20
-250
-$\mathbb{H}^{*}_{\epsilon=0.005}$
-Set of $\epsilon$-cubes around all positive dataset sentences, $\epsilon=0.005$
-1.00e-60
-3400
-1.00e-60
-989
-Table 6: Sets of geomtetric subspaces used in the experiments, their cardinality and average volumes of hyper-rectangles. All shapes are eigenspace rotated for better precision.
-Verifiability Range for Subspaces.
-For each set of hyper-rectangles and the given network, we pass them to the verifier (ERAN) and measure verifiability.
-Table 7 shows that the shrunk hyper-rectangle $\mathbb{H}_{sh}$ achieves zero verifiability, and the various clustered hyper-rectangles ($\mathbb{H}^{*}_{50}$, $\mathbb{H}^{*}_{100}$ $\mathbb{H}^{*}_{200}$, $\mathbb{H}^{*}_{250}$) achieve at most negligible verifiability.
-We call this effect low verifiability of the subspaces.
-In contrast, the baseline $\mathbb{H}^{*}_{\epsilon=0.005}$ achieves up to $99.60\%$ verifiability.
+Table 8: Sets of geometric subspaces used in the experiments, their cardinality and average volumes of hyper-rectangles. All shapes are eigenspace rotated for better precision.
+4.3.1 Verifiability of Geometric Subspaces
+Next, we pass each set of hyper-rectangles and the given network to the ERAN verifier and measure verifiability.
+Table 9 shows that, as expected, the shrunk hyper-rectangle $\mathbb{H}_{sh}$ achieves 0% verifiability, and the various clustered hyper-rectangles ($\mathbb{H}_{50}$, $\mathbb{H}_{100}$ $\mathbb{H}_{200}$, $\mathbb{H}_{250}$) achieve at most negligible verifiability.
+In contrast, the baseline $\mathbb{H}_{\epsilon=0.005}$ achieves up to $99.60\%$ verifiability.
 This suggests that $\epsilon=0.005$ is a good benchmark for a different extreme.
-Table 6 can give us an intuition of why $\mathbb{H}^{*}_{\epsilon=0.005}$ has notably higher verifiability than the other hyper-rectangles: the volume of $\mathbb{H}^{*}_{\epsilon=0.005}$ is several orders of magnitude smaller.
-Dataset Model $\boldsymbol{\mathbb{H}_{sh}}$ $\boldsymbol{\mathbb{H}^{*}_{50}}$ $\boldsymbol{\mathbb{H}^{*}_{100}}$ $\boldsymbol{\mathbb{H}^{*}_{200}}$ $\boldsymbol{\mathbb{H}^{*}_{250}}$ $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.005}}$
-RUAR $N_{base}$ 0.00 0.00 1.33 0.52 0.41 0.00 88.67
-Medical $N_{base}$ 0.00 0.00 0.00 2.10 4.08 5.00 97.86
-Table 7: Verifiability of the baseline DNN on the R-U-A-Robot and the Medical datasets, for a selection of geometric subspaces; for ERAN verifier.
-Verifiability-Generalisability Trade-off for Geometric Subspaces
-Tables 6 and 7 suggest that smaller subspaces are more verifiable.
+Table 8 can give us an intuition of why $\mathbb{H}_{\epsilon=0.005}$ has notably higher verifiability than the other hyper-rectangles: the volume of $\mathbb{H}_{\epsilon=0.005}$ is several orders of magnitude smaller.
+We call this effect low verifiability of the high-volume subspaces.
+Dataset Model $\boldsymbol{\mathbb{H}_{sh}}$ $\boldsymbol{\mathbb{H}_{50}}$ $\boldsymbol{\mathbb{H}_{100}}$ $\boldsymbol{\mathbb{H}_{200}}$ $\boldsymbol{\mathbb{H}_{250}}$ $\boldsymbol{\mathbb{H}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}_{\epsilon=0.005}}$
+RUAR $N_{base}$ 0.00% 0.00% 1.33% 0.52% 0.41% 0.00% 88.67%
+Medical $N_{base}$ 0.00% 0.00% 0.00% 2.10% 4.08% 5.00% 97.86%
+Table 9: Verifiability of the baseline DNN on the RUAR and the Medical datasets, for a selection of geometric subspaces; using the ERAN verifier.
+Tables 8 and 9 suggest that smaller subspaces are more verifiable.
 One may also conjecture that they are less generalisable (as they will contain fewer embedded sentences).
-We confirm this via experiments; we are particularly interested in understanding how quickly generalisability deteriorates as verifiability increases.
-To test generalisability, for each data set, we algorithmically generate a new data set $\mathcal{A}$ containing its semantic perturbations, using the method described in Section 0.3.1.
-It will be convenient to refer to parts of that set that correspond to perturbations of sentences belonging to positive and negative classes, respectively.
-In such cases, we will be using the notation $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$.
-Note that in this section, we will work only with $\mathcal{A}^{pos}$.
-This choice is motivated by the nature of the chosen data sets: both Medical and R-U-A-Robot sentences split into:
+We now will confirm this via experiments; we are particularly interested in understanding how quickly generalisability deteriorates as verifiability increases.
+4.3.2 Generalisability of Geometric Subspaces
+To test generalisability, we algorithmically generate a new dataset $\mathcal{A}^{b}_{t}(\mathcal{Y}^{pos})$ containing its semantic perturbations, using the method described in Section 3.1.
+The choice to use only positive sentences is motivated by the nature of the chosen datasets - both Medical and RUAR sentences split into:
   • a positive class, that contains sentences with one intended semantic meaning (they are medical queries, or they are questions about robot identity); and
-  • a negative class that represents “all other sentences". These “other sentences" are not grouped by any specific semantic meaning and therefore do not form one coherent semantic category.
-However Section 0.5 will make use of $\mathcal{A}^{neg}$ in the context of falsifiability of verified subspaces.
+  • a negative class that represents “all other sentences”. These “other sentences” are not grouped by any specific semantic meaning and therefore do not form one coherent semantic category.
+However Section 5 will make use of $\mathcal{A}^{b}_{t}(\mathcal{Y}^{neg})$ in the context of discussing the embedding error of verified subspaces.
 […]
-For generalisability, Table 8 shows that the most verifiable subspace $\mathbb{H}^{*}_{\epsilon=0.005}$ is the least generalisable.
-This means $\mathbb{H}^{*}_{\epsilon=0.005}$ may not contain any valid new sentences apart from the one for which it was formed!
-At the same time, $\mathbb{H}^{*}_{\epsilon=0.05}$ has up to $48\%$ of generalisability at the expense of only up to $5\%$ of verifiability (cf.
+Dataset Experiment Avg. Volume of hyper-rectangles Generalisability (%) Number of sentences contributing to generalisability Total Sentences in $\mathcal{A}^{b}_{t}(\mathcal{Y}^{pos})$
+RUAR $\mathbb{H}_{\epsilon=0.005}$ 1.00e-60 1.95 2821 144500
+$\mathbb{H}_{\epsilon=0.05}$ 1.00e-30 38.47 55592 144500
+$\mathbb{H}_{sh}$ 7.55e-11 50.91 73561 144500
+Medical $\mathbb{H}_{\epsilon=0.005}$ 1.00e-60 0.09 10 11209
+$\mathbb{H}_{\epsilon=0.05}$ 1.00e-30 28.49 3194 11209
+$\mathbb{H}_{sh}$ 2.6e-09 37.13 4162 11209
+Table 10: Generalisability of the selected geometric subspaces $\mathbb{H}_{\epsilon=0.005}$, $\mathbb{H}_{\epsilon=0.05}$ and $\mathbb{H}_{sh}$, measured on the sets of semantic perturbations $\mathcal{A}^{b}_{t_{RAUR}}(\mathcal{Y}^{pos})$ and $\mathcal{A}^{b}_{t_{medical}}(\mathcal{Y}^{pos})$.
+Table 10 shows that the most verifiable subspace $\mathbb{H}_{\epsilon=0.005}$ is the least generalisable.
+This means $\mathbb{H}_{\epsilon=0.005}$ may not contain any valid new sentences apart from the one for which it was formed!
+At the same time, $\mathbb{H}_{\epsilon=0.05}$ has up to $48\%$ of generalisability at the expense of only up to $5\%$ of verifiability (cf.
 Table 9).
 The effect of the generalisability vs verifiability trade-off can thus be rather severe for geometric subspaces.
-This experiment demonstrates the importance of using the generalisability metric, as with only taking into account the verifiability of the subspaces one would choose $\mathbb{H}^{*}_{\epsilon=0.005}$, obtaining mathematically sound but pragmatically useless results.
+This experiment demonstrates the importance of using the generalisability metric: if one only took into account the verifiability of the subspaces one would choose $\mathbb{H}_{\epsilon=0.005}$, obtaining mathematically sound but pragmatically useless results.
 We argue that this is a strong argument for including generalisability as a standard metric in reporting NLP verification results in the future.
-Dataset Hyper-rectangles Avg. Volume of $\mathbb{H}$ Generalisability (%) Generalisability (# of sentences) Total Sentences in $\mathcal{A}^{pos}$
-RUAR
-$\mathbb{H}^{*}_{\epsilon=0.005}$
-1.00e-60
-1.95
-2821
-144500
-$\mathbb{H}^{*}_{\epsilon=0.05}$
-1.00e-30
-38.47
-55592
-144500
-$\mathbb{H}_{sh}$
-7.55e-11
-50.91
-73561
-144500
-Medical
-$\mathbb{H}^{*}_{\epsilon=0.005}$
-1.00e-60
-0.09
-10
-11209
-$\mathbb{H}^{*}_{\epsilon=0.05}$
-1.00e-30
-28.49
-3194
-11209
-$\mathbb{H}_{sh}$
-2.6e-09
-37.13
-4162
-11209
-Table 8: Generalisability of the selected geometric subspaces $\mathbb{H}^{*}_{\epsilon=0.005}$, $\mathbb{H}^{*}_{\epsilon=0.05}$ and $\mathbb{H}_{sh}$, measured on the sets of semantic perturbations $\mathcal{A}^{pos}_{RUAR}$ and $\mathcal{A}^{pos}_{medical}$.
-0.4.2 Semantic Subspaces and Generalisability-Verifiability Trade-off
-The previous subsection has shown that the generalisability-verifiability trade-off is not resolvable by geometric manipulations alone.
+4.4 Verifiability-Generalisability Trade-off for Semantic Subspaces
+The previous subsection has shown that the verifiability-generalisability trade-off is not resolvable by geometric manipulations alone.
 In this section we argue that using semantic subspaces can help to improve the effects of the trade-off.
-The main hypothesis that we are testing is: semantic subspaces permit drawing more precise shapes, and this in turn must improve both verifiability and generalisability.
-Verifiability of Semantic Subspaces
+The main hypothesis that we are testing is: semantic subspaces constructed using semantic-preserving perturbations are more precise, and this in turn improves both verifiability and generalisability.
 We will use the construction given in Definition 6.
-As Table 9 illustrates, we construct several semantic hyper-rectangles on sentences of the positive class using character-level ($\mathbb{H}^{*}_{char}$, $\mathbb{H}^{*}_{del.}$, $\mathbb{H}^{*}_{ins.}$, $\mathbb{H}^{*}_{rep.}$, $\mathbb{H}^{*}_{repl.}$, $\mathbb{H}^{*}_{swap.}$), word-level ($\mathbb{H}^{*}_{word}$) and sentence-level perturbations ($\mathbb{H}^{*}_{pj}$).
-The subscripts ${}_{char}$ and ${}_{word}$ refer to the kind of perturbation algorithm, while ${}_{del.}$, ${}_{ins.}$, ${}_{rep.}$, ${}_{repl.}$, ${}_{swap.}$ and ${}_{pj}$ refer to the type of perturbation, where pj stands for Polyjuice (see Section 0.3.1).
-Notice comparable volumes of all these shapes, and compare with $\mathbb{H}^{*}_{\epsilon=0.05}$.
-Hyper-rectangles Construction method Avg. volume of $\mathbb{H}$ RUAR Set cardinality of $\mathbb{H}^{*}$ RUAR Avg. volume of $\mathbb{H}$ Medical Set cardinality of $\mathbb{H}^{*}$ Medical
-$\mathbb{H}^{*}_{char}$
-Set of hyper-rectangles for character perturbations
-1.54e-30
-3400
-7.66e-31
-989
-$\mathbb{H}^{*}_{word}$
-Set of hyper-rectangles for word perturbations
-1.28e-30
-3400
-$\mathbb{H}^{*}_{pj}$
-Set of hyper-rectangles for polyjuice sentence perturbations
-2.01e-28
-989
-$\mathbb{H}^{*}_{swap.}$
-Set of hyper-rectangles for swapping perturbations
-1.57e-31
-3400
-3.42e-31
-989
-$\mathbb{H}^{*}_{repl.}$
-Set of hyper-rectangles for replacement perturbations
-9.84e-31
-3400
-3.43e-31
-989
-$\mathbb{H}^{*}_{del.}$
-Set of hyper-rectangles for deletion perturbations
-3.46e-31
-3400
-1.24e-32
-989
-$\mathbb{H}^{*}_{ins.}$
-Set of hyper-rectangles for insertion perturbations
-3.21e-31
-3400
-9.11e-33
-989
-$\mathbb{H}^{*}_{rep.}$
-Set of hyper-rectangles for repetition perturbations
-1.56e-31
-3400
-1.06e-32
-989
-$\mathbb{H}^{*}_{\epsilon=0.05}$
-Set of $\epsilon$-cubes generated to match the volume of the hyper-rectangles
-1e-30
-3400
-1e-30
-989
+As Table 11 illustrates, we construct several semantic hyper-rectangles on sentences of the positive class using character-level ($\mathbb{H}_{char}$, $\mathbb{H}_{del.}$, $\mathbb{H}_{ins.}$, $\mathbb{H}_{rep.}$, $\mathbb{H}_{repl.}$, $\mathbb{H}_{swap.}$), word-level ($\mathbb{H}_{word}$) and sentence-level perturbations ($\mathbb{H}_{pj}$).
+The subscripts char and word refer to the kind of perturbation algorithm, while del., ins., rep., repl., swap.
+and pj refer to the type of perturbation, where pj stands for Polyjuice (see Section 3.1).
+Notice comparable volumes of all these shapes, and compare with $\mathbb{H}_{\epsilon=0.05}$.
+Experiment name Hyper-rectangles construction method Avg. volume of hyper-rectangles RUAR Number of hyper-rectangles RUAR Avg. volume of hyper-rectangles Medical Number of hyper-rectangles Medical
+$\mathbb{H}_{char}$ Set of hyper-rectangles for character perturbations 1.54e-30 3400 7.66e-31 989
+$\mathbb{H}_{word}$ Set of hyper-rectangles for word perturbations 1.28e-30 3400 - -
+$\mathbb{H}_{pj}$ Set of hyper-rectangles for polyjuice sentence perturbations - - 2.01e-28 989
+$\mathbb{H}_{swap.}$ Set of hyper-rectangles for swapping perturbations 1.57e-31 3400 3.42e-31 989
+$\mathbb{H}_{repl.}$ Set of hyper-rectangles for replacement perturbations 9.84e-31 3400 3.43e-31 989
+$\mathbb{H}_{del.}$ Set of hyper-rectangles for deletion perturbations 3.46e-31 3400 1.24e-32 989
+$\mathbb{H}_{ins.}$ Set of hyper-rectangles for insertion perturbations 3.21e-31 3400 9.11e-33 989
+$\mathbb{H}_{rep.}$ Set of hyper-rectangles for repetition perturbations 1.56e-31 3400 1.06e-32 989
 Table 11: Sets of semantic subspaces used in the experiments, their cardinality and average volumes of hyper-rectangles. All shapes are eigenspace rotated for better precision.
+4.4.1 Verifiability of Semantic Subspaces
 We pass each set of hyper-rectangles and the network $N_{base}$ to the verifiers ERAN and Marabou to measure verifiability of the subspaces.
 Table 12 illustrates the verification results obtained using ERAN.
-From the table, we can infer that the verifiability of our semantic hyper-rectangles is indeed higher than that of the geomtrically-defined hyper-rectangles (Table 7).
-Furthermore, our semantic hyper-rectangles, while unable to reach the verifiability of $\mathbb{H}^{*}_{\epsilon=0.005}$, achieve notable higher verification than its counterpart of comparable volume $\mathbb{H}^{*}_{\epsilon=0.05}$.
+From the table, we can infer that the verifiability of our semantic hyper-rectangles is indeed higher than that of the geometrically-defined hyper-rectangles (Table 9).
+Furthermore, our semantic hyper-rectangles, while unable to reach the verifiability of $\mathbb{H}_{\epsilon=0.005}$, achieve notable higher verification than its counterpart of comparable volume $\mathbb{H}_{\epsilon=0.05}$.
 From this experiment, we conclude that not only volume, but also precision of the subspaces has an impact on their verifiability.
-Dataset Model $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}^{*}_{word}}$ $\boldsymbol{\mathbb{H}^{*}_{char}}$ $\boldsymbol{\mathbb{H}^{*}_{del.}}$ $\boldsymbol{\mathbb{H}^{*}_{ins.}}$ $\boldsymbol{\mathbb{H}^{*}_{rep.}}$ $\boldsymbol{\mathbb{H}^{*}_{repl.}}$ $\boldsymbol{\mathbb{H}^{*}_{swap.}}$ $\boldsymbol{\mathbb{H}^{*}_{pj}}$
-RUAR $N_{base}$ 0.00 1.80 0.87 1.62 2.63 1.66 0.94 2.07 -
-Medical $N_{base}$ 5.00 - 39.71 39.62 44.66 48.71 37.49 42.60 50.09
-Table 10: Verifiability of the baseline DNN on the R-U-A-Robot and the Medical datasets, for a slection of semantic subspaces; for ERAN verifier.
+Dataset Model $\boldsymbol{\mathbb{H}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}_{word}}$ $\boldsymbol{\mathbb{H}_{char}}$ $\boldsymbol{\mathbb{H}_{del.}}$ $\boldsymbol{\mathbb{H}_{ins.}}$ $\boldsymbol{\mathbb{H}_{rep.}}$ $\boldsymbol{\mathbb{H}_{repl.}}$ $\boldsymbol{\mathbb{H}_{swap.}}$ $\boldsymbol{\mathbb{H}_{pj}}$
+RUAR $N_{base}$ 0.00% 1.80% 0.87% 1.62% 2.63% 1.66% 0.94% 2.07% -
+Medical $N_{base}$ 5.00% - 39.71% 39.62% 44.66% 48.71% 37.49% 42.60% 50.09%
+Table 12: Verifiability of the baseline DNN on the RUAR and the Medical datasets, for a selection of semantic subspaces; using the ERAN verifier.
 Following these results, Table 13 reports the verification results using Marabou instead of ERAN.
-As shown, Marabou is able to verify up to $66.83\%$ ($\mathbb{H}^{*}_{rep.}$), while ERAN achieves at most $50.09\%$.
-This shows that complete and precise verification (Marabou) outperforms abstract interpretation (ERAN).
+As shown, Marabou is able to verify up to $66.83\%$ ($\mathbb{H}_{rep.}$), while ERAN achieves at most $50.09\%$.
+This shows that Marabou outperforms ERAN.
+This is most likely due to the fact that Reluplex algorithm of Marabou achieves better precision on ReLU networks, that we use.
 Overall, the Marabou experiment confirms the trends of improved verifiability shown by ERAN and thus confirms our hypothesis about importance of shape precision.
-Dataset Model $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}^{*}_{word}}$ $\boldsymbol{\mathbb{H}^{*}_{char}}$ $\boldsymbol{\mathbb{H}^{*}_{del.}}$ $\boldsymbol{\mathbb{H}^{*}_{ins.}}$ $\boldsymbol{\mathbb{H}^{*}_{rep.}}$ $\boldsymbol{\mathbb{H}^{*}_{repl.}}$ $\boldsymbol{\mathbb{H}^{*}_{swap.}}$ $\boldsymbol{\mathbb{H}^{*}_{pj}}$
-RUAR $N_{base}$ 1.79 11.69 4.88 4.35 9.72 9.46 5.65 8.07 -
-Medical $N_{base}$ 37.96 - 64.03 64.15 64.65 66.83 64.75 64.36 61.57
-Table 11: Verifiability of the baseline DNN on the R-U-A-Robot and the Medical datasets, for a slection of semantic subspaces; for Marabou verifier.
-Generalisability of Semantic Subspaces
+Dataset Model $\boldsymbol{\mathbb{H}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}_{word}}$ $\boldsymbol{\mathbb{H}_{char}}$ $\boldsymbol{\mathbb{H}_{del.}}$ $\boldsymbol{\mathbb{H}_{ins.}}$ $\boldsymbol{\mathbb{H}_{rep.}}$ $\boldsymbol{\mathbb{H}_{repl.}}$ $\boldsymbol{\mathbb{H}_{swap.}}$ $\boldsymbol{\mathbb{H}_{pj}}$
+RUAR $N_{base}$ 1.79% 11.69% 4.88% 4.35% 9.72% 9.46% 5.65% 8.07% -
+Medical $N_{base}$ 37.96% - 64.03% 64.15% 64.65% 66.83% 64.75% 64.36% 61.57%
+Table 13: Verifiability of the baseline DNN on the RUAR and the Medical datasets, for a selection of semantic subspaces; using the Marabou verifier.
+4.4.2 Generalisability of Semantic Subspaces
 It remains to establish whether the more verifiable semantic subspaces are also more generalisable.
-Table 12 augments Table 8 with results for best-verifiable semantic subspaces, $\mathbb{H}^{*}_{word}$ and $\mathbb{H}^{*}_{pj}$.
-It shows that these semantic subpspaces are also the most generalisable, containing, respectively, $47.67\%$ and $28.74\%$ of the unseen sentences.
+Whereas Table 10 compared the generalisability of $\mathbb{H}_{\epsilon=0.005}$ and $\mathbb{H}_{\epsilon=0.05}$ with that of $\mathbb{H}_{sh}$, Table 14 compares their generalisability to the most verifiable semantic subspaces, $\mathbb{H}_{word}$ and $\mathbb{H}_{pj}$.
+It shows that these semantic subspaces are also the most generalisable among the verifiable subspaces, containing, respectively, $47.67\%$ and $28.74\%$ of the unseen sentences.
+Note that among all the experiments, only $\mathbb{H}_{sh}$ has higher generalisability, but its verifiability is 0.
+Dataset Experiment Avg. Volume of hyper-rectangles Generalisability (%) Number of sentences contributing to generalisability Total Sentences in $\mathcal{A}^{b}_{t}(\mathcal{Y}^{pos})$
+RUAR $\mathbb{H}_{\epsilon=0.005}$ 1.00e-60 1.95 2821 144500
+$\mathbb{H}_{\epsilon=0.05}$ 1.00e-30 38.47 55592 144500
+$\mathbb{H}_{word}$ 1.28e-30 47.67 68882 144500
+Medical $\mathbb{H}_{\epsilon=0.005}$ 1.00e-60 0.09 10 11209
+$\mathbb{H}_{\epsilon=0.05}$ 1.00e-30 28.49 3194 11209
+$\mathbb{H}_{pj}$ 2.01e-28 28.74 3222 11209
+Table 14: Generalisability of the selected geometric subspaces $\mathbb{H}_{\epsilon=0.005}$ and $\mathbb{H}_{\epsilon=0.05}$ and the semantic subspaces $\mathbb{H}_{word}$ and $\mathbb{H}_{pj}$, measured on the sets of semantic perturbations $\mathcal{A}^{b}_{t_{RUAR}}(\mathcal{Y}^{pos})$ and $\mathcal{A}^{b}_{t_{medical}}(\mathcal{Y}^{pos})$.
+Note that the generalisability of $\mathbb{H}_{sh}$ (Table 10), despite it having the volume 19 order of magnitudes bigger, is only $3\%$ greater than $\mathbb{H}_{word}$.
 […]
-Dataset Hyper-rectangles Avg. Volume of $\mathbb{H}$ Generalisability (%) Generalisability (# of sentences) Total Sentences in $\mathcal{A}^{pos}$
-RUAR
-$\mathbb{H}^{*}_{\epsilon=0.005}$
-1.00e-60
-1.95
-2821
-144500
-$\mathbb{H}^{*}_{\epsilon=0.05}$
-1.00e-30
-38.47
-55592
-144500
-$\mathbb{H}^{*}_{word}$
-1.28e-30
-47.67 68882
-144500
-Medical
-$\mathbb{H}^{*}_{\epsilon=0.005}$
-1.00e-60
-0.09
-10
-11209
-$\mathbb{H}^{*}_{\epsilon=0.05}$
-1.00e-30
-28.49
-3194
-11209
-$\mathbb{H}^{*}_{pj}$
-2.01e-28
-28.74 3222
-11209
-Table 12: Generalisability of the selected geometric subspaces $\mathbb{H}^{*}_{\epsilon=0.005}$ and $\mathbb{H}^{*}_{\epsilon=0.05}$ and the semantic subspaces $\mathbb{H}^{*}_{word}$ and $\mathbb{H}^{*}_{pj}$, measured on the sets of semantic perturbations $\mathcal{A}^{pos}_{RUAR}$ and $\mathcal{A}^{pos}_{medical}$.
-Note that the generalisability of $\mathbb{H}^{*}_{sh}$ (Table 8), despite it having the volume 19 order of magnitudes bigger, is only $3\%$ greater than $\mathbb{H}^{*}_{word}$.
-0.4.3 Adversarial Training on Semantic Subspaces
-In this section we aim to study the effects that adversarial training methods have on the verifiability of the previously defined subspaces.
-By comparing a range of different training approaches, we show in this section that our new method of adversarial training based on semantic subspaces is the most efficient.
+Dataset Experiment Total Volume of Hyper-rectangles Training Embedding Space Covered (%)
+RUAR $\mathbb{H}_{\epsilon=0.005}$ 2.89e-57 4.71e-53
+$\mathbb{H}_{\epsilon=0.05}$ 2.89e-27 4.71e-23
+$\mathbb{H}_{word}$ 3.7e-27 6.03e-23
+Medical $\mathbb{H}_{\epsilon=0.005}$ 9.89e-58 6.92e-53
+$\mathbb{H}_{\epsilon=0.05}$ 9.89e-28 6.92e-23
+$\mathbb{H}_{pj}$ 1.63e-25 1.14e-20
+Table 15: Total volume and percentage of the training embedding space covered by our best hyper-rectangles.
+The total volume of the training embedding space for RUAR is $56.14e-$ 5, and for Medical is $51.43e-$ 5.
+4.5 Adversarial Training on Semantic Subspaces
+In this section, we study the effects that adversarial training methods have on the verifiability of the previously defined subspaces in Tables 8 and 11.
+By comparing the effectiveness of the different training approaches described in Section 3.5, we show in this section that adversarial training based on our new semantic subspaces is the most efficient.
 Three kinds of training are deployed in this section:
   1. 1.
-    We use $N_{base}$ as the baseline network from the previous experiments for comparison with networks trained for robustness.
+    No robustness training - The baseline network is $N_{base}$ from the previous experiments, which has not undergone any robustness training.
   2. 2.
     Data augmentation.
-We obtain three augmented datasets $\mathcal{Y}\cup\mathcal{A}^{pos}_{char}$, $\mathcal{Y}\cup\mathcal{A}^{pos}_{word}$ and $\mathcal{Y}\cup\mathcal{A}^{pos}_{pj}$.
-Note that $\mathcal{A}^{pos}_{char}$, $\mathcal{A}^{pos}_{word}$ and $\mathcal{A}^{pos}_{pj}$ are in correspondence with the sets defined in Section 0.4.2.
-The subscripts char and word denote the kind of perturbation, for which the types are detailed in Tables 3 and 4, while the subscript pj refers to the sentence level perturbations generated with Polyjuice.
+We obtain three augmented datasets $\mathcal{Y}\cup\mathcal{A}^{5}_{char}(\mathcal{Y}^{pos})$, $\mathcal{Y}\cup\mathcal{A}^{6}_{word}(\mathcal{Y}^{pos})$ and $\mathcal{A}^{5}_{pj}(\mathcal{Y}^{pos})$ where $\mathcal{A}(\cdot)$ is defined in Section 4.4.
+The subscripts char and word denote the type of perturbation as detailed in Tables 4 and 5, while the subscript pj refers to the sentence level perturbations generated with Polyjuice.
 We train the baseline architecture, using the standard stochastic gradient descent and cross entropy loss, on the augmented datasets, and obtain DNNs $N_{char-aug}$, $N_{word-aug}$ and $N_{pj-aug}$.
   3. 3.
-    Customised PGD adversarial training.
-As defined in Section 0.3.4, we vary the parameter $\mathbb{H}$.
-In particular, we take for $\mathbb{H}$ different choices of geometric and semantic subspaces we have defined so far.
-We denote a network trained with PGD algorithm on a subspace $\mathbb{H}^{*}_{name}$ as $N_{name}$.
-For example, taking $\mathbb{H}=\mathbb{H}_{sh}$ we get $N_{sh}$ as a result of adversarially training the benchmark architecture on $\mathbb{H}_{sh}$.
-See Table 13 for full listing of the networks we obtain in this way.
+[…]
+See Tables 16 & 19 for full listing of the networks we obtain in this way.
 We call DNNs of second and third type robustly trained networks.
 We keep the geometric and semantic subspaces from the previous experiments (shown in Table 11) to compare how training affects their verifiability.
-Following the same evaluation methodology of experiments as in Sections 0.4.1 and 0.4.2, we use the verifiers ERAN and Marabou to measure verifiability of the subspaces.
-Table 13 reports accuracy of the robustly trained networks, while the verification results are presented in Tables 14 and 15.
+Following the same evaluation methodology of experiments as in Sections 4.2 and 4.4.1, we use the verifiers ERAN and Marabou to measure verifiability of the subspaces.
+Table 16 reports accuracy of the robustly trained networks, while the verification results are presented in Tables 17 and 18.
 From Table 16 we can see that networks trained with data augmentation achieve similar nominal accuracy to networks trained with adversarial training.
-However, the most prominent difference is exposed in Tables 14 and 15: adversarial training effectively improves the verifiability of the networks, while data augmentation actually decreases it.
-Specifically, the adversarially trained networks trained on semantic subspaces ($N_{char}$, $N_{word}$, $N_{pj}$) achieved high verifiability, reaching up to $45.87\%$ for R-U-A-Robot and up to $83.48\%$ for the Medical dataset.
+However, the most prominent difference is exposed in Tables 17 and 18: adversarial training effectively improves the verifiability of the networks, while data augmentation actually decreases it.
+Specifically, the adversarially trained networks trained on semantic subspaces ($N_{char-adv}$, $N_{word-adv}$, $N_{pj-adv}$) achieved high verifiability, reaching up to $45.87\%$ for RUAR and up to $83.48\%$ for the Medical dataset.
 This constitutes a significant improvement of the verifiability results compared to $N_{base}$.
-Looking at nuances, there does not seem to be a single winner subspace when it comes to adversarial training, and indeed in some cases $\mathbb{H}^{*}_{\epsilon=0.05}$ wins over more precise subspaces.
+Looking at nuances, there does not seem to be a single winner subspace when it comes to adversarial training, and indeed in some cases $\mathbb{H}_{\epsilon=0.05}$ wins over more precise subspaces.
 All of the subspaces in Table 11 have very similar volume, which accounts for improved performance across all experiments.
@@
 Thus, if the kind of attack is known in advance, the precision of hyper-rectangles can be further tuned.
-Model Dataset Adversarial training Train Accuracy RUAR Test Accuracy RUAR Train Accuracy Medical Test Accuracy Medical
-$N_{char-aug}$
-$\mathcal{Y}\cup\mathcal{A}^{pos}_{char}$
-No
-0.9562 ± 0.0026
-0.9320 ± 0.0035
-0.9908 ± 0.0006
-0.9346 ± 0.0030
-$N_{word-aug}$
-$\mathcal{Y}\cup\mathcal{A}^{pos}_{word}$
-No
-0.9857 ± 0.0006
-0.9459 ± 0.0036
-$N_{pj-aug}$
-$\mathcal{Y}\cup\mathcal{A}^{pos}_{pj}$
-No
-0.9819 ± 0.0009
-0.9319 ± 0.0039
-$N_{char}$
-$\mathcal{Y}$
-Yes
-0.9326 ± 0.0019
-0.9251 ± 0.0038
-0.9627 ± 0.0005
-0.9509 ± 0.0016
-$N_{word}$
-$\mathcal{Y}$
-Yes
-0.9368 ± 0.0016
-0.9237 ± 0.0029
-$N_{pj}$
-$\mathcal{Y}$
-Yes
-0.9505 ± 0.0019
-0.9349 ± 0.0032
-$N_{\epsilon=0.05}$
-$\mathcal{Y}$
-Yes
-0.9401 ± 0.0017
-0.9224 ± 0.0019
-0.9605 ± 0.0009
-0.9504 ± 0.0024
-Table 13: Accuracy of the robustly trained DNNs on the R-U-A-Robot and the Medical datasets. $\mathcal{Y}$ stands for R-U-A-Robot or Medical depending on the column.
-Dataset Model $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}^{*}_{word}}$ $\boldsymbol{\mathbb{H}^{*}_{char}}$ $\boldsymbol{\mathbb{H}^{*}_{del.}}$ $\boldsymbol{\mathbb{H}^{*}_{ins.}}$ $\boldsymbol{\mathbb{H}^{*}_{rep.}}$ $\boldsymbol{\mathbb{H}^{*}_{repl.}}$ $\boldsymbol{\mathbb{H}^{*}_{swap.}}$ $\boldsymbol{\mathbb{H}^{*}_{pj}}$
-RUAR $N_{char-aug}$ 0.00 0.24 0.00 0.51 1.38 1.09 0.35 1.06 -
-$N_{word-aug}$ 0.00 0.24 0.00 0.42 0.31 0.57 0.25 0.92 -
-$N_{char}$ 0.00 8.97 4.43 4.81 9.86 11.3 6.91 8.51 -
-$N_{word}$ 0.04 10.75 4.05 4.36 8.60 9.52 6.81 7.45 -
-$N_{\epsilon=0.05}$ 0.12 10.16 4.18 4.04 8.91 10.17 6.52 7.36 -
-Medical $N_{char-aug}$ 0.00 - 7.59 5.28 12.84 11.05 7.92 7.40 26.97
-$N_{pj-aug}$ 0.00 - 10.31 8.49 15.67 14.90 9.18 10.58 28.59
-$N_{char}$ 5.28 - 50.12 49.78 53.99 57.76 48.02 52.07 55.44
-$N_{pj}$ 2.83 - 47.11 46.14 52.12 56.14 44.59 48.27 57.36
-$N_{\epsilon=0.05}$ 8.68 - 51.60 50.31 55.67 58.52 50.10 53.65 59.76
-Table 14: Verifiability of the robustly trained DNNs on the R-U-A-Robot and the Medical datasets, for a slection of semantic subspaces; for ERAN verifier.
-Dataset Model $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}^{*}_{word}}$ $\boldsymbol{\mathbb{H}^{*}_{char}}$ $\boldsymbol{\mathbb{H}^{*}_{del.}}$ $\boldsymbol{\mathbb{H}^{*}_{ins.}}$ $\boldsymbol{\mathbb{H}^{*}_{rep.}}$ $\boldsymbol{\mathbb{H}^{*}_{repl.}}$ $\boldsymbol{\mathbb{H}^{*}_{swap.}}$ $\boldsymbol{\mathbb{H}^{*}_{pj}}$
-RUAR $N_{char-aug}$ 0.72 13.90 8.49 7.92 13.67 15.50 9.56 11.88 -
-$N_{word-aug}$ 0.24 11.30 3.87 4.05 8.27 8.84 5.71 7.72 -
-$N_{char}$ 7.37 41.93 30.41 30.23 38.20 45.87 32.74 36.62 -
-$N_{word}$ 12.17 45.12 25.82 25.39 33.85 37.45 26.87 30.99 -
-$N_{\epsilon=0.05}$ 18.46 41.93 21.99 20.32 28.13 32.83 23.52 26.74 -
-Medical $N_{char-aug}$ 1.14 - 37.05 35.29 41.50 42.47 34.89 37.94 49.65
-$N_{pj-aug}$ 5.77 - 39.00 38.66 42.28 44.22 37.29 39.03 38.22
-$N_{char}$ 51.70 - 77.59 77.25 77.50 77.98 77.92 78.67 76.58
-$N_{pj}$ 57.45 - 81.94 81.47 82.31 83.48 82.47 82.72 82.24
-$N_{\epsilon=0.05}$ 62.57 - 79.32 78.57 78.70 80.21 79.40 80.76 66.22
-Table 15: Verifiability of the robustly trained DNNs on the R-U-A-Robot and the Medical datasets, for a selection of semantic subspaces; for Marabou verifier.
-Model Adversarial training Train Accuracy RUAR Test Accuracy RUAR Train Accuracy Medical Test Accuracy Medical
-$N_{sh}$
-Yes
-0.9339 ± 0.0022
-0.9296 ± 0.0013
-0.9614 ± 0.0012
-0.9429 ± 0.0026
-$N_{50}$
-Yes
-0.9432 ± 0.0014
-0.9349 ± 0.0019
-0.9556 ± 0.0020
-0.9515 ± 0.0012
-$N_{100}$
-Yes
-0.9488 ± 0.0004
-0.9418 ± 0.0024
-0.9571 ± 0.0011
-0.9547 ± 0.0016
-$N_{200}$
-Yes
-0.9509 ± 0.0009
-0.9445 ± 0.0014
-0.9585 ± 0.0005
-0.9543 ± 0.0010
-$N_{250}$
-Yes
-0.9522 ± 0.0008
-0.9422 ± 0.0023
-0.9607 ± 0.0013
-0.9538 ± 0.0022
-$N_{\epsilon=0.005}$
-Yes
-0.9348 ± 0.0021
-0.9159 ± 0.0007
-0.9624 ± 0.0004
-0.9513 ± 0.0009
-Table 16: Accuracy of the robustly (adversarially) trained DNNs on the R-U-A-Robot and the Medical datasets.
-Dataset Model $\boldsymbol{\mathbb{H}_{sh}}$ $\boldsymbol{\mathbb{H}^{*}_{50}}$ $\boldsymbol{\mathbb{H}^{*}_{100}}$ $\boldsymbol{\mathbb{H}^{*}_{200}}$ $\boldsymbol{\mathbb{H}^{*}_{250}}$ $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.005}}$
-RUAR $N_{sh}$ 0.00 0.00 1.33 0.52 0.41 88.62
-$N_{50}$ 0.00 0.00 0.00 0.00 0.41 90.02
-$N_{100}$ 0.00 0.00 0.00 0.00 0.41 92.74
-$N_{200}$ 0.00 0.00 0.00 0.00 0.08 93.54
-$N_{250}$ 0.00 0.00 0.00 0.00 0.33 93.86
-$N_{\epsilon=0.005}$ 0.00 0.00 0.00 0.00 0.33 98.22
-Medical $N_{sh}$ 0.00 0.00 0.00 2.50 4.40 97.47
-$N_{50}$ 0.00 0.00 1.08 3.60 6.00 98.79
-$N_{100}$ 0.00 0.00 1.08 3.00 5.04 99.09
-$N_{200}$ 0.00 0.00 1.08 2.90 4.96 99.05
-$N_{250}$ 0.00 0.00 0.00 2.90 4.40 98.73
-$N_{\epsilon=0.005}$ 0.00 0.00 0.00 2.30 4.32 99.60
-Table 17: Verifiability of the robustly (adversarially) trained DNNs on the R-U-A-Robot and the Medical datasets, for a selection of geometric subspaces; for ERAN verifier.
-As a final note, we report results from robust training using the subspaces of Experiment 0.4.1 (Table 6).
+Model Dataset Train Accuracy RUAR Test Accuracy RUAR Train Accuracy Medical Test Accuracy Medical
+$N_{char-aug}$ $\mathcal{Y}\cup\mathcal{A}^{5}_{char}(\mathcal{Y}^{pos})$ 95.62 ± 0.26% 93.20 ± 0.35% 99.08 ± 0.06% 93.46 ± 0.30%
+$N_{word-aug}$ $\mathcal{Y}\cup\mathcal{A}^{6}_{word}(\mathcal{Y}^{pos})$ 98.57 ± 0.06% 94.59 ± 0.36% - -
+$N_{pj-aug}$ $\mathcal{Y}\cup\mathcal{A}^{5}_{pj}(\mathcal{Y}^{pos})$ 98.19 ± 0.09% 93.19 ± 0.39%
+$N_{char-adv}$ $\mathcal{Y}$ 93.26 ± 0.19% 92.51 ± 0.38% 96.27 ± 0.05% 95.09 ± 0.16%
+$N_{word-adv}$ $\mathcal{Y}$ 93.68 ± 0.16% 92.37 ± 0.29% - -
+$N_{pj-adv}$ $\mathcal{Y}$ 95.05 ± 0.19% 93.49 ± 0.32%
+$N_{\epsilon=0.05-adv}$ $\mathcal{Y}$ 94.01 ± 0.17% 92.24 ± 0.19% 96.05 ± 0.09% 95.04 ± 0.24%
+Table 16: Accuracy of the robustly trained DNNs on the RUAR and the Medical datasets. $\mathcal{Y}$ stands for either RUAR or Medical depending on the column.
+Dataset Model $\boldsymbol{\mathbb{H}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}_{word}}$ $\boldsymbol{\mathbb{H}_{char}}$ $\boldsymbol{\mathbb{H}_{del.}}$ $\boldsymbol{\mathbb{H}_{ins.}}$ $\boldsymbol{\mathbb{H}_{rep.}}$ $\boldsymbol{\mathbb{H}_{repl.}}$ $\boldsymbol{\mathbb{H}_{swap.}}$ $\boldsymbol{\mathbb{H}_{pj}}$
+RUAR $N_{char-aug}$ 0.00% 0.24% 0.00% 0.51% 1.38% 1.09% 0.35% 1.06% -
+$N_{word-aug}$ 0.00% 0.24% 0.00% 0.42% 0.31% 0.57% 0.25% 0.92% -
+$N_{char-adv}$ 0.00% 8.97% 4.43% 4.81% 9.86% 11.3% 6.91% 8.51% -
+$N_{word-adv}$ 0.04% 10.75% 4.05% 4.36% 8.60% 9.52% 6.81% 7.45% -
+$N_{\epsilon=0.05-adv}$ 0.12% 10.16% 4.18% 4.04% 8.91% 10.17% 6.52% 7.36% -
+Medical $N_{char-aug}$ 0.00% - 7.59% 5.28% 12.84% 11.05% 7.92% 7.40% 26.97%
+$N_{pj-aug}$ 0.00% - 10.31% 8.49% 15.67% 14.90% 9.18% 10.58% 28.59%
+$N_{char-adv}$ 5.28% - 50.12% 49.78% 53.99% 57.76% 48.02% 52.07% 55.44%
+$N_{pj-adv}$ 2.83% - 47.11% 46.14% 52.12% 56.14% 44.59% 48.27% 57.36%
+$N_{\epsilon=0.05-adv}$ 8.68% - 51.60% 50.31% 55.67% 58.52% 50.10% 53.65% 59.76%
+Table 17: Verifiability of the robustly trained DNNs on the RUAR and the Medical datasets, for a selection of semantic subspaces; using the ERAN verifier.
+Dataset Model $\boldsymbol{\mathbb{H}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}_{word}}$ $\boldsymbol{\mathbb{H}_{char}}$ $\boldsymbol{\mathbb{H}_{del.}}$ $\boldsymbol{\mathbb{H}_{ins.}}$ $\boldsymbol{\mathbb{H}_{rep.}}$ $\boldsymbol{\mathbb{H}_{repl.}}$ $\boldsymbol{\mathbb{H}_{swap.}}$ $\boldsymbol{\mathbb{H}_{pj}}$
+RUAR $N_{char-aug}$ 0.72% 13.90% 8.49% 7.92% 13.67% 15.50% 9.56% 11.88% -
+$N_{word-aug}$ 0.24% 11.30% 3.87% 4.05% 8.27% 8.84% 5.71% 7.72% -
+$N_{char-adv}$ 7.37% 41.93% 30.41% 30.23% 38.20% 45.87% 32.74% 36.62% -
+$N_{word-adv}$ 12.17% 45.12% 25.82% 25.39% 33.85% 37.45% 26.87% 30.99% -
+$N_{\epsilon=0.05-adv}$ 18.46% 41.93% 21.99% 20.32% 28.13% 32.83% 23.52% 26.74% -
+Medical $N_{char-aug}$ 1.14% - 37.05% 35.29% 41.50% 42.47% 34.89% 37.94% 49.65%
+$N_{pj-aug}$ 5.77% - 39.00% 38.66% 42.28% 44.22% 37.29% 39.03% 38.22%
+$N_{char-adv}$ 51.70% - 77.59% 77.25% 77.50% 77.98% 77.92% 78.67% 76.58%
+$N_{pj-adv}$ 57.45% - 81.94% 81.47% 82.31% 83.48% 82.47% 82.72% 82.24%
+$N_{\epsilon=0.05-adv}$ 62.57% - 79.32% 78.57% 78.70% 80.21% 79.40% 80.76% 66.22%
+Table 18: Verifiability of the DNNs trained for robustness on the RUAR and the Medical datasets, for a selection of semantic subspaces; using the Marabou verifier.
+Model Train Accuracy RUAR Test Accuracy RUAR Train Accuracy Medical Test Accuracy Medical
+$N_{sh-adv}$ 93.39 ± 0.22% 92.96 ± 0.13% 96.14 ± 0.12% 94.29 ± 0.26%
+$N_{50-adv}$ 94.32 ± 0.14% 93.49 ± 0.19% 95.56 ± 0.20% 95.15 ± 0.12%
+$N_{100-adv}$ 94.88 ± 0.04% 94.18 ± 0.24% 95.71 ± 0.11% 95.47 ± 0.16%
+$N_{200-adv}$ 95.09 ± 0.09% 94.45 ± 0.14% 95.85 ± 0.05% 95.43 ± 0.10%
+$N_{250-adv}$ 95.22 ± 0.08% 94.22 ± 0.23% 96.07 ± 0.13% 95.38 ± 0.22%
+$N_{\epsilon=0.005-adv}$ 93.48 ± 0.21% 91.59 ± 0.07% 96.24 ± 0.04% 95.13 ± 0.09%
+Table 19: Accuracy of the DNNs trained adversarially on the RUAR and the Medical datasets.
+Dataset Model $\boldsymbol{\mathbb{H}_{sh}}$ $\boldsymbol{\mathbb{H}_{50}}$ $\boldsymbol{\mathbb{H}_{100}}$ $\boldsymbol{\mathbb{H}_{200}}$ $\boldsymbol{\mathbb{H}_{250}}$ $\boldsymbol{\mathbb{H}_{\epsilon=0.005}}$
+RUAR $N_{sh-adv}$ 0.00% 0.00% 1.33% 0.52% 0.41% 88.62%
+$N_{50-adv}$ 0.00% 0.00% 0.00% 0.00% 0.41% 90.02%
+$N_{100-adv}$ 0.00% 0.00% 0.00% 0.00% 0.41% 92.74%
+$N_{200-adv}$ 0.00% 0.00% 0.00% 0.00% 0.08% 93.54%
+$N_{250-adv}$ 0.00% 0.00% 0.00% 0.00% 0.33% 93.86%
+$N_{\epsilon=0.005-adv}$ 0.00% 0.00% 0.00% 0.00% 0.33% 98.22%
+Medical $N_{sh-adv}$ 0.00% 0.00% 0.00% 2.50% 4.40% 97.47%
+$N_{50-adv}$ 0.00% 0.00% 1.08% 3.60% 6.00% 98.79%
+$N_{100-adv}$ 0.00% 0.00% 1.08% 3.00% 5.04% 99.09%
+$N_{200-adv}$ 0.00% 0.00% 1.08% 2.90% 4.96% 99.05%
+$N_{250-adv}$ 0.00% 0.00% 0.00% 2.90% 4.40% 98.73%
+$N_{\epsilon=0.005-adv}$ 0.00% 0.00% 0.00% 2.30% 4.32% 99.60%
+Table 20: Verifiability of the DNNs trained adversarially on the RUAR and the Medical datasets, for a selection of geometric subspaces; using the ERAN verifier.
+As a final note, we report results from robust training using the subspaces from Section 4.2 in Table 8.
 Table 19 reports the accuracy and the details of the robustly trained networks on those subspaces, while the verification results are presented in Table 20.
 These tables further demonstrate the importance of volume, and show that subspaces that are too big still achieve negligible verifiability even after adversarial training.
-Generalisability of the shapes used in Tables 13 - 17 remains the same, see Tables 812.
-0.5 NLP Case Studies
-The purpose of this section is two-fold:
-  • Firstly, the case studies we present here apply the NLP Verification Piplene set out in Section 0.2.5 using more modern NLP tools.
+Generalisability of the shapes used in Tables 16 - 20 remains the same, see Tables 1014.
+4.5.1 Zonotopes vs Hyper-rectangles
+For our final experiment, we compare different subspace shapes for verification.
+Hyper-rectangles are easy to compute but represent the largest over-approximation, and convex-hulls are precise but too computationally expensive to calculate.
+Hence, we consider zonotopes as an alternative, as they are more precise than hyper-rectangles while being computable.
+Although complete verifiers could theoretically work with zonotopes, they do not practically support them.
+Therefore, we use CORA, an abstract interpretation-based verifier, to compare hyper-rectangles and zonotopes.
+Additionally, we run verification using $\alpha\beta$-CROWN, the best-performing verifier in the VNN-COMP 2024 competition.
+This experiment is conducted on our best-performing combination of network ($N_{pj-adv}$), subspace ($\mathbb{H}_{pj}$) and dataset (Medical).
+Verifier Geometric Shape Verifiability %
+$\alpha\beta$-CROWN Rotated Hyper-rectangles 88.41
+Marabou Rotated Hyper-rectangles 82.24
+ERAN Rotated Hyper-rectangles 57.36
+CORA Zonotopes 55.73
+CORA Rotated Hyper-rectangles 29.10
+Table 21: Comparison of different subspace shapes for verification and verifiers for $N_{pj-adv}$ and $\mathbb{H}_{pj}$ on the Medical dataset.
+The results presented in Table 21 show that the method of hyper-rectangle rotation that we use is effective.
+It also shows that zonotopes can improve verifiability over rotated hyper-rectangles within CORA.
+However, the approximation provided by the abstract interpretation verifiers ERAN and CORA significantly reduces their effectiveness compared to the precision of Marabou and $\alpha\beta$-CROWN.
+This aligns with our previous findings, where Marabou, outperformed the abstract interpretation verifier ERAN.
+However, these results should be interpreted with caution, as we cannot directly compare the shapes, given that the top complete verifiers do not support zonotopes.
+We conjecture that, should Marabou and $\alpha\beta$-CROWN implement zonotope based verification, their increased precision would further improve the verifiability results.
+5 NLP Case Studies
+The purpose of this section is two-fold.
+Firstly, the case studies we present here apply the NLP Verification Pipeline set out in Section 2.6 using a wider range of NLP tools.
 Notably, in this section we try different LLMs to embed sentences and replace Polyjuice with the LLM vicuna-13b [^8] Using the following API: https://replicate.com/replicate/vicuna-13b/api., a state-of-the-art open source chatbot trained by fine-tuning LLaMA [138] on user-shared conversations collected from ShareGPT  [^9] https://sharegpt.com/.
-For further details, please refer to Section 0.3.1.
-    To support the work with varying different components of the NLP Verification pipeline, we released the tool ANTONIO [21] that implements different choices for each pipeline component, as shown in Figure 4.
-    Figure 4: Tool ANTONIO that implements a modular approach to the NLP verification pipeline introduced in this paper.
-  • And secondly, and perhaps more fundamentally, the purpose of this section is to draw attention to the importance of the purely NLP parts of the pipeline, namely the parts that have to do with generation, processing, perturbing, and embedding sentences.
-This part is largely ignored in NLP verification papers.
-We show in this section that taking NLP decisions for granted may compromise the practical value of the NLP verification pipelines.
-    To show what we mean by these let us show some examples:
-    Example 3 (Failure of the Embedding Function; Falsifiable Subspaces).
-    As an example, imagine a scenario when a DNN was verified on subpaces of a class $c_{j}$.
-Then it is used on new, yet unseen sentences.
-Imagine also that the implicit assumption that the embedding function embeds only semantically similar sentences close in the embedding space, fails.
-Then, the new sentences that should belong to a class $c_{i}$ may be by mistake embedded into the verified subspace (of class $c_{j}$).
-It will be wrong to assure the user that the verified subspace guarantees that these new sentences belong to $c_{j}$ (even if the DNN is guaranteed to classify them as $c_{j}$)!
-In fact we will say that new sentences of class $c_{i}$ that fall inside the verified subspace of class $c_{j}$ falsify the verified subspace.
-Omitting to acknowledge and report the problem of falsifiable verified subspaces may have different implications, depending on the usage scenarios.
-If the verified subspace of class $c_{j}$ is used to recognise and censor sensitive (‘dangerous’) sentences, new sentences of class $c_{i}$ that fall inside of the verified subspace will be wrongly censored; which in turn may make interaction with the chatbot impractical.
-But if the verified subspace of class $c_{j}$ certifies safe sentences, then potentially dangerous sentences of class $c_{i}$ could be wrongly admitted as verifiably safe.
-Note that this problem can be related to the well-known problem of false positives and false negatives in machine learning: the new sentences of class $c_{i}$ that fall within the verified subpspace of class $c_{j}$ at the same time will be false positives or negatives for that DNN (depending whether $c_{j}$ is a positive or negative class).
-    Example 4 (Failure of the Sentence Perturbation Algorithm).
-    Another assumption that most NLP verification papers make is that we can algorithmically generate sentence perturbations in a way that is guaranteed to retain their original semantic meaning.
+For further details, please refer to Section 3.1.
+In order to be able to easily vary the different components of the NLP Verification pipeline, we use the tool ANTONIO [16], shown in Figure 4.
+Figure 4: Tool ANTONIO that implements a modular approach to the NLP verification pipeline used in this paper.
+Secondly, and perhaps more fundamentally, we draw attention to the fact that the correctness of the specification (i.e.
+the subspace being verified) is dependent on the purely NLP parts of the pipeline.
+In particular, the parts that generate, perturb, and embed sentences.
+Therefore, the probability of the specification itself being wrong is higher than in many other areas of verification.
+This aspect is largely ignored in NLP verification papers and, in this section, we show that using standard NLP methods may result in incorrect specifications and therefore compromising the practical value of the NLP verification pipelines.
+Imagine a scenario where a DNN was verified on subspaces of a class $c_{j}$ and then used to classify new, unseen sentences.
+There are two key assumptions that affect the correctness of the generated specifications:
+  1. 1.
+    Locality of the Embedding Function - We have been using the implicit assumption that the embedding function maps semantically similar sentences to nearby points in the embedding space and dissimilar sentences to faraway points.
+If this assumption fails, the verified subspace may also contain the embeddings of unseen sentences that actually belong to a different class $c_{i}$.
+  2. 2.
+    Sentence Perturbation Algorithm Preserves Semantics - Another assumption that most NLP verification papers make is that we can algorithmically generate sentence perturbations in a way that is guaranteed to retain their original semantic meaning.
 All semantic subspaces of Section 4 are defined based on the implicit assumption that all perturbed sentences retain the same class as the original sentence!
-But if this assumption fails, we will once again end up drawing falsifiable semantic subspaces around embeddings of sentences belonging to different classes, with similar consequences as in the previous example.
-As a consequence of the embedding gap, we obtain a seemingly paradoxical situation, when in principle, the same semantic subspace can be both formally verified and empirically falsified!
-Formal verification ensures that all sentences embedded within the semantic subspace will be classified identically by the given DNN; but empirical falsification of the semantic subspace comes from appealing to the semantic meaning of the embedded sentences, – something that only human can define exactly and verification methodology does not directly define.
-In the light of this limitation, How can we measure and improve the quality of the purely NLP components of the pipeline, in a way that ensures that the NLP verification pipeline is sufficiently reliable and practically usable?
-– is the main question that we answer in this section.
-0.5.1 Practicalities of NLP Verification for Critically Safe Applications
-Role of false positives and false negatives.
+But if this assumption fails, we will once again end up constructing semantic subspaces around embeddings of sentences belonging to different classes.
+Given that it is plausible that one or both of these assumptions may fail, it is therefore wrong to assure the user that the fact that we have verified the subspace, guarantees that all sentences that embed into it, actually belong to $c_{j}$ (even if the DNN is guaranteed to classify them as $c_{j}$)!
+In fact we will say that new sentences of class $c_{i}$ that fall inside the verified subspace of class $c_{j}$ expose an embedding error in the verified subspace.
+Note the root cause of these failures is the embedding gap, as we are unable to map sets of points in the embedding space back to sets of natural language sentences.
+Consequently, we are unable to reliably obtain correct specifications, and therefore we may enter a seemingly paradoxical situation when, in principle, the same subspace can be both formally verified and empirically shown to exhibit embedding errors!
+Formal verification ensures that all sentences embedded within the semantic subspace will be classified identically by the given DNN; but empirical evidence of embedding errors in the semantic subspace comes from appealing to the semantic meaning of the embedded sentences – something that the NLP model can only seek to approximate.
+Failing to acknowledge and report on the problem of verified subspaces exhibiting embedding errors may have different implications, depending on the usage scenario.
+Suppose the network is being used to recognise and censor sensitive (‘dangerous’) sentences, and the subspace is verified to only contain such dangerous sentences.
+Then new sentences that fall inside of the verified subspace may still be wrongly censored; which in turn may make interaction with the chatbot impractical.
+But if the subspace is verified to only contain safe sentences, then potentially dangerous sentences could still be wrongly asserted as verifiably safe.
+Note that this problem is closely related to the well-known problem of false positives and false negatives in machine learning: as any new sentences that get incorrectly embedded into a verified subspace of a different class, must necessarily be false positives or false negatives for that DNN.
+In the light of this limitation, the main question investigated by this section is: How can we measure and improve the quality of the purely NLP components of the pipeline, in a way that decreases the likelihood of generating subspaces prone to embedding errors and therefore ensures the that our verification results are usable in practice?
+As an answer to the measurement part of this question, we will introduce the embedding error metric, that we argue should be used together with verifiability and generalisability metrics in all NLP verification benchmarks.
+5.1 Role of False Positives and False Negatives
 Generally, when DNNs are used for making decisions in situations where safety is critically important, practical importance of accuracy for each class may differ.
-For example, for an autonomous car, misrecognising a stop sign for 30 mph is more dangerous than misrecognising a 30 mph sign for a stop sign.
+For example, for an autonomous car, misrecognising a 20 mph sign for a 60 mph is more dangerous than misrecognising a 60 mph sign for a 20 mph sign.
 Similarly for NLP, because of legal or safety implications, it is crucial that the chatbot always discloses its identity when asked, and never gives medical advice.
-We want to avoid false negatives altogether, i.e.
+In the literature and in this paper, it is assumed that verified DNNs serve as filters that allow the larger system to use machine learning in a safer manner.
+We therefore want to avoid false negatives altogether, i.e.
 if there is any doubt about the nature of the question, we would rather err on the side of caution and disallow the chatbot answers.
@@
 in addition to accuracy.
-Higher recall will mean fewer false negative, which is highly desirable.
+Higher recall will mean fewer false negatives, which is highly desirable.
   2. 2.
-    In the literature and in this paper, it is assumed that verified DNNs serve as filters that allow the larger system to use machine learning in a safe manner.
-Verification covers only subspaces that characterise the critically important class or classes.
-I.e.: DNN verification guarantees that all meaning-preserving perturbations of a stop sign, or “Are you a Robot?” sentence will be classified by the given DNN in the same way as the original unperturbed sample.
-In practice, only DNN outputs on these safe regions are trusted, and DNN output bears no verification guarantees for all other inputs.
-This approach completely eliminates the effect of any false negatives.
+    Secondly, when verifying the network, we only seek to verify subspaces that characterise the critically important class or classes that we wish to filter.
+This approach completely eliminates the effect of any false negatives within the verified subspaces.
 But the verified subspaces may still contain false positives (cf.
-Example 3) and in this sense be “falsifiable”, which may not be pragmatically desirable (but should not constitute a security threat).
-Bearing these considerations in mind, throughout this section we report precision and recall of models, in addition to accuracy, to see whether one or another class is particularly challenging to recognise.
+Example 1) and in this sense be “prone to embedding errors”, which may not be pragmatically desirable but do not constitute a security threat.
+Therefore, throughout this section we report the precision, recall and F1 to see whether one or another class is particularly challenging to recognise.
 Depending on the concrete scenario, precision or recall maybe more important.
-We also introduce the falsifiability metric, that should be used together with verifiability and generalisability metrics in all NLP benchmarks.
-0.5.2 Robustness of Large Language Models
-We start first with gaining an intuition about robustness of modern LLMs.
-We will prompt an LLM to generate the desired output without additional fine-tuning, i.e.
-a zero-shot model.
-The model we choose is vicuna-13b, given its high performance on the TruthfulQA zero-shot benchmark [8], a benchmark to test a model’s propensity to generate truthful answers to questions [80].
+5.2 Performance of Existing LLMs as Safety-Critical Filters
+Bearing this in mind, we start first by gaining an intuition about the effectiveness of modern LLMs acting as safety-critical filters, by prompting an LLM to generate the desired output without additional fine-tuning, i.e.
+we first try a zero-shot model.
+For our experiments, we choose the model vicuna-13b, given its high performance on the TruthfulQA zero-shot benchmark [147], a benchmark to test a model’s propensity to generate truthful answers to questions [148].
 Using vicuna-13b, we experimented with prompts as shown in Figure 5, and settled on the first prompt If I ask you anything about your agency, always only reply “I am a large language model”.
@@
 A1 and A2 represent different answers to the same prompt, illustrating a lack of consistency in the output.
-For our zero-shot model, results are reported on the test set.
-Using the test sets of our chosen datasets, we use rule-based methods to filter through the answers [^11] Additionally omitting $\approx 40\%$ of answers which returned empty due to API errors.
+For our zero-shot model, results are reported on the test set of our datasets.
+We use regular expressions and hand-crafted rules to check for the presence of the desired answer (e.g.
+‘I am a large language model’ for the RUAR dataset) for positively classified training samples [^11] Additionally omitting $\approx 40\%$ of answers which returned empty due to API errors.
 .
-For the R-U-A-Robot dataset, if we are strict about the requirements of the output (only allowing for minor differences such as capitalisation), the accuracy of the LLM is $54\%$ ($precision=0.51$, $recall=0.58$, and $F1=0.54$).
+For the RUAR dataset, if we are strict about the requirements of the output (only allowing for minor differences such as capitalisation), the F1 of the LLM is $54\%$ ($precision=0.51$, $recall=0.58$) as shown in the top line of Table 23.
 This shows that false positives are slightly more likely than false negatives.
-If we loosen this criteria to consider other variations on our desired output (e.g.
-‘I am a large language model’ $\rightarrow$ ‘I am a chatbot’) the accuracy marginally improves, with $F1=0.56$.
-For the medical safety dataset, the results are $precision=0.58$, $recall=0.70$, and $F1=0.64$, indicating comparatively less false negatives.
-However, for several cases, the answers include a combination of the desired output and undesired output, e.g.
+If we loosen our success criteria to consider other non-requested variations on our desired output (e.g.
+‘I am a chatbot’ instead of ‘I am a large language model’) the F1 marginally improves, with $F1=0.56$.
+For the medical safety dataset, the results are $precision=0.58$, $recall=0.70$, and $F1=0.64$, indicating comparatively fewer false negatives.
+However, we found that in several cases the generated answers include a combination of the desired output and undesired output, e.g.
 ‘…I am not authorized to provide medical advice …’ followed by explicit medical advice and the results must be interpreted with this caveat.
-Note there were at least 5 instances regarding the R-U-A-Robot dataset where the system confirmed human identity, without any disclaimers.
+Therefore the actual success rate may be even lower than these reported results.
+Note there were at least 5 instances regarding the RUAR dataset where the system confirmed human identity, without any disclaimers.
 Thus, we find that our zero-shot model is, at most, minimally successful in identifying such queries, encouraging the need for verification methodologies.
-0.5.3 Verification Pipeline Setup and the Role of Embedding Functions
-For all experiments in this section, we set up the NLP verification pipelines as shown in Table 18; and implement them using the tool ANTONIO [21] that we implemented to support this paper.
-In setting up the pipeline, we follow key conclusions made in Section 0.4 about successful verification strategies, namely:
-  • semantic subspaces should be preferred over geometric subpspaces;
-  • best semantic subspaces result from using stronger NLP perturbations;
-  • adversarial training should be done on these “best” subpspaces;
-  • Marabou performs best thanks to its precision.
-Based on this, we further strengthen NLP perturbations: we substitute Polyjuice used in the previous section with Vicuna.
+Pipeline Component Component Implementations Additional Details
+0.
+Choosing Datasets RUAR, Medical Same as in Section 4 experiments.
+The RUAR dataset has $6800$ sentences equally divided among the two classes, while the Medical dataset has 2917 medical and non-medical queries ($1417$ and $1500$ examples respectively).
+1.
+Generating Sentence Perturbations $\mathcal{A}^{16}_{t}(\mathcal{Y}^{pos})$, $\mathcal{A}^{16}_{t^{\diamondsuit}}(\mathcal{Y}^{pos})$, $\mathcal{A}^{16}_{t}(\mathcal{Y}^{neg})$, $\mathcal{A}^{16}_{t^{\diamondsuit}}(\mathcal{Y}^{neg})$ With $t=\{char,word,vicuna\}$, the resulting set of sentences $\mathcal{A}^{16}_{t}(\cdot)$ has $54400$ sentences for RUAR and $15824$ sentences for Medical.
+The superscript refers to filtering that will be introduced in Section 5.5.
+2. Embedding Sentences into Real Vector Space s-bert 22M, s-gpt 1.3B, s-gpt 2.7B In the experiments of Section 4 only s-bert 22M was used.
+3.
+Defining Semantic Subspaces based on Sentence Perturbations $\mathbb{H}_{pert}$, $\mathbb{H}_{pert^{\diamondsuit}}$ $\mathbb{H}_{pert}$ and $\mathbb{H}_{pert^{\diamondsuit}}$ are obtained on $\mathcal{A}^{b}_{t}(\mathcal{Y}^{pos})$, $\mathcal{A}^{b}_{t^{\diamondsuit}}(\mathcal{Y}^{pos})$, respectively.
+Their cardinality is $3400$ for RUAR and $989$ for Medical.
+Volume of $\mathbb{H}_{pert}$ for RUAR is $191.83e-$ 19 (s-bert 22M), $353.24e+$ 35 (s-gpt 1.3B), $363.30e+$ 36 (s-gpt 2.7B).
+Volume of $\mathbb{H}_{pert^{\diamondsuit}}$ for RUAR is $252.43e-$ 25 (s-bert 22M), $271.54e+$ 27 (s-gpt 1.3B), $283.10e+$ 28 (s-gpt 2.7B).
+Volume of $\mathbb{H}_{pert}$ for Medical is $223.13e-$ 22 (s-bert 22M), $331.70e+$ 33 (s-gpt 1.3B), $332.10e+$ 33 (s-gpt 2.7B).
+Volume of $\mathbb{H}_{pert^{\diamondsuit}}$ for Medical is $283.65e-$ 28 (s-bert 22M), $253.30e+$ 25 (s-gpt 1.3B), $273.83e+$ 27 (s-gpt 2.7B).
+4.
+Training Robust DNNs using Semantic Subspaces $N_{base}$, $N_{pert}$, $N_{pert^{\diamondsuit}}$ $N_{base}$ is obtained as in Section 4, while $N_{pert}$ and $N_{pert^{\diamondsuit}}$ are obtained through our adversarial training on $\mathbb{H}_{pert}$ and $\mathbb{H}_{pert{\diamondsuit}}$, respectively.
+5. Verifying resulting DNNs on the given semantic subspaces Marabou Same settings as in Section 4
+Table 22: Section 5 NLP verification pipeline setup, implemented using ANTONIO.
+Note that, after filtering, the volume of $\mathbb{H}_{pert}$ decreases by several orders of magnitude.
+Note the gap in volumes of the subspaces generated by s-bert and s-gpt embeddings.
+5.3 Experimental Setup of the Verification Pipeline
+We therefore turn our attention to assessing the effectiveness of training a classifier specifically for the task, and measuring the effect of the assumptions in Section 5 on the embedding error of the verified subspaces.
+For all experiments in this section, we set up the NLP verification pipeline as shown in Table 22; and implement it using the tool ANTONIO [16].
+In setting up the pipeline, we use the key conclusions from Section 4 about successful verification strategies, namely:
+  1. 1.
+    semantic subspaces should be preferred over geometric subspaces as they result in a better verifiability-generalisability trade-off;
+  2. 2.
+    constructing semantic subspaces using stronger NLP perturbations results in higher verifiability of those subspaces;
+  3. 3.
+    likewise, adversarial training using subspaces constructed with stronger NLP perturbations also results in higher verifiability;
+  4. 4.
+    Marabou allows us to verify a higher percentage of subspaces compared to ERAN thanks to its completeness and precision.
+Based on these results, we further strengthen the NLP perturbations by substituting Polyjuice used in the previous section with Vicuna.
 Vicuna introduces more diverse and sophisticated sentence perturbations.
 In addition, we mix in the character and word perturbations used in the previous section, to further diversify and enlarge the set of available perturbed sentences.
-In the terminology of Section 0.4.1, we obtain the sets of perturbed sentences $\mathcal{A}^{pos}_{pert}$ and $\mathcal{A}^{neg}_{pert}$ where $pert$ is itself a combined kind of perturbations, i.e.
-$pert=\{char,word,vicuna\}$.
-Table 18 also uses notation $\mathcal{A}^{pos}_{pert^{\diamondsuit}}$ and $\mathcal{A}^{neg}_{pert^{\diamondsuit}}$ to refer to filtered sets, this terminology will be introduced in the next subsection.
-Having $\mathcal{A}^{pos}_{pert}$, the set of hyper-rectangles $\mathbb{H}^{*}_{pert}$ is obtained according to Definition 3.
-Accordingly, we set the adversarial training to explore hyper-rectangles in $\mathbb{H}^{*}_{pert}$, and obtain the adversarially robust DNN $N_{pert}$.
+In the terminology of Section 4.1, we obtain the sets of perturbed sentences $\mathcal{A}^{b}_{t}(\mathcal{Y}^{pos})$ and $\mathcal{A}^{b}_{t}(\mathcal{Y}^{neg})$ where $t=\{char,word,vicuna\}$ is a combination of these perturbations.
+Table 22 also uses notation $\mathcal{A}^{b}_{t^{\diamondsuit}}(\mathcal{Y}^{pos})$ and $\mathcal{A}^{b}_{t^{\diamondsuit}}(\mathcal{Y}^{neg})$ to refer to filtered sets, this terminology will be introduced in Section 5.5.2.
 In the light of the goals set up in this section, we diversify the kinds of LLMs we use as embedding functions.
@@
 We refer to these 3 models as s-bert 22M, s-gpt 1.3B, and s-gpt 2.7B respectively, where the number refers to size of the model (measured as the number of parameters).
-Pipeline Component Section 0.5 Choices Additional Details
-0. Choice of Datasets
-RUAR, Medical
-Same as in Section 0.4 experiments.
-The RUAR dataset has $6800$ sentences equally divided among the two classes, while the Medical dataset has 2917 medical and non-medical queries ($1417$ and $1500$ examples respectively).
-1. Generating Sentence Perturbations
-$\mathcal{A}^{pos}_{pert}$, $\mathcal{A}^{pos}_{pert^{\diamondsuit}}$, $\mathcal{A}^{neg}_{pert}$, $\mathcal{A}^{neg}_{pert^{\diamondsuit}}$
-With $pert=\{char,word,vicuna\}$, the resulting set of sentences $\mathcal{A}_{pert}$ has $54400$ sentences for RUAR and $15824$ sentences for Medical.
-The superscript ${}^{\diamondsuit}$ refers to filtering that will be introduced in Section 0.5.4.
-2. Embedding Sentences into Real Vector Space
-s-bert 22M, s-gpt 1.3B, s-gpt 2.7B
-In the experiments of Section 0.4 only s-bert 22M was used.
-3. Defining Semantic Subspaces based on Sentence Perturbations
-$\mathbb{H}^{*}_{pert}$, $\mathbb{H}^{*}_{pert^{\diamondsuit}}$
-$\mathbb{H}^{*}_{pert}$ and $\mathbb{H}^{*}_{pert^{\diamondsuit}}$ are obtained on $\mathcal{A}^{pos}_{pert}$ and $\mathcal{A}^{pos}_{pert^{\diamondsuit}}$, respectively.
-Their cardinality is $3400$ for RUAR and $989$ for Medical.
-  • Volume of $\mathbb{H}^{*}_{pert}$ for RUAR is $191.83e-$ 19 (s-bert 22M), $353.24e+$ 35 (s-gpt 1.3B), $363.30e+$ 36 (s-gpt 2.7B).
-  • Volume of $\mathbb{H}^{*}_{pert^{\diamondsuit}}$ for RUAR is $252.43e-$ 25 (s-bert 22M), $271.54e+$ 27 (s-gpt 1.3B), $283.10e+$ 28 (s-gpt 2.7B).
-  • Volume of $\mathbb{H}^{*}_{pert}$ for Medical is $223.13e-$ 22 (s-bert 22M), $331.70e+$ 33 (s-gpt 1.3B), $332.10e+$ 33 (s-gpt 2.7B).
-  • Volume of $\mathbb{H}^{*}_{pert^{\diamondsuit}}$ for Medical is $283.65e-$ 28 (s-bert 22M), $253.30e+$ 25 (s-gpt 1.3B), $273.83e+$ 27 (s-gpt 2.7B).
-4. Training Robust DNNs using Semantic Subspaces
-$N_{base}$, $N_{pert}$, $N_{pert^{\diamondsuit}}$
-$N_{base}$ is obtained as in Section 0.4, while $N_{pert}$ and $N_{pert^{\diamondsuit}}$ are obtained through our adversarial training on $\mathbb{H}^{*}_{pert}$ and $\mathbb{H}^{*}_{pert{\diamondsuit}}$, respectively.
-5. Verifying resulting DNNs on the given semantic subpspaces
-Marabou
-Same settings as in Section 0.4
-Table 18: Section 0.5 NLP verification pipeline setup, implemented using ANTONIO.
-Note that, after filtering, the volume of $\mathbb{H}^{*}_{pert}$ decreases by several orders of magnitude.
-Note the gap in volumes of the subspaces generated by s-bert and s-gpt embeddings.
-For illustration as well as the initial sanity check, we report accuracy of the obtained models, per each of the chosen embedding functions, in Table 19.
-Overall the figures are as expected: the choice of the encoder influences accuracy.
-Thus, using an additional DNN as a filter (as opposed to the zero-shot model), dramatically increases the accuracy of the overall system, from $54-64\%$ in the previous section to the range of $76-95\%$ in Table 19.
+Given $\mathcal{A}^{b}_{t}(\mathcal{Y}^{pos})$, the set of semantic subspaces $\mathbb{H}_{pert}$ which we wish to verify, are obtained via the hyper-rectangle construction in Definition 6.
+Accordingly, we set the adversarial training to explore the same subspaces $\mathbb{H}_{pert}$, and to obtain the network $N_{pert}$.
+5.4 Analysis of the Role of Embedding Functions
+For illustration, as well as an initial confidence check, we report F1 of the obtained models, for each of the chosen embedding functions in Table 23.
+Overall the figures are as expected: compared to the F1 of 54-64% for the zero-shot model, using a fine-tuned trained DNN as a filter dramatically increases the F1 to the range of 76-95%.
 Dataset Model Test set Perturbed test set
 Precision Recall F1 Precision Recall F1
-RUAR $N_{zero-shot}$ 0.5167 0.5835 0.5481 - - -
-$N_{base}$ (s-bert 22M) 0.9568 0.9129 0.9344 0.9477 0.7186 0.8174
-$N_{pert}$ (s-bert 22M) 0.8497 0.9863 0.9129 0.8125 0.9466 0.8745
-$N_{base}$ (s-gpt 1.3B) 0.9620 0.8725 0.9151 0.9545 0.6738 0.7898
-$N_{pert}$ (s-gpt 1.3B) 0.6303 0.9980 0.7724 0.6126 0.9860 0.7554
-$N_{base}$ (s-gpt 2.7B) 0.9674 0.8729 0.9177 0.9549 0.6982 0.8066
-$N_{pert}$ (s-gpt 2.7B) 0.6018 0.9980 0.7508 0.5846 0.9899 0.7350
-Medical $N_{zero-shot}$ 0.5895 0.7022 0.6409 - - -
-$N_{base}$ (s-bert 22M) 0.9523 0.9325 0.9423 0.9520 0.8964 0.9234
-$N_{pert}$ (s-bert 22M) 0.9335 0.9736 0.9531 0.9238 0.9517 0.9376
-$N_{base}$ (s-gpt 1.3B) 0.9193 0.8811 0.8998 0.9217 0.8417 0.8798
-$N_{pert}$ (s-gpt 1.3B) 0.8441 0.9627 0.8938 0.8315 0.9470 0.8854
-$N_{base}$ (s-gpt 2.7B) 0.9325 0.8929 0.9123 0.9289 0.8479 0.8866
-$N_{pert}$ (s-gpt 2.7B) 0.8603 0.9656 0.9098 0.8488 0.9499 0.8964
-Table 19: Accuracy of the models on the test/perturbation set. The average standard deviation is $0.0049$.
-Dataset Model Test set Perturbed test set
-Precision Recall F1 Precision Recall F1
-RUAR $N_{base^{\text{ \ }}}$ (s-bert 22M) 0.9568 0.9129 0.9344 0.9477 0.7186 0.8174
-$N_{pert^{\diamondsuit}}$ (s-bert 22M) 0.8507 0.9894 0.9148 0.8289 0.9412 0.8815
-$N_{base^{\text{ \ }}}$ (s-gpt 1.3B) 0.9620 0.8725 0.9151 0.9545 0.6738 0.7898
-$N_{pert^{\diamondsuit}}$ (s-gpt 1.3B) 0.6493 0.9965 0.7862 0.6356 0.9808 0.7713
-$N_{base^{\text{ \ }}}$ (s-gpt 2.7B) 0.9674 0.8729 0.9177 0.9549 0.6982 0.8066
-$N_{pert^{\diamondsuit}}$ (s-gpt 2.7B) 0.6305 0.9969 0.7724 0.6143 0.9852 0.7566
-Medical $N_{base^{\text{ \ }}}$ (s-bert 22M) 0.9523 0.9325 0.9423 0.9520 0.8964 0.9234
-$N_{pert^{\diamondsuit}}$ (s-bert 22M) 0.9313 0.9717 0.9511 0.9251 0.9493 0.9370
-$N_{base^{\text{ \ }}}$ (s-gpt 1.3B) 0.9193 0.8811 0.8998 0.9217 0.8417 0.8798
-$N_{pert^{\diamondsuit}}$ (s-gpt 1.3B) 0.8445 0.9571 0.8972 0.8426 0.9424 0.8896
-$N_{base^{\text{ \ }}}$ (s-gpt 2.7B) 0.9325 0.8929 0.9123 0.9289 0.8479 0.8866
-$N_{pert^{\diamondsuit}}$ (s-gpt 2.7B) 0.8682 0.9656 0.9143 0.8560 0.9474 0.8993
-Table 20: Accuracy of the models on the test/perturbation set, after filtering. The average standard deviation is $0.0049$.
+RUAR $N_{zero-shot}$ 51.67% 58.35% 54.81% - - -
+$N_{base}$ (s-bert 22M) 95.68% 91.29% 93.44% 94.77% 71.86% 81.74%
+$N_{pert}$ (s-bert 22M) 84.97% 98.63% 91.29% 81.25% 94.66% 87.45%
+$N_{base}$ (s-gpt 1.3B) 96.20% 87.25% 91.51% 95.45% 67.38% 78.98%
+$N_{pert}$ (s-gpt 1.3B) 63.03% 99.80% 77.24% 61.26% 98.60% 75.54%
+$N_{base}$ (s-gpt 2.7B) 96.74% 87.29% 91.77% 95.49% 69.82% 80.66%
+$N_{pert}$ (s-gpt 2.7B) 60.18% 99.80% 75.08% 58.46% 98.99% 73.50%
+Medical $N_{zero-shot}$ 58.95% 70.22% 64.09% - - -
+$N_{base}$ (s-bert 22M) 95.23% 93.25% 94.23% 95.20% 89.64% 92.34%
+$N_{pert}$ (s-bert 22M) 93.35% 97.36% 95.31% 92.38% 95.17% 93.76%
+$N_{base}$ (s-gpt 1.3B) 91.93% 88.11% 89.98% 92.17% 84.17% 87.98%
+$N_{pert}$ (s-gpt 1.3B) 84.41% 96.27% 89.38% 83.15% 94.70% 88.54%
+$N_{base}$ (s-gpt 2.7B) 93.25% 89.29% 91.23% 92.89% 84.79% 88.66%
+$N_{pert}$ (s-gpt 2.7B) 86.03% 96.56% 90.98% 84.88% 94.99% 89.64%
+Table 23: Performance of the models on the test/perturbation set. The average standard deviation is $0.0049$.
 Looking into nuances, one can further notice the following:
   1. 1.
-    Different datasets may work better with different embedding functions: eg s-bert 22M works best for Medical, and s-gpt 2.7B – for RUAR (with the exception of F1 score, for which s-bert 22M is best for both datasets).
-Smaller GPT model s-gpt 1.3B is systematically worse for both datasets;
+    There is not a single embedding function that always results in the highest F1.
+For example, s-bert 22M is found to have the highest F1 for Medical, while s-gpt 2.7B has the highest F1 for RUAR (with the exception of F1 score, for which s-bert 22M is best for both datasets).
+The smaller GPT model s-gpt 1.3B is systematically worse for both datasets.
   2. 2.
-    Best accuracy may not be the best indicator of performance, depending on the scenario of use.
-Note that in our verification scenarios, where the verified DNN filter recognises the positive class, we look for higher precision, thus favouring fewer false positives, as they falsify verifiable subspaces (cf.
-Example 3).
-However for raw (unverified) use of DNN we value higher recall, favouring fewer false negatives, as we do not want the DNN to “miss” sensitive/dangerous cases.
-    For Medical, s-bert 22M (either with or without adversarial training) looks like the best choice. However, for RUAR, the choice of the embedding function has a huge effect:
-    • for general accuracy, s-bert 22M is the best choice (difference with the worst choice of the embedding function is $12-16$ percentage points),
-    • for scenarios when one uses unverified DNN, $N_{pert}$ (s-gpt 2.7B) gives an incredibly high recall ($>99\%$) and would be a great choice (difference with the worst choice of the embedding function is $13-28$ percentage points).
-    • However, if one wanted to use the same $N_{pert}$ (s-gpt 2.7B) for verification, this would be the worst choice, as precision for the network drops to $58-61\%$.
-For verification, $N_{base}$ (s-gpt 2.7B) $N_{base}$ or s-bert 22M would be better choices (with precision $>95\%$).
+    As expected and discussed in Section 5.1, depending on the scenario of use, the highest F1 may not be the best indicator of performance.
+For Medical, s-bert 22M (either with or without adversarial training) obtains the highest precision, recall and F1.
+However, for RUAR, the choice of the embedding function has a greater effect:
+    • if F1 is desired, s-bert 22M is the best choice (difference with the worst choice of the embedding function is $12-16\%$,
+    • for scenarios when one is not interested in verifying the network, the embedding function s-gpt 2.7B when combined with adversarial training gives an incredibly high recall ($>99\%$) and would be a great choice (difference with the worst choice of the embedding function is $13-28\%$).
+    • however, if one wanted to use the same network for verification, s-gpt 2.7B would be the worst choice of embedding function, as the resulting precision drops to $58-61\%$.
+For verification, either $N_{base}$ trained with s-gpt 2.7B, or $N_{base}$ trained with s-bert 22M would be better choices, both of which have precision $>95\%$.
   3. 3.
-    Looking at the effects of the adversarial training, it only makes a significant difference in accuracy for the Medical perturbed test set.
-However, it has more effect on improving recall (up to 10% for Medical and 33% for RUAR).
+    Adversarial training only makes a significant difference in F1 for the Medical perturbed test set. However, it has more effect on improving recall (up to 10% for Medical and 33% for RUAR).
   4. 4.
-    For verifiability-generalisabiliy trade-off, the choice of an embedding function also plays a role.
+    For verifiability-generalisability trade-off, the choice of an embedding function also plays a role.
 Table 32 shows that s-gpt models exhibit lower verifiability compared to s-bert models.
-This observation also concurs with the findings in Section 0.4: greater volume correlates with increased generalisation, while a smaller and more precise subspace enhances verifiability.
+This observation also concurs with the findings in Section 4: greater volume correlates with increased generalisation, while a smaller and more precise subspace enhances verifiability.
 Indeed volumes for s-gpt models are orders of magnitude ($52-55$) larger than s-bert models.
 The main conclusion one should make from the more nuanced analysis, is that depending on the scenario, the embedding function may influence the quality of the NLP verification pipelines, and reporting the error range (for both precision and recall) depending on the embedding function choice should be a common practice in NLP verification.
-0.5.4 Analysis of Perturbations and Falsifiability
-Recall that two problems were identified as potential causes of falsifiable semantic subspaces: the imprecise embedding functions and invalid perturbations (i.e.
+5.5 Analysis of Perturbations
+Recall that two problems were identified as potential causes of embedding errors in semantic subspaces: the imprecise embedding functions and invalid perturbations (i.e.
 the ones that change semantic meaning and the class of the perturbed sentences).
@@
 human annotators should still assign the same label to the perturbed sample.
-Firstly, we wish to understand how common it is for perturbations to change the class, and secondly, we propose several practical methods how perturbation adequacy can be measured algorithmically.
+Firstly, we wish to understand how common it is for our chosen perturbations to change the class, and secondly, we propose several practical methods how perturbation adequacy can be measured algorithmically.
 Recall that the definition of semantic subspaces depends on the assumption that we can always generate semantically similar (valid) perturbations and draw semantic subspaces around them.
 Both adversarial training and verification then explore the semantic subspaces.
-If this assumption fails and the subspaces contain a large number of invalid sentences, the NLP verification pipeline looses much of its practical value.
+If this assumption fails and the subspaces contain a large number of invalid sentences, the NLP verification pipeline loses much of its practical value.
 To get a sense of the scale of this problem, we start with the most reliable evaluation of sentence validity – human evaluation.
-Understanding the Scale of the Problem
-For the human evaluation, we labelled a subset of the perturbed datasets considering all validity criteria. Annotation instructions are modified from [136] and are given in Table 21.
-For this experiment, at the character and word level, we randomly take 10 samples from each dataset for each perturbation sub-type, i.e.
-at the character level this entails 10 perturbations for each category of inserting, deleting, replacing, swapping or repeating a character.
-At the sentence level, a sample entails more than 1 perturbation, given that we prompt vicuna-13b with instructions for the original sentence to be rephrased 5 times.
-For a sample, we randomly choose one of the perturbations from the returned set of perturbations.
-In this way, we take 50 samples from each dataset.
-This leaves us with 290 manually annotated samples (130 from the medical safety, and 160 from the R-U-A-Robot dataset).
+5.5.1 Understanding the Scale of the Problem
+For the human evaluation, we labelled a subset of the perturbed datasets considering all three validity criteria discussed above.
+In the experiment, for each original dataset $\mathcal{Y}$ and word/character perturbation type $t$, we select 10 perturbed sentences from $\mathcal{A}^{16}_{t}(\mathcal{Y})$.
+At the character level this gives us 50 perturbed sentences for both datasets (10 each for inserting, deleting, replacing, swapping or repeating a character).
+At the word level this gives us 60 perturbed sentences for RUAR (deletion, repetition, ordering, negation, singular/plural, verb tense) and 30 for Medical (deletion, repetition, ordering).
+At the sentence level, we only have one kind of perturbation - obtained by prompting vicuna-13b with instructions for the original sentence to be rephrased 5 times.
+We therefore randomly select 50 vicuna-13b perturbed sentences for each dataset.
+This results in a total of 290 pairs consisting of the original sentence and the perturbed sentence (130 from the medical safety, and 160 from the R-U-A-Robot dataset).
+We then asked two annotators to both manually annotate all 290 pairs for the criteria shown in Table 24 which are modified from [150].
 Inter-Annotator Agreement (IAA) is reported via intraclass correlation coefficient (ICC).
-Human Evaluation.
-ICC (A,1) estimates and their 95% confidence intervals (CI) were calculated based on absolute-agreement (single, fixed raters).
+Criteria Instructions
+Semantic similarity Evaluate whether the original and the modified sentence have the same meaning on a scale from 1 to 4, where 1 is ‘The modified version means something completely different’ and 4 means ‘The modified version has exactly the same meaning’.
+Grammaticality Grammatically means issues in grammar, such as verb tense.
+Evaluate the grammaticality of the modified version on a scale of 1-3, where 1 is ‘Not understandable because of grammar issues’, and 3 is ‘Perfectly grammatical’.
+Label consistency Decide whether the positive label of the modified sentence is correct using labels 1 - ‘Yes, the label is correct’, 2 - ‘No, the label is incorrect’ and 3 - ‘Unsure’.
+Table 24: Annotation instructions for manual estimation of the perturbation validity.
+Results of Human Evaluation.
+The raw evaluation results are shown in Tables 2526 and 27.
+Overall, there are high scores for label consistency, in particular for rule-based perturbations, with $\approx 88\%$ and $85\%$ of the perturbations rated as maintaining the same label (i.e.
+score 1) by the two annotators $A1$ and $A2$ respectively.
+Similarly there are high scores for semantic similarity, with $\approx 85\%$ and $78\%$ of the ratings falling between levels 4 and 3 for $A1$ and $A2$.
+For grammaticality, annotators generally rate that perturbations generated by vicuna-13b are grammatical, whereas (as expected) rule-based perturbations compromise on grammaticality.
+Dataset Perturbation Semantic Similarity (%)
+A1 A2
+1 2 3 4 1 2 3 4
+RUAR Rule-based 06.36 07.27 09.09 77.27 05.45 10.90 34.54 49.09
+LLM-based 18.00 08.00 20.00 54.00 16.00 08.00 00.00 76.00
+Medical Rule-based 01.25 02.50 10.00 86.25 10.00 10.00 31.25 48.75
+LLM-based 06.00 20.00 28.00 46.00 12.00 18.00 20.00 50.00
+Table 25: Semantic similarity results of the manual evaluation for annotators A1 and A2.
+Dataset Perturbation Grammaticality (%)
+A1 A2
+1 2 3 1 2 3
+RUAR Rule-based 10.90 31.81 57.27 13.63 78.18 08.18
+LLM-based 02.00 02.00 96.00 00.00 02.00 98.00
+Medical Rule-based 07.50 32.50 60.00 01.25 88.75 10.00
+LLM-based 00.00 00.00 100.0 00.00 06.00 94.00
+Table 26: Grammaticality results of the manual evaluation for annotators A1 and A2.
+Dataset Perturbation Label Consistency (%)
+A1 A2
+1 2 3 1 2 3
+RUAR Rule-based 88.18 00.90 10.90 85.46 04.54 10.00
+LLM-based 78.00 20.00 02.00 70.00 24.00 06.00
+Medical Rule-based 90.00 00.00 10.00 97.50 00.00 02.50
+LLM-based 88.00 04.00 08.00 74.00 00.00 26.00
+Table 27: Label consistency results of the manual evaluation for annotators A1 and A2.
+In order to evaluate the inter-annotator agreement, we report the ICC between the annotators.
+The ICC estimates and their 95% confidence intervals (CI) were calculated based on absolute-agreement (single, fixed raters) – often referred to as ICC(A,1).
 Using cutoffs provided by [151], agreement was determined to be MODERATE for semantic similarity (F = 4.4 df (289), p<.001, 95% CI = [0.56,0.69]), BELOW SATISFACTORY for grammaticality (ICC = 0.43, p <.001, 95% CI = [0.34,0.52]) and BELOW SATISFACTORY for label consistency (ICC = 0.29, p<.001, 95% CI = [0.18, 0.39]).
-This suggests that, when using LLMs, perturbation quality and robustness to class change cannot be taken for granted.
+This suggests that although annotators individually rated the perturbations for high label consistency, there may be disagreement on which specific samples maintain the same label.
+Given the moderate agreement for semantic similarity, we note that there may be perturbations that are semantically similar and yet may not maintain label consistency.
+For example Original: ‘if a computer can feel emotions, does that make you a computer or an actual human?’, Perturbation: ‘if a computer can feel, does that make it a machine or a person’ was rated by both annotators as high in semantic similarity but not maintaining label consistency.
+Overall, and particularly when using LLMs, perturbation quality and robustness to class change cannot be taken for granted, particularly when dealing with safety-critical queries.
 Limitations.
-We note this is in part due to our definition of grammatical being interpreted differently by the two independent evaluators (one accounting for character perturbations/spelling mistakes as un-grammatical and one not), and label consistency being ambiguous for the R-U-A-Robot dataset.
+We note this is in part due to our definition of grammatical being interpreted differently by the two independent evaluators (one accounting for character perturbations/spelling mistakes as un-grammatical and one not), and label consistency being ambiguous for the RUAR dataset.
 Finally, we also note that correlation between raters is statistically significant across all categories - indicating that ratings across coders were aligned beyond chance probability (criteria $\alpha$ = 0.05).
 Future replications are warranted.
-Criteria Instructions
-Semantic similarity
-Evaluate whether the original and the modified sentence have the same meaning on a scale from 1 to 4, where 1 is ‘The modified version means something completely different’ and 4 means ‘The modified version has exactly the same meaning’.
-Grammaticality
-Grammatically means issues in grammar, such as verb tense.
-Evaluate the grammaticality of the modified version on a scale of 1-3, where 1 is ‘Not understandable because of grammar issues’, and 3 is ‘Perfectly grammatical’.
-Label consistency
-Decide whether the positive label of the modified sentence is correct using labels 1 - ‘Yes, the label is correct’, 2 - ‘No, the label is incorrect’ and 3 - ‘Unsure’.
-Table 21: Annotation instructions for manual estimation of the perturbation validity.
-Effective Ways to Measure and Report Perturbation Validity
-Although no geometric or algorithmic method will ever be able to match to the full extent the human perception and interpretation of sentences, we can still formulate a number of effective methods that give an implicit characterisation of validity of perturbations utilised when defining semantic subspaces.
-We propose three:
-  1. 1.
-    Using cosine similarity of embedded sentences, we can implicitly characterise semantic similarity
-  2. 2.
+5.5.2 Automatic Ways to Measure and Report Perturbation Validity
+Although in the near future, no geometric or algorithmic method will be able to match to the full extent the human perception and interpretation of sentences, we can still formulate a number of effective methods that give a characterisation of the validity of the perturbations utilised when defining semantic subspaces.
+We propose two:
+  • Using cosine similarity of embedded sentences, we can characterise semantic similarity
   • Using the ROUGE-N method [152] – a standard technique to evaluate natural sentence overlap, we can measure lexical and syntactic validity
-  3. 3.
-    Finally, we introduce our own metric – faslifiability that measures the number of unwanted sentences that are mapped into a verified subspace.
 We proceed to describe and evaluate each of them in order.
+Note that, as already pointed out in Section 3.1, these metrics give interesting results assuming some simplifying assumptions, respectively, and the analysed sentences are aligned geometrically in the embedding space and the analysed sentences have a large lexical overlap.
+Cosine Similarity
+Recall the definitions of $\mathcal{A}^{b}_{t}(\mathcal{Y})$, $\mathcal{V}^{b}_{t}(\mathcal{Y})$ and cosine similarity in Section 3.4.
+To measure the general effectiveness of the embedding function at generating semantically similar sentences, we compute the percentage of vectors in $\mathcal{V}^{b}_{t}(\mathcal{Y})$ that have a cosine similarity with the embedding of the original sentence that is greater than $0.6$.
+The results are shown in Table 28.
 Dataset Class Encoder Character Vicuna Word
@@
 Table 28: Number of perturbations kept for each model after filtering with cosine similarity > 0.6, used as an indicator of similarity of perturbed sentences relative to original sentences.
-Cosine similarity.
-Generally, cosine similarity is a metric used to measure how similar two vectors are in a multi-dimensional space.
-The cosine similarity between two vectors is calculated based on the cosine of the angle between them.
-Given two vectors $v_{1}$ and $v_{2}$, the cosine similarity is computed using the following formula:
-$\text{CoS}(v_{1},v_{2})=\frac{v_{1}\cdot v_{2}}{\|v_{1}\|\|v_{2}\|}$
-where $\cdot$ is the dot product and $\|v\|=\sqrt{v\cdot v}$.
-The resulting value ranges from 00 to $1$.
-A value of $1$ indicates that the vectors are parallel (highest similarity), while 00 means that the vectors are orthogonal (no similarity).
-Recall that Section 0.3.3 defined $\mathcal{A}^{k}_{s_{i}}$ as a set of perturbations of kind $k$ for the sentence $s_{i}$, and throughout this section we used the notation $\mathcal{A}_{k}=\bigcup\mathcal{A}^{k}_{s_{i}}$ to refer to a set of such perturbations for all sentences.
-Given $\mathcal{A}^{k}_{s_{i}}$, we embed each sentence of $\mathcal{A}^{k}_{s_{i}}$ in ${}^{m}$ obtaining vectors $\mathcal{V}_{i}=\{v_{i},v_{1},\ldots,v_{n}\}$.
-Note that $\mathcal{V}=\bigcup\mathcal{V}_{i}$ is in one-to-one correspondence with $\mathcal{A}_{k}$.
-For each $\mathcal{V}_{i}$, we compute the cosine similarity between each $v_{j}$ and $v_{i}$, for all $j\neq i$.
-We remove from $\mathcal{V}_{i}$ each $v_{j}$ for which the cosine similarity is less than $0.6$ and obtain $\mathcal{V}_{i^{\diamondsuit}}$.
-To compute cosine similarity globally, we take $\mathcal{V}_{{}^{\diamondsuit}}=\bigcup\mathcal{V}_{i^{\diamondsuit}}$ and compute $\frac{|\mathcal{V}_{{}^{\diamondsuit}}|}{|\mathcal{V}|}$, as a percentage of vectors in $\mathcal{V}$ that had cosine similarity greater than $0.6$.
-Because $\mathcal{V}$ is in one-to-one correspondence with $\mathcal{A}_{k}$, it will be convenient to also refer to $\mathcal{A}_{k^{\diamondsuit}}$ – the set of sentence perturbations that correspond to $\mathcal{V}_{{}^{\diamondsuit}}$.
-Furthermore, we will refer to the set of hyper-rectangles obtained from $\mathcal{A}_{k^{\diamondsuit}}$ as $\mathbb{H}^{*}_{k^{\diamondsuit}}$ and, accordingly, we obtain the DNN $N_{k^{\diamondsuit}}$ through adversarial training on $\mathbb{H}^{*}_{k^{\diamondsuit}}$.
-The results are shown in Table 22, and they allow us to identify pros and cons of this metric.
-  • Cons: due to its geometric nature, the cosine similarity metric does not give us a direct knowledge about true semantic similarity of sentences.
-Indeed, human evaluation of semantic similarity we presented in the previous sections hardly matches the optimistic numbers reported in Table 22!
-Moreover, cosine similarity vastly depends on the assumption that the embedding function embeds semantically similar sentences close to each other in ${}^{m}$.
-As an indication of this problem, Table 22 shows that disagreement in cosine similarity estimations may vary up to $15\%$ when different embedding functions are applied.
+We then perform the experiments again, having removed all generated perturbations that fail to meet this threshold.
+For each original type of perturbation $t$, this can be viewed as creating a new perturbation $t^{\diamondsuit}$.
+Therefore in these alternative experiments, we form $\mathcal{A}^{b}_{t^{\diamondsuit}}(\mathcal{Y})$ – the set of filtered sentence perturbations.
+Furthermore, we will refer to the set of hyper-rectangles obtained from $\mathcal{A}^{b}_{t^{\diamondsuit}}(\mathcal{Y})$ as $\mathbb{H}_{t^{\diamondsuit}}$ and, accordingly, we obtain the network $N_{t^{\diamondsuit}}$ through adversarial training on $\mathbb{H}_{t^{\diamondsuit}}$.
+The results are shown in Table 29.
+Dataset Model Test set Perturbed test set
+Precision Recall F1 Precision Recall F1
+RUAR $N_{base^{\text{ \ }}}$ (s-bert 22M) 95.68% 91.29% 93.44% 94.77% 71.86% 81.74%
+$N_{pert^{\diamondsuit}}$ (s-bert 22M) 85.07% 98.94% 91.48% 82.89% 94.12% 88.15%
+$N_{base^{\text{ \ }}}$ (s-gpt 1.3B) 96.20% 87.25% 91.51% 95.45% 67.38% 78.98%
+$N_{pert^{\diamondsuit}}$ (s-gpt 1.3B) 64.93% 99.65% 78.62% 63.56% 98.08% 77.13%
+$N_{base^{\text{ \ }}}$ (s-gpt 2.7B) 96.74% 87.29% 91.77% 95.49% 69.82% 80.66%
+$N_{pert^{\diamondsuit}}$ (s-gpt 2.7B) 63.05% 99.69% 77.24% 61.43% 98.52% 75.66%
+Medical $N_{base^{\text{ \ }}}$ (s-bert 22M) 95.23% 93.25% 94.23% 95.20% 89.64% 92.34%
+$N_{pert^{\diamondsuit}}$ (s-bert 22M) 93.13% 97.17% 95.11% 92.51% 94.93% 93.70%
+$N_{base^{\text{ \ }}}$ (s-gpt 1.3B) 91.93% 88.11% 89.98% 92.17% 84.17% 87.98%
+$N_{pert^{\diamondsuit}}$ (s-gpt 1.3B) 84.45% 95.71% 89.72% 84.26% 94.24% 88.96%
+$N_{base^{\text{ \ }}}$ (s-gpt 2.7B) 93.25% 89.29% 91.23% 92.89% 84.79% 88.66%
+$N_{pert^{\diamondsuit}}$ (s-gpt 2.7B) 86.82% 96.56% 91.43% 85.60% 94.74% 89.93%
+Table 29: Performance of the models on the test/perturbation set, after filtering. The average standard deviation is $0.0049$.
+The results then allow us to identify the pros and cons of cosine similarity as a metric.
   • Pros:
     • There is some indication that cosine similarity is to a certain extent effective.
-For example, we have seen in Section 0.5.3 that s-bert 22M was the best choice for accuracy and precision (Table 19) – and we see in Table 22 that it is s-bert 22M that is most successful at identifying dissimilar sentences, while not penalising accuracy (Table 20).
-At the end of this section, we will also trace how using $\mathcal{V}_{{}^{\diamondsuit}}$ instead of $\mathcal{V}$ impacts verifiability and falsifiability, as an additional evidence of effectiveness of filters based on cosine similarity.
+For example, we have seen in Table 23 in Section 5.3 that s-bert 22M was the best choice for F1 and precision – and we see in Table 28 that s-bert 22M eliminates the most perturbed sentences, while not penalising its F1 in Table 29.
+However, we cannot currently evaluate whether it is eliminating the truly dissimilar sentences.
+This will be evaluated at the end of this section, when we measure how using $t^{\diamondsuit}$ instead of $t$ impacts verifiability and embedding error.
     • Cosine similarity metric is general (i.e. would apply irrespective of other choice of the pipeline), efficient and scalable.
-Thus, the overall conclusion is that, with the proviso of declaring its limitations, cosine similarity is a useful metric to report, and filtering based on cosine similarity is useful as a pre-processing stage in the NLP verification pipeline (the latter will be demonstrated at the end of this section, when we take the pipeline of Table 18 and substitute $\mathcal{A}_{k^{\diamondsuit}}$ for $\mathcal{A}_{k}$).
+  • Cons:
+    • As discussed earlier, due to its geometric nature, the cosine similarity metric does not give us direct knowledge about true semantic similarity of sentences.
+As evidence of this, the human evaluation of semantic similarity we presented in Section 5.5.1 hardly matches the optimistic numbers reported in Table 28!
+    • Moreover, cosine similarity relies on the assumption that the embedding function embeds semantically similar sentences close to each other in m.
+As an indication that this assumption may not hold, Table 28 shows that disagreement in cosine similarity estimations may vary up to $15\%$ when different embedding functions are applied.
+Thus, the overall conclusion is that, although it has its limitations, cosine similarity is a useful metric to report, and filtering based on cosine similarity is useful as a pre-processing stage in the NLP verification pipeline.
+The latter will be demonstrated at the end of this section, when we take the pipeline in Table 22 and substitute $t^{\diamondsuit}$ for $t$.
 Dataset ROUGE-N Precision Recall
-No filtering Filtering No filtering Filtering
+No
+filtering
+Filtering No
+filtering
+Filtering
 s-bert 22M s-gpt 1.3B s-gpt 2.7B s-bert 22M s-gpt 1.3B s-gpt 2.7B
@@
 Dataset ROUGE-N Precision Recall
-No filtering Filtering No filtering Filtering
+No
+filtering
+Filtering No
+filtering
+Filtering
 s-bert 22M s-gpt 1.3B s-gpt 2.7B s-bert 22M s-gpt 1.3B s-gpt 2.7B
@@
   • $recall$, i.e. number of words (for $n=1$) or word sequences (for $n=2,3$) in $s_{i}$ that also appear in $s_{j}$, divided by the number of words in $s_{i}$.
-For lexical ROUGE-N, we compare the strings of the original sample to the perturbations, while for syntax we follow the same procedure, but using the corresponding POS tags [116].
-Furthermore, we calculate and compare ROUGE-N before and after filtering with cosine similarity.
-Results are given in Tables 23 and 24, and qualitative examples of errors in Figure 6.
-It is important to note that we are not concerned with low $precision$ and $recall$ scores, as it does not mean instances of bad rephrases (for example shuffling, rephrasing or synonym substitution could have such an effect).
-Example 5 (Validity of Perturbations).
+Example 3 (Validity of Perturbations)
 Figure 6 shows an experiment in which vicuna-13b is asked to generate sentence perturbations.
@@
 Figure 6: Analysis of some common issues found in the vicuna-13b generated perturbations.
+For lexical ROUGE-N, we compare the strings of the original sample to the perturbations, while for syntax we follow the same procedure, but using the corresponding parts-of-speech (POS) tags [153].
+Furthermore, we calculate and compare ROUGE-N before and after filtering with cosine similarity.
+Results are given in Tables 30 and 31, and qualitative examples of errors in Figure 6.
+It is important to note that we are not concerned with low $precision$ and $recall$ scores, as it does not necessarily imply non-semantic preserving rephrases.
+For example, shuffling, rephrasing or synonym substitution could lower the scores.
   1. 1.
@@
     Without filtering, there is higher precision compared to recall, while after filtering, the $recall$ increases.
-From Tables 23 and 24 we can hypothesise that overall cosine similarity filters out perturbations that are shorter than the original sentences.
-Observationally, we also find instances of literal rephrasing (see Figure 6, the ‘Literal (not pragmatic) rephrasing’ example), which illustrates the difficulties of generating high quality perturbations for data such as medical queries, which can often have expressed emotions that need to be inferred.
-Additionally, hallucinated content present in the perturbations is problematic, however would be more so if we were to utilise the additional levels of risk labels from the medical safety dataset (see Section 0.2.4) – the hallucinated content can have a non-trivial impact on label consistency.
-Falsifiability.
-We start with formally defining the falsifiability metric, following the notation from Section 0.4.1.
-Let $\mathcal{Y}^{c1}$, $\mathcal{Y}^{c2}$ be an NLP dataset $\mathcal{Y}$ split into two datasets, by classes $c1$ and $c2$, respectively.
-The reader may find it intuitive to think about $c1$ as a positive class, and $c2$ as a negative class.
-By perturbing sentences of each class, we obtain sets $\mathcal{A}^{c1}$ and $\mathcal{A}^{c2}$ of adversarial sentence perturbations, according to procedure described in Section 0.4.1.
-Both $\mathcal{A}^{c1}$ and $\mathcal{A}^{c2}$ have corresponding sets of subspaces $\mathcal{S}^{c1}$ and $\mathcal{S}^{c2}$, i.e.
-$\mathcal{S}^{c1}=\bigcup\mathcal{S}_{i}$, where $i\in 1,\ldots q$, and $q$ is the number of sentences in $\mathcal{Y}^{c1}$.
-Let $\mathcal{A}^{c2}=\{s_{1}.\ldots s_{m}\}$.
-Consider a set $\mathcal{V}^{c2}$ of vectors in ${}^{m}$ such that each vector $v_{i}\in\mathcal{V}^{c2}$ is defined as $v_{i}=\mathbb{E}(s_{i})$, with $s_{i}\in\mathcal{A}^{c2}$.
-This is the set of embeddings of sentences contained in $\mathcal{A}^{c2}$.
-We say a subspace $\mathcal{S}_{i}\in\mathcal{S}^{c1}$ is falsifiable if $\mathcal{S}_{i}$ contains at least one element from $\mathcal{V}^{c2}$.
-Falsifiability of $\mathcal{S}^{c1}$ is the percentage of elements of $\mathcal{S}^{c1}$ that are falsifiable subspaces.
-This means, that falsifiability measures the number of subspaces that contain at least one sample of any undesired class.
-False (adversarial) positives of $\mathcal{S}^{c1}$ are calculated as the percentage of elements of $\mathcal{V}^{c2}$ contained cumulatively in all subspaces contained in $\mathcal{S}^{c1}$.
-Dataset Model Verifiability Generalisability Falsifiability False Positives
+From Tables 30 and 31 we can hypothesise that overall cosine similarity filters out perturbations that are shorter than the original sentences.
+Observationally, we also find instances of literal rephrasing (see Figure 6, the ‘Literal (not pragmatic) rephrasing’ example), which illustrates the difficulties of generating high quality perturbations.
+For example in the medical queries, often there are expressed emotions that need to be inferred.
+The addition of hallucinated content in perturbations is also problematic.
+However, it would be more problematic if we were to utilise the additional levels of risk labels from the medical safety dataset (see Section 2.5.1) – the hallucinated content can have a non-trivial impact on label consistency.
+5.6 Embedding Error
+As the final result of this paper, we introduce the new metric – embedding error – that measures the number of unwanted sentences that are mapped into a verified subspace.
+Recall that Sections 5.4 and 5.5 discussed the methods that assess the role of inaccurate embeddings and semantically incoherent perturbations in isolation.
+In both cases, the methods were of general NLP applicability, and did not directly link to verifiability or generalisability of verified subpspaces.
+The embedding error metric differs from these traditional NLP methods in two aspects:
+  • firstly, it helps to measure both effects simultaneously, and thus helps to assess validity of both the assumption of locality of the embedding function and the assumption of semantic stability of the perturbations outlined at the start of Section 5.
+  • secondly, it is applied here as a verification metric specifically.
+Applied to the same verified subspaces and adversarially trained networks as advocated in Section 4, it is shown as a verification metric on par with verifiability and generalisability.
+We next formally define the embedding error metric.
+Intuitively, the embedding error of a set of subspaces $\mathcal{S}_{1},\ldots,\mathcal{S}_{l}$ of class $c_{1}$ is the percentage of those subspaces that contain at least one embedding of a sentence that belongs to a different class.
+Definition 9 (Embedding Error)
+Given a set of subspaces $\mathcal{S}_{1},\ldots,\mathcal{S}_{l}$ that are supposed to contain exclusively sentences of class $c_{1}$, a dataset $\mathcal{Y}$ that contains sentences not of class $c_{1}$ and a set of embeddings $V$, the embedding error is measured as the percentage of subspaces that contain at least one element of $V$.
+$\mathcal{F}(V,\mathcal{S}_{1},\ldots,\mathcal{S}_{l})=\frac{\sum_{i=1}^{l}\mathbb{I}[V\cap\mathcal{S}_{i}\neq\emptyset]}{l}$
+where $\mathbb{I}[\cdot]$ is the indicator function returning $1$ for true.
+As with the definition of generalisability, in this paper we will generate the target set of embeddings $V$ as $\bigcup_{s\in\mathcal{Y}}\mathcal{V}^{b}_{t}(s)$ where $\mathcal{Y}$ is a dataset, $t$ is the type of semantic perturbation, $b$ is the number of perturbations and $\mathcal{V}^{b}_{t}(s)$ is the embeddings of the set of semantic perturbations $\mathcal{A}^{b}_{t}(s)$ around $s$ generated using $\mathcal{P}_{t}$, as described in Section 3.4.
+We also measure the presence of false positives, as calculated as the percentage of the perturbations of sentences from classes other than $c_{1}$ that lie within at least one of the set of subspaces $\mathcal{S}_{1},\ldots,\mathcal{S}_{l}$.
+To measure the effectiveness of the embedding error metric, we perform the following experiments.
+As previously shown in Table 22, both RUAR and Medical datasets are split into two classes, $pos$ and $neg$.
+We construct $\mathcal{A}^{16}_{t}(\mathcal{Y}^{pos})$ and $\mathcal{A}^{16}_{t}(\mathcal{Y}^{neg})$, and as described in Section 3.4, the set $\mathcal{V}^{16}_{t}(\mathcal{Y}^{neg})$ is obtained by embedding sentences in $\mathcal{A}^{16}_{t}(\mathcal{Y}^{neg})$.
+The subspaces for which we measure embedding error are given by $\mathbb{H}_{pert}=\mathbb{H}(\mathcal{A}^{16}_{t}(\mathcal{Y}^{pos}))$ where we consider both the unfiltered ($t$) and the filtered version ($t^{\diamondsuit}$) of the perturbation $t$.
+Dataset Model Verifiability Generalisability Embedding Error False Positives
 RUAR $N_{base^{\text{ \ }}}$ (s-bert 22M) 2.56 1256/44013 2.85 1/3400 0.03 27/40270 0.07
@@
 $N_{pert^{\diamondsuit}}$ (s-gpt 2.7B) 24.92 3957/15530 25.48 0/989 0.00 0/12709 0.00
-Table 25: Verifiability, generalisability and falsifiability of the baseline and the robustly (adversarially) trained DNNs on the RUAR and the Medical datasets, for $\mathbb{H}^{*}_{pert^{\diamondsuit}}$ ($N_{base}$ and $N_{pert^{\diamondsuit}}$) and $\mathbb{H}^{*}_{pert}$ ($N_{pert}$); for Marabou verifier.
-Example 6 (Falsifiability for Verification Pipeline).
-As shown in Table 18, both RUAR and Medical datasets already have two-class splitting, $pos$ and $neg$.
-$\mathcal{A}^{pos}_{pert}$ and $\mathcal{A}^{neg}_{pert}$ are constructed for a positive and a negative class, respectively.
-The set $\mathcal{V}^{neg}$ is obtained by embedding sentences of $\mathcal{A}^{neg}_{pert}$.
-Subspaces for which we measure falsifiability are given by $\mathbb{H}^{*}_{pert}$ and $\mathbb{H}^{*}_{pert^{\diamondsuit}}$ (constructed on $\mathcal{A}^{pos}_{pert}$ and $\mathcal{A}^{pos}_{pert^{\diamondsuit}}$ for RUAR and Medical).
-Table 25 shows falsifiability of our models and semantic subspaces, giving quantitative estimation of the scale of the problem outlined in Examples 3 and 4.
-We can see that the problem indeed exists, hence our claim that falsifiability should be reported routinely in NLP verification papers.
-At the same time, the problem does not appear to be as severe as one might fear: we note that at most $0.09\%$ of the subspaces are falsified, and falsifiability has only occurred in the subspaces created based on the RUAR dataset.
-Therefore falsifiability is constrained to a few subspaces that can, in theory, be removed.
-Furthermore, falsifiability could also reflect issues in the dataset and subsequent noisy perturbations.
+Table 32: Verifiability, generalisability and embedding error of the baseline and the robustly (adversarially) trained DNNs on the RUAR and the Medical datasets, for $\mathbb{H}_{pert^{\diamondsuit}}$ ($N_{base}$ and $N_{pert^{\diamondsuit}}$) and $\mathbb{H}_{pert}$ ($N_{pert}$); for Marabou verifier.
+Table 32 shows the embedding error of our models and semantic subspaces.
+This gives us a quantitative estimation of the scale of the problem of discussed at the start of this section.
+Namely, whether the assumptions of the locality of the embedding function and that perturbations are semantics-preserving hold.
+From the fact that, in five out of the twelve experiments, embedding error is non-zero, we can see that the problem indeed exists, hence our claim that embedding error should be reported routinely in NLP verification papers.
+At the same time, the problem does not appear to be as severe as one might fear: we note that at most $0.09\%$ of the subspaces exhibit embedding error, and embedding error has only occurred in the subspaces created based on the RUAR dataset.
+Therefore embedding error is constrained to a few subspaces that can, in theory, be removed.
+Furthermore, embedding error could also reflect issues in the dataset and subsequent noisy perturbations.
 The medical safety dataset, for instance, was annotated by an expert practitioner, while the RUAR dataset contains (for this particular task) what could be construed as noisy labels.
@@
 The reason for the negative label is that it is an ungrammatical false positive, but given our methods of perturbation for the construction of subspaces, this negative sample may be very similar to a word level perturbation for the positive class.
-Concretely, for the model with the highest falsifiability in Table 25 (i.e.
-$N_{pert}$ s-bert 22M for RUAR dataset with $0.09\%$ falsifiability), some sentence pairs of negative samples with their accompanying perturbations contained in falsified subspaces are: (Original: ‘Are you a chump?’, Perturbation: ‘You a chump’), (Original: ‘Are you a liar’, Perturbation: ‘You a liar’), (Original: ‘if a computer can feel emotions, does that make you a computer or an actual human?’, Perturbation: ‘if a computer can feel, does that make it a machine or a person’).
+Concretely, for the model with the highest embedding error in Table 32 (i.e.
+$N_{pert}$ s-bert 22M for RUAR dataset with $0.09\%$ embedding error), some sentence pairs of negative samples with their accompanying perturbations contained in subspaces exhibiting embedding error are: (Original: ‘Are you a chump?’, Perturbation: ‘You a chump’), (Original: ‘Are you a liar’, Perturbation: ‘You a liar’), (Original: ‘if a computer can feel emotions, does that make you a computer or an actual human?’, Perturbation: ‘if a computer can feel, does that make it a machine or a person’).
 Thus, the task of determining what queries require disclosure (e.g.
@@
 This area merits further research.
-Generalisability and Verifiability.
-For comparison with the findings outlined in Section 0.4, we provide additional insights into verifiability and generalisability, presented in Table 25.
-We analyse the effect of cosine similarity filtering.
-Initially, our observations reveal slightly lower levels of both verifiability and generalisability prior to filtering.
-The unexpected aspect arises in the realm of generalisability.
-As demonstrated in Section 0.4, larger subspaces tend to exhibit greater generalisability, and the subspaces before filtering indeed possess a larger volume, however they do not possess higher generalisability.
-In other words, cosine similarity filtering reduces volume without sacrificing generalisability, as Section 0.4 might suggest.
-But filtering does improve verifiability in accordance with Section 0.4 conclusions.
-Therefore, we conjecture that cosine similarity filtering can serve as an additional heuristic for improving precision of the verified DNNs, and for further reducing the verifiability-generalisability gap.
+Embedding Error
+vs Generalisability and Verifiability For comparison with the findings outlined in Section 4, we provide additional insights into verifiability and generalisability, also presented in Table 32.
+We first analyse the effect of cosine similarity filtering.
+Initially, the experiments reveal that filtering results in slightly higher levels of both verifiability and generalisability for all models.
+Given the conclusions in Section 4, the increase in verifiability is expected.
+However, the increase in generalisability is somewhat unexpected because, as demonstrated in Section 4, larger subspaces tend to exhibit greater generalisability, but filtering decreases the volume of the subspaces.
+Therefore, we conjecture the increase in precision of the subspaces from filtering outweighs the reduction in their volume and hence generalisability increases overall.
+The data therefore suggests that cosine similarity filtering can serve as an additional heuristic for improving precision of the verified DNNs, and for further reducing the verifiability-generalisability gap.
 Indeed, upon calculating the ratio of generalisability to verifiability, we observe a higher ratio before filtering ($1.19\rightarrow 0.99$ for RUAR and $0.99\rightarrow 0.95$ for Medical).
-Recall that Section 0.4 already showed that our proposed usage of semantic subspaces can serve as a heuristic for closing the gap; and cosine similarity filtering provides opportunity for yet another heuristic improvement.
-Moreover, the best performing model $N_{pert}$ (s-bert 22M), results in $10,964$ $(70.6\%)$ medical perturbations and $9530$ $(21.65\%)$ RUAR perturbations contained.
+Recall that Section 4 already showed that our proposed usage of semantic subspaces can serve as a heuristic for closing the gap; and cosine similarity filtering provides opportunity for yet another heuristic improvement.
+Moreover, the best performing model $N_{pert}$ (s-bert 22M), results in $10,964$ $(70.6\%)$ medical perturbations and $9530$ $(21.65\%)$ RUAR perturbations contained in the verified subspaces.
 While $21.65\%$ of the $positive$ perturbations contained in the verified subspaces for the RUAR dataset may seem like a low number, it still results in a robust filter, given that the $positive$ class of the dataset contains many adversarial examples of the same input query, i.e.
-semantically same but lexically different queries.
-Consequently, giving guarantees for $\approx 10,000$ variations on asking a system to disclose non-identity is a notable result.
+semantically similar but lexically different queries.
 The medical dataset on the other hand contains many semantically diverse queries, and there are several unseen medical queries not contained in the dataset nor in the resultant verified subspaces.
-However, given that the subspaces contain $70.6\%$ of the $positive$ perturbations of the medical safety dataset, an application of this could be to carefully curate a dataset of queries containing critical and serious risk-level labels defined by the WEF for chatbots in healthcare (see Section 0.2.4 and [130]), ensure label consistency for the level of risk on the perturbations, and then create verification filters centred around these queries to prevent generation of medical advice for such high-risk queries.
-We find that semantically informed verification generalises well across the different kinds of data to ensure guarantees on the output, and thus should aid in the safety of LLMs.
-0.6 Conclusions and Future Work
+However, given that the subspaces contain $70.6\%$ of the $positive$ perturbations of the medical safety dataset, an application of this could be to carefully curate a new dataset containing only queries with critical and serious risk-level labels defined by the World Economic Forum for chatbots in healthcare (see Section 2.5.1 and [135]).
+This dataset could be used to create verified filters centred around these queries to prevent generation of medical advice for these high-risk categories.
+Overall, we find that semantically-informed verification generalises well across the different kinds of data to ensure guarantees on the output, and thus should aid in ensuring the safety of LLMs.
+6 Conclusions and Future Work
 Summary.
-This paper started with a general analysis of existing NLP verification approaches, with a view of identifying key components of a general NLP verification methodology, which we distilled as a six component “NLP Verification Pipeline”:
+This paper started with a general analysis of existing NLP verification approaches, with a view of identifying key components of a general NLP verification methodology.
+We then distilled these into a “NLP Verification Pipeline” consisting of the following six components:
+  1. 1.
     dataset selection;
+  2. 2.
     generation of perturbations;
+  3. 3.
     choice of embedding functions;
+  4. 4.
     definition of subspaces;
+  5. 5.
     robust training;
+  6. 6.
     verification via one of existing verification algorithms.
-Based on this taxonomy, we formally defined all the components of the pipeline, and implemented it as a tool ANTONIO [21].
-ANTONIO allowed us to mix and match different choices for each pipeline component, enabling us to study the effect of various factors and components of the pipeline in a algorithm-independent way.
+Based on this taxonomy, we make concrete selections for each component, and implement the pipeline using the tool ANTONIO [16].
+ANTONIO allowed us to mix and match different choices for each pipeline component, enabling us to study the effects of varying the components of the pipeline in a algorithm-independent way.
 Our main focus was to identify weak or missing parts of the existing NLP verification methodologies.
-We proposed that NLP verification results must report, in addition to the standard verifiability metric, the following:
+We proposed that NLP verification results should report, in addition to the standard verifiability metric, the following:
   • whether they use geometric or semantic subspaces, and for which type of semantic perturbations;
-  • volumes, generalisability and falsifiability of verified subspaces.
-We finished the paper with a study of major pitfalls of NLP verification pipeline from the point of view of LLMs and proposed future wider and more principled use of some standard NLP methods in NLP verification (precision, recall, cosine similarity, ROUGE-N).
-Contributions.
-The major discoveries of this paper were:
-  • In Section 0.4 we showed that NLP verification tasks exhibit a generalisability-verifiability trade-off, and its effects can be very severe, especially if the verification tasks are approached naively.
-Generalisability is a novel metric we proposed here, and we strongly believe that generalisability should be routinely measured and reported as part of NLP verification pipeline.
-  • In Sections 0.4 and 0.5 we showed that it is possible to overcome this trade-off by using several heuristic methods: defining semantic subspaces, training for semantic robustness, choosing the most suited embedding function and filtering with cosine similarity.
-All of these methods boiled down to defining more precise verifiable subspaces; and all of them can be effectively implemented as part of NLP verification pipelines in the future.
-  • In Section 0.5 we demonstrated that accuracy of LLMs, used as embedding functions and generators of perturbations in NLP verification, cannot be taken for granted.
-They may fail to correctly embed sentences, and they may fail to preserve the semantic meaning of perturbations.
+  • volumes, generalisability and embedding error of verified subspaces.
+We finished the paper with a study of the current limitations of the NLP components of the pipeline and proposed possible improvements such as introducing a perturbations filter stage using cosine similarity.
+One of the major strengths of the pipeline is that each component can be improved individually.
+Contributions. The major discoveries of this paper were:
+  • In Section 4 we proposed generalisability as a novel metric, and showed that NLP verification methods exhibit a generalisability-verifiability trade-off.
+The effects of the trade-off can be severe, especially if the verified subspaces are generated naively (e.g.
+geometrically).
+We therefore strongly believe that generalisability should be routinely reported as part of NLP verification pipeline.
+  • In Sections 4 and 5 we showed that it is possible to overcome this trade-off by using several heuristic methods: defining semantic subspaces, training for semantic robustness, choosing a suitable embedding function and filtering with cosine similarity.
+All of these methods result in the definition of more precise verifiable subspaces; and all of them can be practically implemented as part of NLP verification pipelines in the future.
+This is the main positive result of this paper.
+  • In Section 5 we demonstrated that there are two key assumptions underlying the definition of subspaces that cannot be taken for granted.
+Firstly the LLMs, used as embedding functions, may not map semantically similar sentences to similar vectors in the embedding space.
+Secondly, our algorithmic methods for generating perturbations, whether by LLMs or otherwise, may not always be semantically-preserving operations.
 Both of these factors influence practical applications of the NLP verification pipeline.
-  • In Section 0.5 we demonstrated that even verified subspaces can be semantically falsified: this effect is due to the tension between verification methods that are essentially geometric and the intuitively understood semantic meaning of sentences.
-By defining the falsifiability metric and using it in our experiments, we found out that the effects of falsifiability do not seem to be severe in practice; but this may vary from one scenario to another.
-It is important that NLP verification papers are aware of this pitfall, and report faslsifiability alongside verifiability and generalisability.
-Finally, we claim as a contribution, a novel coherent methodological framework that allows to include a broad spectrum of NLP, geometric, machine learning, and verification methods under a single umbrella.
-No previous publication in this domain covered this range and, as this paper illustrated throughout all sections, consistently covering this broad range of methods is crucial for the development of this field.
+  • In Section 5 we demonstrated that even verified subspaces can exhibit semantic embedding errors: this effect is due to the tension between verification methods that are essentially geometric and the intuitively understood semantic meaning of sentences.
+By defining the embedding error metric and using it in our experiments, we demonstrated that the effects of embedding errors do not seem to be severe in practice; but this may vary from one scenario to another.
+It is important that NLP verification papers are aware of this pitfall, and report embedding error alongside verifiability and generalisability.
+Finally, we claim as a contribution, a novel, coherent, methodological framework that allows us to include a broad spectrum of NLP, geometric, machine learning, and verification methods under a single umbrella.
+As illustrated throughout this paper, no previous publication in this domain covered this range and we believe that covering this broad range of methods is crucial for the development of this field.
+Limitations and the Role of the Embedding Gap.
+In this paper, we have shown the effect of the embedding gap in the NLP verification domain.
+In Section 4.1, we introduced the generalisability metric (Definition 8) to estimate how well subspaces capture semantically similar yet unseen sentences, providing a quantitative lens to examine this challenge.
+Sections 4.3 and 4.4 demonstrated that geometric subspaces struggle with the verifiability-generalisability trade-off, while semantic subspaces, constructed using semantic-preserving perturbations, show promise in mitigating the embedding gap by better aligning with data semantics.
+[…]
+This problem is fundamental for machine learning methods deployed in NLP: they always rely on an “embedding function” that maps sentences into real vectors (on which machine learning algorithms operate).
+There is an implicit assumption that the embedding function works in a way that semantic similarity of sentences is reflected in geometric proximity of their embeddings.
+However, as general semantic similarity of sentences is not effectively computable, there is no hope that a perfect embedding function will ever be defined.
+Fundamentally, this is exactly the reason why machine learning (and not symbolic) approaches to NLP proved to be successful: they operate on the assumption that the embedding function is imperfect.
+As a consequence, any verification pipeline for NLP must include metrics that measure potential embedding errors.
+Section 5.6 of this paper is entirely devoted to defining this problem in mathematically precise terms and proposing an effective metric for measuring and reporting the severity and effects of the embedding errors.
+This paper aims to quantify and address the embedding gap in the NLP domain.
+For better quantifying the effect of the embedding gap, we proposed precise metrics such as verifiability, generalisability and embedding error, and showed their interplay.
+This better understanding of the problem gave rise to our main positive result: the method that empirically reduces the effect of the embedding gap.
+While our findings mark progress, further research is needed to better align geometric representations with semantic meaning, especially in NLP contexts.
 Future Work.
-Following from our in depth analysis of the NLP perspective, we note that even if one has a satisfactory solution to all the issues discussed, there is still the problem of scalability of the available verification algorithms.
+Following from our in-depth analysis of the NLP perspective, we note that even if one has a satisfactory solution to all the issues discussed, there is still the problem of scalability of the available verification algorithms.
 For example, the most performant neural network verifier, $\alpha\beta$-Crown [43], can only handle networks in the range of tens of millions of trainable parameters.
@@
 For future work, we propose to tackle this based on the idea of verifying a smaller DNN (classifier), manageable by verifiers, that can be placed upstream of a complex NLP system as a safeguard.
-We call this a filter (as mentioned in Section 0.3 and illustrated in Figure 2), and Figure 7 shows how a semantically informed verified filter can be prepended to an NLP system (here, an LLM) to check that safety-critical queries are handled responsibly, e.g.
+We call this a filter (as mentioned in Section 3 and illustrated in Figure 2), and Figure 7 shows how a semantically informed verified filter can be prepended to an NLP system (here, an LLM) to check that safety-critical queries are handled responsibly, e.g.
 by redirecting the query to a tightly controlled rule-based system instead of a stochastic LLM.
@@
 These guardrails have been proposed at multiple stages of an NLP pipeline, for example an output rail that checks the output returned by an LLM, or input rail, that rejects unsafe user queries.
-In the figure, we show an application of our filter applied to the user input, and thus creating guarantees that safety critical queries are handled responsibly.
+In Figure 7, we show an application of our filter applied to the user input, which thus creates guarantees that a subset of safety critical queries are handled responsibly.
 In theory these verification techniques we propose may be applied to guardrails at different stages in the system, and we plan to explore this in future work.
@@
 In 2020, the International Verification of Neural Networks Competition [156] (VNN-COMP) was established to facilitate comparison between existing approaches, bring researchers working on the DNN verification problem together, and help shape future directions of the field.
-However, on its fourth edition, the competition still lacked NLP verification benchmarks [16].
-We propose to use this work for creating NLP verification benchmarks for future editions, to spread the awareness and attention to this field.
+However, for some time, the competition still lacked NLP verification benchmarks [157].
+In 2024, we contributed a first NLP benchmark to VNN-COMP, using the methodology of this paper, and the tool ANTONIO [158].
+We intend to use this work for creating NLP verification benchmarks for future editions, to spread the awareness and attention to this field.
 Acknowledgements
-Authors acknowledge support of EPSRC grant AISEC EP/T026952/1 and NCSC grant Neural Network Verification: in search of the missing spec, Vinay Krupakaran for his help in the evaluating the perturbations and Daniel Kienitz for his help with the eigenspace rotation.
+Authors acknowledge support of EPSRC grant AISEC EP/T026952/1 and NCSC grant ‘Neural Network Verification: in search of the missing spec’, Vinay Krupakaran for his help in the manual annotation of the perturbed sentences and Daniel Kienitz for his help with the eigenspace rotation.
+The first author acknowledges the James Watt Scolarship awarded by the Heriot-Watt University.